### v1

# Online Convex Optimization with Memory and Limited Predictions

Abstract
We study the problem of online convex optimization with memory and predictions over a horizon $T$. At each time step, a decision maker is given some limited predictions of the cost functions from a finite window of future time steps, i.e., values of the cost function at certain decision points in the future. The decision maker then chooses an action and incurs a cost given by a convex function that depends on the actions chosen in the past. We propose an algorithm to solve this problem and show that the dynamic regret of the algorithm decays exponentially with the prediction window length. Our algorithm contains two general subroutines that work for wider classes of problems. The first subroutine can solve general online convex optimization with memory and bandit feedback with $\sqrt{T}$-dynamic regret with respect to $T$. The second subroutine is a zeroth-order method that can be used to solve general convex optimization problems with a linear convergence rate that matches the best achievable rate of first-order methods for convex optimization. The key to our algorithm design and analysis is the use of truncated Gaussian smoothing when querying the decision points for obtaining the predictions. We complement our theoretical results using numerical experiments.
1 Introduction
Online decision making is a ubiquitous framework that captures a wide range of real-world applications, such as machine learning, control systems and finance [1]. In this framework, a decision maker interacts with an unknown environment by sequentially choosing a decision point $x_{t}$ at each time step and incurs a cost $f_{t}(x_{t})$ given by some underlying cost function $f_{t}(\cdot)$, and the goal is to minimize the accumulative cost over a horizon $t=1,\dots,T$. For convex cost functions $f_{t}(\cdot)$, the problem is known as Online Convex Optimization (OCO) which has been widely studied in the past decades (see, e.g., [2, 3] for comprehensive survey). The OCO framework may be further split into two settings. In the full information setting, the decision maker receives complete knowledge of the cost function $f_{t}(\cdot)$ after choosing $x_{t}$ for time step $t$ [4]; in the bandit information setting, the decision maker only receives $f_{t}(x_{t})$ (and potentially $f_{t}(\bar{x}_{t})$ at some perturbed points $\bar{x}_{t}$ from $x_{t}$) after choosing $x_{t}$ for time step $t$ [5, 6, 7]. Since the decision maker needs to choose the decision points in an online manner (i.e., before knowing all the cost functions), regret is a typical performance metric used for algorithms proposed to solve OCO, which compares the accumulative cost incurred by the algorithm to the minimum achievable cost in hindsight [4]. For non-stationary environment [8], the minimum achievable cost typically requires time-varying decisions over the horizon $T$ and the corresponding regret is referred to as dynamic regret [9, 10, 11].
In many online decision making problems, the cost incurred at the current time step depends not only on the decision made for the current time step but also on the decisions made in the past. For example, when the underlying environment is characterized by a dynamical system whose state evolves according to the chosen actions, the cost incurred in the current step is governed by the current action and system state which in turn depends on past actions and system states. Such a framework of OCO (also known as OCO with memory) has been studied in [12, 13, 14] and specialized to the online control problem of linear dynamical system [15, 16, 17, 18, 19]. Due to the long-term dependency of the cost on the decision history, i.e., the current decision also influences cost at future steps, there is a line of research that also considers predictions of the cost in a window of future time steps [20, 21, 22, 23, 24]. We refer to this framework as OCO with memory and predictions. However, the works mentioned above have certain limitations. In [21, 22], the cost is assumed to be only a function of decisions in the current step and one step in the past. The classical control algorithm, model predictive control (MPC) [25, 26, 27], considers the receding horizon optimization for the next $W$ stages. However, similar to [24, 23], the predictions of the full information of the cost functions are assumed to be available in the MPC algorithms.
Motivated by these limitations in the existing works, we study the problem of OCO with memory and limited (i.e., partial) predictions. Specifically, we consider the scenario that at each time step, the decision maker is allowed to only query a limited number of points and obtain potentially noisy predictions of the cost functions from a finite-length prediction window at these points, and the incurred cost has a long-term dependency on the past decisions. Studying the partial prediction setting is natural since it indeed corresponds to the partial (i.e., bandit) information setting that has been widely studied in the classic OCO framework (without memory or prediction) as we described above. Our approaches and contributions are summarized as follows.
$\bullet$ Algorithm Design and Regret Analysis. In Section 3, we propose an algorithm (Algorithm 1) to solve the problem of OCO with memory and limited predictions. Our algorithm extends that in [23] for full information setting to the partial information setting when receiving the predictions. We prove in Section 4 that the algorithm enjoys a dynamic regret that decays exponentially with the length of the prediction window and matches the dynamic regret of the algorithm in [23]. Our algorithm contains two general subroutines that can be applied to solve wider classes of problems as we describe next.
$\bullet$ OCO with Memory and Bandit feedback. The first subroutine of Algorithm 1 can be applied to the general OCO problem with memory and bandit feedback. We show in Theorem 1 our algorithm achieves a dynamic regret of $\tilde{\mathcal{O}}(\sqrt{T})$, where $\mathcal{O}(\cdot)$ hides logarithmic factors in $T$. Our dynamic regret matches that of the algorithm proposed in [8] for OCO with memory and full information of cost functions (up to logarithmic factors in $T$).
$\bullet$ Faster Convergence Rate of Zeroth-Order Method. The second subroutine of Algorithm 1 can be viewed as a zeroth-order (i.e., derivative-free) method for solving general (offline) convex minimization problems. Zeroth-order methods have typically been used to solve convex minimization problems when one only has access to a function value oracle while computing the derivative of the function is not allowed [28, 29]. For smooth and strongly convex cost functions, our algorithm achieves the same linear convergence rate as the first-order method using gradient descent [30, 31], which improves the convergence rate of the zeroth-order method proposed in [28]. We achieve this speedup by introducing a novel truncated Gaussian smoothing (formally introduced in Definition 1) when choosing the query points for the function value oracle, i.e., we perturb the decision point $x_{t}$ at time step $t$ by truncated Gaussian random vectors and query the perturbed points to obtain predictions of the function values.
$\bullet$ Noisy Function Value Oracle. Algorithm 1 and results in Theorems 1-2 hold for the setting when there is noise in the function value oracle used in Algorithm 1 to obtain the predictions. Such a noisy model subsumes the exact function value oracle as a special case.
More Related Work. First, Bandit OCO and zeroth-order methods for convex minimization are naturally related to each other since in both scenarios, a general recursive way to update the action has the form: $x_{t+1}=x_{t}-\alpha_{t}g_{t}$, where $g_{t}$ is obtained from the function value oracle and can be viewed as an estimate of the true gradient at $x_{t}$. A major difference is that the performance of algorithms for bandit OCO is characterized by regret while the performance of algorithms for zeroth-order methods is characterized by convergence rate. However, it has been shown that an algorithm for bandit OCO with regret guarantee can be converted to an algorithm for zeroth-order methods with convergence guarantee [32, 29]. In this work, we propose an algorithm that combines a bandit OCO algorithm and a zeroth-order algorithm to solve the OCO problem with memory and limited predictions. Second, our algorithm design and analysis rely on the use of truncated Gaussian smoothing. The methods of Gaussian smoothing or smoothing over a sphere have been used in bandit OCO or zeroth-order methods [5, 30, 33, 34]. We show that the truncated Gaussian smoothing inherits the desired properties of Gaussian smoothing while also allowing us to prove a faster convergence rate of the zeroth-order method for smooth and strongly convex cost functions. Finally, several extensions of the framework in this paper can be considered, including extending to OCO with unbounded memory (i.e., the current cost depends on all the past decisions) [35] and nonsmooth or nonconvex cost functions [36].
Notation and terminology. The sets of integers and real numbers are denoted as $\mathbb{Z}$ and $\mathbb{R}$, respectively. For a set $\mathcal{X}$, $\mathcal{X}^{n}$ denotes the corresponding Cartesian product $\mathcal{X}\times\cdots\times\mathcal{X}$. For a real number $a$, let $\lfloor a\rfloor$ be the largest integer that is smaller than or equal to $a$. For a vector $x\in\mathbb{R}^{n}$, let $\left\|x\right\|=\sqrt{x^{\top}x}$ be its euclidean norm. Let $\left\|P\right\|_{F}=\sqrt{\operatorname{\mathrm{tr}}(PP^{\top})}$ be the Frobenius norm of $P\in\mathbb{R}^{n\times m}$. Let $I_{n}$ be an $n$-dimensional identity matrix. Given any integer $n\geq 1$, $[n]\triangleq\{1,\dots,n\}$.
2 Problem Formulation
We consider an Online Convex Optimization (OCO) framework over $T\in\mathbb{Z}_{\geq 1}$ time steps. At each $t\in[T]$, a decision maker (i.e., agent) chooses an action $x_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ and incurs a cost $f_{t}(x_{t-h+1},\dots,x_{t})$, where $f_{t}:\mathbb{R}^{dh}\to\mathbb{R}$ is convex and $h\in\mathbb{Z}_{\geq 1}$. In other words, the cost incurred at any time step $t\in\mathbb{Z}_{\geq 1}$ has memory that depends on the current action $x_{t}$ and the past actions $x_{t-1},x_{t-2},\dots,x_{t-h+1}$. Denote
$C_{T}({\bm{x}})\triangleq\sum_{t=1}^{T}f_{t}(x_{t-h+1:t}),$ (1)
where ${\bm{x}}=(x_{1},\dots,x_{T})$, $x_{t-h+1:t}=(x_{t-h+1},x_{t-h+2},\dots,x_{t-1},x_{t})$ and assume throughout this paper that $x_{t}=\bar{x}_{0}$ for all $t\leq 1$ for some initial condition $\bar{x}_{0}\in\mathcal{X}$. Now, given the complete knowledge of the functions $f_{1}(\cdot),\dots,f_{T}(\cdot)$, one can solve
${\bm{x}}^{\star}\in\operatorname*{arg\,min}_{x_{1},\dots,x_{T}\in\mathcal{X}}C_{T}(\bm{x}),$ (2)
where $\bm{x}^{\star}=(x_{1}^{\star},\dots,x_{T}^{\star})$. In the online framework described above, since the agent does not have the complete knowledge of $f_{1}(\cdot),\dots,f_{T}(\cdot)$ before making the decision $x_{t}\in\mathcal{X}$ at time step $t\in\mathbb{Z}_{\geq 1}$, solving (2) is impossible. The goal is then to design an online algorithm $\mathcal{A}$ for the agent such that the following regret $R(\mathcal{A})$ is minimized:
$R(\mathcal{A})\triangleq C_{T}(\bm{x}^{\mathcal{A}})-C_{T}(\bm{x}^{\star}),$ (3)
where $\bm{x}^{\mathcal{A}}=(x_{1},\dots,x_{T})$ is chosen by the algorithm.
Even though solving (3) requires complete knowledge of $f_{1}(\cdot),\dots,f_{T}(\cdot)$ a priori, in practice one may have access to predictions of $f_{t}(\cdot),\cdots,f_{t+W-1}(\cdot)$ for some $W\in\mathbb{Z}_{\geq 1}$ when choosing $x_{t}\in\mathcal{X}$ at time step $t\in\mathbb{Z}_{\geq 1}$ [21, 22, 26, 27]. However, the available predictions can be inaccurate and scarce (i.e., limited). Unlike existing works as discussed in the introduction, in this paper, we focus on a general framework to capture the inaccuracy and scarcity of the predictions. Specifically, at each time step $t\in\mathbb{Z}_{\geq 1}$, the predictions can be obtained by first running sequences of actions $\{(x_{t}^{j},\dots,x_{t+W-1}^{j})\}_{j=1}^{p}$ on a simulator and then receiving the corresponding costs $\{l_{t}(x_{t-h+1:t}^{j}),\dots,l_{t+W-1}(x_{t+W-h:t+W-1}^{j})\}_{j=1}^{p}$, where $l_{k}:\mathcal{X}^{h}\to\mathbb{R}$ for all $k\in\{t,\dots,t+W-1\}$. In other words, the simulator is associated with a model $l_{t}(\cdot)$ that describes the true cost $f_{t}(\cdot)$ with some model misspecification and can be viewed as a noisy oracle of the function values of $f_{t}(\cdot)$. Based on the available information, the agent then chooses an action $x_{t}\in\mathcal{X}$. Note that for each time step $t\in\mathbb{Z}_{\geq 1}$, one may obtain multiple cost prediction trajectories from the simulator by running multiple sequences of actions on the simulator in parallel (or running a single trajectory on multiple simulators). To summarize, the agent does the following:
  • At the beginning of time step $t\in\mathbb{Z}_{\geq 1}$, run sequences of actions $\{(x_{t}^{j},\dots,x^{j}_{t+W-1})\}_{j=1}^{p}$ on simulator $l_{t}(\cdot)$ and receive the predictions $\{(l_{t}(x^{j}_{t-h+1:t}),\dots,l_{t+W-1}(x^{j}_{t+W-h:t+W-1})\}_{j=1}^{p}$.
  • Compute $x_{t}$ based on the predictions.
[…]
Remark 1.
Based on the above arguments, we see that the information setting considered in our framework corresponds to the bandit information in the classic OCO framework (without memory or predictions) [5, 7, 37, 38, 11], since only values of the cost function are available when the agent makes decisions.
3 Algorithm Design
To tackle the limited predictions (i.e., bandit information) in our problem setting, we adopt the zeroth-order (i.e., derivative-free) method that has been used for both bandit OCO and (offline) convex optimization problems (e.g., [5, 28]). To this end, we extend the algorithm design in [24, 20] for full information setting to bandit information setting. The detailed steps are summarized in Algorithm 1. Using similar arguments to those in [23], one can argue that the overall algorithm can be viewed as using a zeroth-order method for solving (2), which consists of two phases: initialization and zeroth-order method update. Specifically, denoting $\bm{x}^{j}\triangleq(x_{1}^{j},\dots,x_{T}^{j})$ for any $j\in\{0,\dots,K-1\}$, the algorithm first initializes $\bm{x}^{0}$ using an OCO scheme, and then iteratively computes the update $\bm{x}^{j+1}=\Pi_{\mathcal{X}}(\bm{x}^{j}-\alpha\bm{g}^{j})$ for $j=0,\dots,K-1$, where $\Pi_{\mathcal{X}}(\cdot)$ denotes the projection onto the convex set $\mathcal{X}^{T}$ and $\bm{g}^{j}=(\bm{g}^{j}_{1},\dots,\bm{g}^{j}_{T})\in\mathbb{R}^{dT}$ with $\bm{g}^{j}_{s}=\sum_{k=s}^{s+h-1}g_{k}^{j}$ for all $s\in[T]$. In particular, the term $\bm{g}^{j}$ is obtained from $l_{1}(\cdot),\dots,l_{T}(\cdot)$ queried at certain points (see lines 6 and 16 of Algorithm 1) which can be viewed as an estimate of the true gradient $\nabla C_{T}(\bm{x}^{j})$. Note that we simply set $f_{t}(\bm{x})=l_{t}(\bm{x})=0$ for all $t\leq 0$, all $t>T$ and all $\bm{x}\in\mathcal{X}^{h}$ in Algorithm 1, and recall that we have set $x_{t}^{0}=\bar{x}_{0}\in\mathcal{X}$ for all $t\leq 1$. We summarize several key features of Algorithm 1.
Input: $\bar{x}_{0}\in\mathcal{X},\eta_{t},\alpha,\delta,\delta^{\prime}$.
1 Pick $u^{0}\sim\widetilde{\mathcal{N}}(0,I_{d})$
2 for $t=2-W$ to $T$ do
       /* Initialize $x_{t+W}^{0}$ using OCO */
3       Let $r=t+W-1$.
4       Set $\bar{x}_{r}^{0}=x_{r}^{0}+\delta u^{0}$ and $\tilde{x}_{r}^{0}=x_{r}^{0}+\delta u^{0}$.
5       Receive predictions $l_{r}(\bar{x}^{0}_{r-h+1:r})$ and $l_{r}(\tilde{x}^{0}_{r-h+1:r})$.
6       Set $g_{r}^{0}=\frac{1}{2\delta}\big{[}l_{r}(\bar{x}_{r-h+1:r}^{0})-l_{s}(\tilde{x}_{r-h+1:r}^{0})\big{]}u^{0}$.
7       Update $x_{r+1}^{0}=\mathbf{\Pi}_{\mathcal{X}}(x_{r}^{0}-\eta_{r}g_{r}^{0})$.
       /* Obtain $x_{t}^{K}$ using zeroth-order method */
8       Let $K=\lfloor\frac{W}{h-1}\rfloor$.
9       for $j=0,\dots,K-1$ do
10             Let $s=t+W-(j+1)(h-1)-(W-(h-1)K)=t+(K-j-1)(h-1)$.
11             if $j=0$ then
12                   Pick $u_{s}^{j}\overset{i.i.d}{\sim}\widetilde{\mathcal{N}}(0,I_{d})$.
13                   Set $\bar{x}_{s+h-1}^{j}=x_{s+h-1}^{j}+\delta^{\prime}u_{s}^{j}$ and $\tilde{x}_{s+h-1}^{j}=x_{s+h-1}^{j}+\delta^{\prime}u_{s}^{j}$.
14                   Receive predictions $l_{s+h-1}(\bar{x}^{j}_{s:s+h-1})$ and $l_{s+h-1}(\tilde{x}^{j}_{s:s+h-1})$.
15
16            for $k=s,\dots,s+h-1$ do
17                   Set $g_{k}^{j}=\frac{1}{2\delta^{\prime}}\big{[}l_{k}(\bar{x}_{k-h+1:k}^{j})-l_{k}(\tilde{x}_{k-h+1:k}^{j})\big{]}u_{s}^{j}$.
18            Update $x_{s}^{j+1}=\mathbf{\Pi}_{\mathcal{X}}\big{(}x_{s}^{j}-\alpha\sum_{k=s}^{s+h-1}g_{k}^{j}\big{)}$.
19             Pick $u_{s}^{j+1}\overset{i.i.d}{\sim}\widetilde{\mathcal{N}}(0,I_{d})$.
20             Set $\bar{x}_{s}^{j+1}=x_{s}^{j+1}+\delta^{\prime}u_{s}^{j+1}$ and $\tilde{x}_{s}^{j+1}=x_{s}^{j+1}+\delta^{\prime}u_{s}^{j+1}$.
21             Receive $l_{s}(\bar{x}^{j+1}_{s-h+1:s})$ and $l_{s}(\tilde{x}^{j+1}_{s-h+1:s})$.
22
23      Play $x_{t}^{K}$ and incur the cost $f_{t}(x_{t-h+1:t}^{K})$.
Output: $x_{1}^{K},\dots,x_{T}^{K}$.
Algorithm 1 OCO with Memory and Limited Predictions
Two-Point Prediction via $l_{t}(\cdot)$. Note that in line 5, $\bar{x}^{0}_{r-h+1:r}=(x^{0}_{r-h+1},x^{0}_{r-h+2},\dots,x^{0}_{r-1},\bar{x}^{0}_{r})$ and $\tilde{x}^{0}_{r-h+1:r}=(x^{0}_{r-h+1},x^{0}_{r-h+2},\dots,x^{0}_{r-1},\tilde{x}^{0}_{r})$, where $\bar{x}^{0}_{r}$, $\tilde{x}^{0}_{r}$ are set in line 4. Similar arguments hold for lines 14 and 20. For each call of line 5, line 14 or line 20, Algorithm 1 will query two points for $l_{t}(\cdot)$ to obtain the corresponding predictions. For any $t\in\{2-W,\dots,T\}$, one can check that to obtain $x_{t}^{K}$ in line 21 of Algorithm 1, the algorithm needs to use function values of $l_{t}(\cdot)$ contained in the set $\mathcal{I}_{t}$ defined as
$\mathcal{I}_{t}\triangleq\big{\{}l_{k}(\bar{x}_{k-h+1:k}^{j}),l_{k}(\tilde{x}_{k-h+1:k}^{j}):k\in\{t+(K-j-1)h,\dots,t+W-1\},j\in\{0,\dots,K-1\}\big{\}}.$ (4)
One can also check that all the function values in $\mathcal{I}_{t}$ are available via the predictions in lines 5, 14 and 20 of Algorithm 1. Moreover, at any $t\in\{2-W,\dots,T\}$, Algorithm 1 only needs to maintain in its memory those functions values in $\mathcal{I}_{t}$ (that are needed to obtain $x_{t}^{K}$) and remove the other function values from its memory. In other words, the memory of Algorithm 1 may be recursively updated through receiving fresh predictions via $l_{t}(\cdot)$ and removing the function values of $l_{t}(\cdot)$ that are no longer useful. Finally, we note that for $t=2-W,\dots,T$, the points associated with the predictions in lines 5, 14 and 20 of Algorithm 1 correspond to consecutive points on $2(K+1)$ trajectories given by $\{\bar{x}_{1}^{j},\bar{x}_{2}^{j},\dots\}_{j=0}^{K}$ and $\{\tilde{x}_{1}^{j},\tilde{x}_{2}^{j},\dots\}_{j=0}^{K}$, and thus we do not need to restart the simulator $l_{t}(\cdot)$ during the course of Algorithm 1.
Truncated Gaussian Smoothing. In Algorithm 1 (lines 4, 13 and 19), we use a truncated Gaussian random vector denoted as $u_{s}^{j}\sim\widetilde{\mathcal{N}}(0,I_{d})$ to perturb the point $x_{s}^{j}$ to obtain $\bar{x}_{s}^{j}$ and $\tilde{x}_{s}^{j}$ which are then fed to the simulator $l_{t}(\cdot)$. Specifically, for any $i\in[d]$, we let the $i$th margin $u_{s,i}^{j}$ of $u_{s}^{j}$ to be a truncated Gaussian random variable obtained from the normal distribution such that $u_{s,i}^{j}$ belongs to the interval $[-(1/2d^{2}(2h-1))^{1/4},1/(2d^{2}(2h-1))^{1/4}]$ and the pdf of $u_{s,i}^{j}$ (i.e., the $i$th margin of $\widetilde{\mathcal{N}}(0,I_{d})$) is given by
$p(x)=\frac{1}{\sqrt{2\pi}\kappa}e^{-\frac{x^{2}}{2}}\ \text{and}\ \kappa\ \text{is such that}\ \int_{-1/(2d^{2}(2h-1))^{\frac{1}{4}}}^{1/(2d^{2}(2h-1))^{\frac{1}{4}}}p(x)=1.$ (5)
It follows that $\lVert u_{s}^{j}\rVert\leq 1/(2(2h-1))^{1/4}$. Formally, we introduce the following [39].
Definition 1.
We use $u\sim\widetilde{\mathcal{N}}(0,I_{d})$ to represent an $n$-dimensional truncated Gaussian random vector whose $i$th margin $u_{i}$ satisfies $u_{i}\in[-1/(2d^{2}(2h-1))^{1/4},1/(2d^{2}(2h-1))^{1/4}]$ and the pdf of $u_{i}$ is given by (5) for all $i\in[d]$.
There exist efficient methods for sampling from a truncated Gaussian distribution (see e.g. [40]). As we discussed in the introduction, using the truncated Gaussian allows us to prove the desired regret guarantee of Algorithm 1 in the next section.
4 Regret Analysis
We now analyze the regret of Algorithm 1. To proceed, we introduce the following definitions and assumptions, which are standard in the OCO literature [6, 15, 8, 5].
Definition 2.
A continuous differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $\beta$-smooth if
$\left\|\nabla f(x)-\nabla f(y)\right\|\leq\beta\left\|x-y\right\|,\ \forall x,y\in\mathbb{R}^{d}.\$
Definition 3.
A differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $\mu$-strongly convex if
$f(x)-f(y)\leq\langle x-y,\nabla f(x)\rangle-\frac{\mu}{2}\left\|x-y\right\|^{2},\ \forall x,y\in\mathbb{R}^{d}.$
Definition 4.
A function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $G$-Lipschitz if
$|f(x)-f(y)|\leq G\left\|x-y\right\|,\ \forall x,y\in\mathbb{R}^{d}.$
Assumption 1.
For any $t\in[T]$, $f_{t}(\cdot)$ is (a) $\beta$-smooth, (b) $\mu$-strongly convex and (c) $G$-Lipschitz.
Assumption 2.
The feasible set $\mathcal{X}\subseteq\mathbb{R}^{d}$ is convex and satisfies that $\max_{x,y\in\mathcal{X}}\left\|x-y\right\|\leq D$.
We will also assume that the model error of $l_{t}(\cdot)$ is uniformly bounded.
Assumption 3.
For any $t\in[T]$, there exists $\varphi_{t}\in\mathbb{R}_{\geq 0}$ such that $|l_{t}(\bm{x})-f_{t}(\bm{x})|\leq\varphi_{t}$ for all $\bm{x}\in\mathcal{X}^{h}$.
Remark 2.
Note that the noisy model in Assumption 3 is equivalent to assuming that $l_{t}(\bm{x})=f_{t}(\bm{x})+\phi_{t}$ for all $\bm{x}\in\mathcal{X}^{h}$ with $\phi_{t}\in\mathbb{R}$ and $\left|\phi_{t}\right|\leq\varphi_{t}$, where $\phi_{t}$ can be either deterministic or stochastic. The bounded noise model has also been considered in e.g. [33]. The potentially unbounded stochastic noise model has been considered in e.g. [41] and we leave the extension to such a noise model to future work.
We define several auxiliary functions that will be useful in our later analysis.
Definition 5.
(a) For any $x\in\mathbb{R}^{d}$ and any $t\in[T]$, define the $h$-unitary function $f_{t}^{c}(x)\triangleq f_{t}(x,\dots,x)$ and its smoothed version $\hat{f}_{t}^{c}(x)\triangleq\mathbb{E}_{v}[f_{t}^{c}(x+\delta v)]$, where $v\sim\widetilde{\mathcal{N}}(0,I_{d})$ and $\delta\in\mathbb{R}_{>0}$. [^2] Throughout the paper, we use $\mathbb{E}_{x}[\cdot]$ (resp., $\mathbb{E}[\cdot]$) to denote the expectation with respect to a certain random variable (or vector) $x$ (resp., all the randomness in Algorithm 1).
(b) For any $\bm{x}=(x_{1},\dots,x_{T})\in\mathbb{R}^{dT}$, $\hat{C}_{T}(\bm{x})\triangleq\mathbb{E}_{\bm{v}}[C(\bm{x}+\delta^{\prime}\bm{v})]$, where $\bm{v}\sim\widetilde{\mathcal{N}}(0,I_{dT})$ and $\delta^{\prime}\in\mathbb{R}_{>0}$.
The following result shows that the functions $\hat{f}_{t}^{c}(\cdot)$ and $\hat{C}_{T}(\cdot)$ inherit the properties from $f_{t}(\cdot)$; the proof can be found in Appendix A.
Lemma 1.
Suppose Assumption 1(a) holds for $f_{t}(\cdot)$ for all $t\in[T]$. Then, (a) $\hat{f}_{t}^{c}(\cdot)$ is $\beta$-smooth and $\mu$-strongly convex; (b) $C_{T}(\cdot)$ is $\beta h$-smooth, $\mu$-strongly convex and $G\sqrt{Th}$-Lipschitz; and (c) $\hat{C}_{T}(\cdot)$ is $\beta h$-smooth and $\mu$-strongly convex.
More importantly, one can form unbiased gradient estimators of $\hat{f}_{t}^{c}(\cdot)$ and $\hat{C}_{T}(\cdot)$ using only the function values. Formally, we have the following result; the proof can be found in Appendix A.
Lemma 2.
Under Assumption 1(a), the following hold:
(a) For $f_{t}^{c}(\cdot)$ and $\hat{f}_{t}^{c}(\cdot)$ given by Definition 5(a), any $x\in\mathbb{R}^{d}$ and any $\delta\in\mathbb{R}_{>0}$,
$\nabla\hat{f}_{t}^{c}(x)=\mathbb{E}_{v}\Big{[}\frac{f_{t}^{c}(x+\delta v)-f_{t}^{c}(x-\delta v)}{2\delta}v\Big{]}$ (6)
and
$|f_{t}^{c}(x)-\hat{f}_{t}^{c}(x)|\leq\frac{\delta^{2}}{2}\beta d.$ (7)
(b)For $C_{T}(\cdot)$ and $\hat{C}_{T}(\cdot)$ given by Definition 5(b), any $\bm{x}=(x_{1},\dots,x_{T})\in\mathbb{R}^{dT}$ and any $\delta^{\prime}\in\mathbb{R}_{>0}$,
$\frac{\partial\hat{C}_{T}(\bm{x})}{\partial x_{s}}=\mathbb{E}_{v_{s}}\Big{[}\sum_{k=s}^{s+h-1}\frac{f_{k}(\bar{x}_{k-h+1:k})-f_{k}(\tilde{x}_{k-h+1:k})}{2\delta^{\prime}}v_{s}\Big{]}\ \forall s\in[T],$ (8)
$|C_{T}(\bm{x})-\hat{C}_{T}(\bm{x})|\leq\frac{\delta^{\prime 2}}{2}\beta hTd,$ (9)
and
$\left\|\nabla\hat{C}_{T}(\bm{x})-\nabla C_{T}(\bm{x})\right\|\leq\frac{\delta^{\prime}}{2}\beta h(Td+3)^{3/2},$ (10)
where $\bm{v}=(v_{1},\dots,v_{T})\sim\widetilde{\mathcal{N}}(0,I_{dT})$, $\bar{x}_{t}=x_{t}+\delta^{\prime}v_{t}$ and $\tilde{x}_{t}=x_{t}-\delta^{\prime}v_{t}$.
Recall from Eq. (3) that the regret of Algorithm 1 is given by $R(\mathcal{A})=C_{T}(\bm{x}^{\mathcal{A}})-C_{T}(\bm{x}^{\star})$, where $\bm{x}^{\mathcal{A}}=(x_{1}^{K},\dots,x_{T}^{K})$ is chosen by the algorithm and $\bm{x}^{\star}=(x_{1}^{\star},\dots,x_{T}^{\star})$ is the optimal solution to (1). Since Algorithm 1 contains two phases, we will split our analysis accordingly.
Initialization Phase: We first upper bound the contribution of the initialization step in Algorithm 1 to $R(\mathcal{A})$, where we note from the definition of Algorithm 1 that $x_{1}^{0},\dots,x_{T}^{0}$ are chosen by the initializing phase of the algorithm. The proof of Theorem 1 can be found in Appendix B.
Theorem 1.
Suppose Assumptions 1-3 hold. Let $\eta_{t}=\frac{1}{t\mu}$ and $\delta=1/\sqrt{T}$ in Algorithm 1. Then, the initialization step of Algorithm 1 satisfies that
$\mathbb{E}\Big{[}\sum_{t=1}^{T}f_{t}(x_{t-h+1:t}^{0})-f_{t}(x_{t-h+1:t}^{\star})\Big{]}\leq\Big{(}\frac{\sqrt{2}D}{\mu}+Gh^{2}\Big{)}V_{T}+T\delta^{2}\beta d$
$+(\frac{8G^{2}h^{2}+hG^{2}}{2\mu(2h-1)}+\frac{h^{3}G^{2}D}{\delta(2(2h-1))^{1/2}}+\frac{3G^{2}h^{3}}{\mu(2(2h-1))^{1/2}})(1+\ln T)$
$+\frac{2}{\delta^{2}\mu\sqrt{2h-1}}\sum_{t=1}^{T}\varphi_{t}^{2}+(\frac{\sqrt{2}h^{2}GD}{2\delta(2(2h-1))^{1/2}}+\frac{\sqrt{2}Gh^{2}+D\mu}{\delta\mu(2(2h-1))^{1/4}})\sum_{t=1}^{T}\varphi_{t}$ (11)
with $V_{T}\triangleq\mathbb{E}\big{[}\sum_{t=1}^{T-1}\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|\big{]}$, where $\mathbb{E}[\cdot]$ denotes the expectation with respect to the randomness in Algorithm 1 and $(x_{1}^{\star},\dots,x_{T}^{\star})$ is an optimal solution to (1).
Recalling the definition of Algorithm 1, the initialization phase of Algorithm 1 can be viewed as being applied to the cost functions $f_{1}(\cdot),\dots,f_{T}(\cdot)$ (without using any prediction) and returning $x_{1}^{0},\dots,x_{T}^{0}$. To be more specific, the initialization phase of Algorithm 1 solves the problem of OCO with memory, where the cost functions are given by $f_{t}(\cdot),\dots,f_{T}(\cdot)$ and one can check that by line 7 of Algorithm 1, the decision points $x_{1}^{0},\dots,x_{T}^{0}$ are obtained as
$x_{t+1}^{0}=\Pi_{\mathcal{X}}(x_{t}^{0}-\eta_{t}g_{t}^{0}),$
where $g_{t}^{0}$ is given by line 6 of Algorithm 1. As we argued in Remark 1, our OCO with memory framework corresponds to the bandit information setting, i.e., $g_{t}^{0}$ is computed based on only the (noisy) cost function values. The resulting dynamic regret of the initialization step applied to the problem of bandit OCO with memory is then given by (11) in Theorem 1. The OCO with memory framework has been studied in [12, 13, 8]. We extend the existing algorithm design and regret analysis to the bandit OCO with memory framework, where the function value oracle can also be noisy. In particular, setting $\varphi_{t}=0$ for all $t\geq 0$ (i.e., there is no noise in the oracle of the function values of $f_{t}(\cdot)$), the regret bound in (11) reduces to $\tilde{\mathcal{O}}(V_{T}\sqrt{T})$ with $\tilde{\mathcal{O}}(\cdot)$ hiding logarithmic factors in $T$, where the scaling of $T$ matches the regret bound for the algorithm proposed in [8] for (full information) OCO with memory (i.e., the cost functions are fully accessible). Note that the $V_{t}$ factor in the regret bound in (11) can potentially be refined to $\sqrt{V_{T}}$ using the techniques from [10, 8]. Specifically, supposing $V_{T}$ (or an upper bound on $V_{T}$) is known, then setting the step size $\eta=\frac{1}{t\mu\sqrt{V_{T}}}$ in Algorithm 1 yields a regret bound of $\tilde{\mathcal{O}}(\sqrt{V_{T}T})$.
Zeroth-Order Update Phase: Next, we upper bound the contribution of the zeroth-order method used in Algorithm 1 to $R(\mathcal{A})$. As we argued before in Section 3, the iterates $\bm{x}^{j}=(x_{1}^{j},\dots,x_{T}^{j})$ for $j=0,\dots,K-1$ in Algorithm 1 are obtained from the update rule $\bm{x}^{j+1}=\bm{x}^{j}-\alpha\bm{g}^{j}$ starting from the initial point $\bm{x}^{0}$ (given by the initialization phase of Algorithm 1), where $\bm{g}^{j}=(\bm{g}^{j}_{1},\dots,\bm{g}^{j}_{T})\in\mathbb{R}^{dT}$ with $\bm{g}^{j}_{s}=\sum_{k=s}^{s+h-1}g_{k}^{j}$ for all $s\in[T]$ and $\bm{g}^{j}$ can be viewed as an estimate of the true gradient $\nabla C_{T}(\bm{x}^{j})$ based on the function values of $l_{t}(\cdot)$ evaluated at certain points perturbed by the truncated Gaussian smoothing (see our arguments in Section 3). Thus, the following result indeed characterizes the convergence of the above update rule to the optimal solution to problem (1), given the initial point $\bm{x}^{0}=(x_{1}^{0},\dots,x_{T}^{0})$ returned by the initialization phase of Algorithm 1. The proof of Theorem 2 can be found in Appendix B.
Theorem 2.
Consider the same setting as Theorem 1 and further let $\alpha=1/\beta^{\prime}$ in Algorithm 1, where $\beta^{\prime}=\beta h$ with $\beta$ given by Assumption 1. The output $\bm{x}^{K}=(x_{1}^{K},\dots,x_{T}^{K})$ of Algorithm 1 satisfies that
$\mathbb{E}\big{[}C_{T}(\bm{x}^{K})-C_{T}(\bm{x}^{\star})\big{]}\leq\Big{(}\frac{1}{1+\gamma}\Big{)}^{K}\big{(}C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})\big{)}+\frac{\varepsilon}{\gamma},$ (12)
where $\gamma\triangleq\mu/(\beta^{\prime}-\mu)$ [^3] For any $\beta$-smooth and $\mu$-strongly convex function $f(\cdot)$, it is known that $\beta>\mu$ holds (e.g., [30]). and
$\varepsilon\triangleq\frac{D\delta^{\prime}\beta h\sqrt{T}}{2\sqrt{2}(2(2h-1))^{3/4}}+\sqrt{h\beta GT\delta^{\prime}}D(Td+3)^{3/4}+\frac{Dh\sum_{t=1}^{T}\varphi_{t}}{\delta^{\prime}(2(2h-1))^{1/4}}.$
Thus, the convergence of the first term on the right-hand-side of (12) achieves the same linear convergence rate (i.e., $(1-\mu/\beta^{\prime})^{K}$) as standard first-order method based on gradient descent for smooth and strongly convex cost functions [30]. Supposing $\varphi_{t}=0$ for all $t\geq 0$, the second term on the right-hand-side of (12) can be made arbitrarily small by setting $\delta^{\prime}$. A careful inspection of the proof of Theorem 2 reveals that the result in Theorem 2 still holds if we remove Assumption 1(c). The reason is that in the proof, we only need an upper bound on $\nabla C_{T}(\bm{x})$ for all $\bm{x}\in\mathcal{X}^{dT}$. Since $C_{T}(\cdot)$ is $\beta\sqrt{2h}$-smooth from Lemma 1, i.e., $\nabla C_{T}(\cdot)$ is continuous, and $\mathcal{X}^{dT}$ is bounded by Assumption 2, such an upper bound on $\nabla C_{T}(\cdot)$ always exists. Based on the above arguments, under the smoothness and strong convexity assumption and setting $\varphi_{t}=0$ for all $t\geq 0$, our convergence rate given in Theorem 2 for the zeroth-order subroutine in Algorithm 1 improves that of the zeroth-order method provided in [28]. [^4] Since [28] considers the setting of $\varphi_{t}=0$, we also $\varphi_{t}=0$ for comparison. Specifically, the zero-order method in [28] also has the form $\bm{x}^{j+1}=\bm{x}^{j}-\alpha\bm{g}^{j}$ as we described above, where $\bm{g}^{j}$ is obtained based on function values at certain points perturbed by a Gaussian smoothing, i.e., $u_{s}^{j+1}$ in line 14 of Algorithm 1 is drawn from a Gaussian distribution $\mathcal{N}(0,I_{d})$. Setting the step size to satisfy $\alpha=\frac{1}{4(T+4)\beta^{\prime}}$, where $Td$ is the dimension of the variable $\bm{x}^{j}$, [28] provides a convergence
$\mathbb{E}\big{[}C_{T}(\bm{x}^{K})-C_{T}(\bm{x}^{\star})\big{]}\leq\Big{(}1-\frac{\mu}{4(T+4)\beta^{\prime}}\Big{)}^{K}\big{(}C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})\big{)}+\mathcal{O}(\frac{\delta^{\prime}T^{2}d^{2}\beta^{\prime}}{\mu}).$ (13)
Comparing (13) to (12) and noting that $\gamma=\mu/(\beta^{\prime}-\mu)$, we see that Algorithm 1 achieves a faster convergence around a neighborhood of the optimal solution $C_{T}(\bm{x}^{\star})$ and the convergence rate of Algorithm 1 does not degrade as the horizon length $T$ increases.
Overall Bound on the Regret of Algorithm 1. Combining the result in Theorems 1-2 gives the desired dynamic regret of Algorithm 1 for solving OCO with memory and limited predictions.
5 Numerical Experiments
This section provides numerical results to complement our theoretical analysis. Recalling that Algorithm 1 uses two-point feedback via the simulator $l_{t}(\cdot)$ and receives the predictions, for comparison, we also consider our algorithm in the single-point feedback case which has also been considered in the classic bandit OCO framework [5, 38]. Specifically, in lines 4 and 5 in Algorithm 1, we set $\bar{x}_{s}^{0}=x_{s}^{0}+\delta u^{0}$ and receive the prediction $l_{s}(\bar{x}^{0}_{s-h+1:s})$. Next, set $g_{s}^{0}=\frac{1}{\delta}\big{[}l_{s}(\bar{x}_{s-h+1:s}^{0})\big{]}u^{0}$ and then update in the same way as line 7. Similarly, for lines 9-20 of Algorithm 1, we modify the algorithm into the case of single-point feedback. Note that in our experiments we simply set $\varphi_{t}=0$ in Assumption 3, i.e., $l_{t}(\bm{x})=f_{t}(\bm{x})$ for all ${\bm{x}}\in\mathcal{X}^{d}$ and all $t\geq 0$.
Now we consider an unconstrained quadratic programming problem given by $\min_{x}C(x)=\frac{1}{2}x^{\top}Px+q^{\top}x$. Let $C(x)=\sum_{t=1}^{T}f_{t}(x_{t-h+1:t})$ and $f_{t}(x_{t-h+1:t})=\frac{1}{2}x_{t-h+1:t}^{\top}A_{t}x_{t-h+1:t}+B_{t}^{\top}x_{t-h+1:t}$. In our experiments, we consider $x_{t}\in\mathbb{R}$ and set $h=2$ and then randomly generate positive definite matrices $A_{t}$ and $B_{t}$ to make sure that Assumption 1(a) and (b) is true. Next, we set $x_{0}=0.5,\eta_{t}=0.2/t,\alpha=0.05,\delta=0.2,\delta^{\prime}=0.0001$ which satisfy the requirements stated in Theorems 1-2. Then, we consider the following three distributions for smoothing. (a) Gaussian distribution: In line 3 and line 14 of Algorithm 1, we instead pick a random variable $u\sim{\mathcal{N}}(0,1)$. (b) Truncated Gaussian distribution: To obtain the simulation results, we slightly modify Definition 1 and pick a truncated Gaussian from the interval $[-2,2]$ in lines 4 and 16 of Algorithm 1. (c) Bernoulli distribution: Note that for scalar $x_{t}\in\mathbb{R}$, the method of smoothing over a sphere is equivalent to that the random variable $u$ is drawn from a Bernoulli distribution with a $\frac{1}{2}$ probability of getting $1$ or $-1$. Hence, we pick a random variable $u$ from the Bernoulli in line 3 and line 14 of Algorithm 1. The results obtained below for Gaussian smoothing and Bernoulli smoothing are averaged over $200$ independent experiments, and the results for truncated Gaussian smoothing are averaged over $50$ independent experiments. The shaded areas in all the plots below represent quartiles.
(a) a
(b) c
(c) e
(d) b
(e) d
(f) f
Figure 1: Regret results for the initialization phase of Algorithm 1.
To verify Theorem 1, we gradually increase $T$ from $5$ to $20$ to get plots of $Reg/\sqrt{T}$, $Reg/T$ with $T$ when the random variable used for smoothing is drawn from Gaussian distribution and truncated Gaussian distribution and in both the single-point feedback and two-point feedback cases (See Figure 1). The $Reg$ here represents the regret corresponding to the initialization phase of Algorithm 1, i.e., when OCO is used in initialization. Since the regret bound in Theorem 1 scales as $\tilde{\mathcal{O}}(\sqrt{T}\ln{T})$, we first divide the regret value by $\sqrt{T}$ to get a curve that approximates the log function and then, we divide the regret value by $T$ to get a declining curve. By comparing these figures, it can be seen that the result of two-point feedback is better than that of single-point feedback, and the result when the random variable is drawn from the truncated Gaussian distribution is better than that from the Gaussian distribution. Moreover, the noise fluctuation in the Gaussian distribution is much larger than that in the truncated Gaussian distribution. Finally, the curve in Figure 1(e) changes very slowly, and roughly aligns with the curve of the logarithmic function $\ln{T}$. Figure 1(f) shows a declining curve that satisfies the characteristics of the function $\ln{T}/\sqrt{T}$. Hence, our experimental results are consistent with the theoretical analysis.
(a) a
(b) c
(c) e
(d) b
(e) d
(f) f
Figure 2: Overall regret results of Algorithm 1.
To verify Theorem 2, we set $T=20$ and $W$ is gradually increased to obtain the results that are plots of the value $\log{(Reg)}$ with $W$ when the random variable used for smoothing is drawn from the three distributions described above and in both the single-point feedback and two-point feedback settings. The $Reg$ here represents the overall regret of Algorithm 1 which consists of the two phases. As can be seen from Figure 2, the performance of two-point feedback is better than that of single-point feedback in general, and the performance of truncated Gaussian distribution is better than that of Gaussian distribution. According to Figure 2(e), the regret of our algorithm decays approximately linearly on a log scale, indicating the exponential decay with $W$. Comparing Figure 2(d) and Figure 2(e), truncated Gaussian distribution achieves better regret value and faster convergence rate than Gaussian distribution as we discussed at the end of Section 4, and all these above are consistent with the theoretical analysis of the algorithm. Finally, the performance of using the Bernoulli random variable for smoothing is comparable to the case of using the truncated Gaussian random variable, however, there is no theoretical guarantee of the regret for Bernoulli smoothing which is left to future work.
6 Conclusion
We studied the problem of OCO with memory and predictions over a horizon $T$. We proposed an algorithm to solve this problem and showed that the dynamic regret of the algorithm decays exponentially with the prediction window length. The first subroutine of the algorithm can be used to solve general online convex optimization with memory and bandit feedback with $\sqrt{T}$-dynamic regret with respect to $T$. The second subroutine of the algorithm is a zeroth-order method that can be used to solve general convex optimization problems with a linear convergence rate that improves the rate of zeroth-order algorithms and matches the rate of first-order algorithms for convex optimization problems. Our algorithm design and analysis rely on truncated Gaussian smoothing when querying the decision points for obtaining the predictions. Future work includes extending the algorithm design and analysis to more general (i.e., nonconvex or nonsmooth) cost functions and exploring other possible smoothing methods for zeroth-order algorithms.
Appendix A Proofs of Preliminary Results
A.1 Proof of Lemma 1
We first prove the properties of $\hat{f}_{t}^{c}(\cdot)$. Note that the $\mu$-strongly convexity of $f^{c}_{t}(\cdot)$ follows directly from that of $f_{t}(\cdot)$. Thus, for any $x,y\in\mathbb{R}^{d}$,
$f_{t}^{c}(x+\delta v)-f_{t}^{c}(y+\delta v)\leq\nabla f_{t}^{c}(x+\delta v)^{\top}(x-y)-\frac{\mu}{2}\left\|x-y\right\|^{2}.$
Taking the expectation with respect to $v\sim\widetilde{\mathcal{N}}(0,I_{d})$ on both sides of the above inequality, we get
$\hat{f}_{t}^{c}(x)-\hat{f}_{t}^{c}(y)$ $\leq\mathbb{E}_{v}\big{[}\nabla f_{t}^{c}(x+\delta v)\big{]}^{\top}(x-y)-\frac{\mu}{2}\left\|x-y\right\|^{2}$
$\leq\nabla\hat{f}_{t}^{c}(x)^{\top}(x-y)-\frac{\mu}{2}\left\|x-y\right\|^{2},$ (A.1)
where we can swap the order of $\mathbb{E}_{v}$ and $\nabla$ since $\nabla f_{t}^{c}(\cdot)$ is continuous in $x$. It follows from (A.1) that $\hat{f}_{t}^{c}(\cdot)$ is $\mu$-strongly convex. Similarly, the $\beta$-smoothness of $f_{t}^{c}(\cdot)$ follows from that of $f_{t}(\cdot)$. For any $x,y\in\mathbb{R}^{d}$, we have
$\left\|\nabla f_{t}^{c}(x+\delta v)-\nabla f_{t}^{c}(y+\delta v)\right\|\leq\beta\left\|x-y\right\|.$
Again, taking the expectation $\mathbb{E}_{v}[\cdot]$ on both sides of the above inequality yields
$\mathbb{E}_{v}\big{[}\left\|\nabla f_{t}^{c}(x+\delta v)-\nabla f_{t}^{c}(y+\delta v)\right\|\big{]}\leq\beta\left\|x-y\right\|,$
$\overset{(a)}{\Rightarrow}$ $\left\|\mathbb{E}_{v}[\nabla f_{t}^{c}(x+\delta v)]-\mathbb{E}_{v}[\nabla f_{t}^{c}(y+\delta v)]\right\|\leq\beta\left\|x-y\right\|,$
$\Rightarrow$ $\left\|\nabla\hat{f}_{t}^{c}(x)-\nabla\hat{f}_{t}^{c}(y)\right\|\leq\beta\left\|x-y\right\|,$
where (a) follows from Jensen’s inequality (i.e., $\mathbb{E}[\left\|a\right\|]\geq\left\|\mathbb{E}[a]\right\|$ for any $a\in\mathbb{R}^{d}$), which shows that $\hat{f}_{t}^{c}(\cdot)$ is $\beta$-smooth.
Next, we prove the properties of $C_{T}(\cdot)$. Recall that $C_{T}(\bm{x})=\sum_{t=1}^{T}f_{t}(x_{t-h+1:t})$, where $\bm{x}=(x_{1},\dots,x_{T})\in\mathbb{R}^{dT}$ and $x_{t}=x_{0}$ for all $t\leq 0$. Considering any $\bm{x}=(x_{1},\dots,x_{T})\in\mathbb{R}^{dT}$ and any $\bm{y}=(y_{1},\dots,y_{T})\in\mathbb{R}^{dT}$, we have
$\left\|\bm{x}-\bm{y}\right\|^{2}\leq\sum_{t=1}^{T}\left\|x_{t-h+1:t}-y_{t-h+1:t}\right\|^{2}.$ (A.2)
Moreover, we have
$\sum_{t=1}^{T}\left\|x_{t-h+1:t}-y_{t-h+1:t}\right\|^{2}\leq h\left\|\bm{x}-\bm{y}\right\|^{2},$ (A.3)
$\Rightarrow$ $\sqrt{\sum_{t=1}^{T}\left\|x_{t-h+1:t}-y_{t-h+1:t}\right\|^{2}}\leq\sqrt{h}\left\|\bm{x}-\bm{y}\right\|,$
$\overset{(a)}{\Rightarrow}$ $\sum_{t=1}^{T}\left\|x_{t-h+1:t}-y_{t-h+1:t}\right\|\leq\sqrt{Th}\left\|\bm{x}-\bm{y}\right\|,$ (A.4)
where (a) uses the fact that $\sqrt{\sum_{t=1}^{T}a_{t}^{2}}\geq\frac{\sum_{t=1}^{T}a_{t}}{\sqrt{T}}$ for any $a_{t}\in\mathbb{R}$. By the $\mu$-strongly convexity of $f_{t}(\cdot)$, for any $t\in[T]$,
$f_{t}(x_{t-h+1:t})-f_{t}(y_{t-h+1:t})\leq\nabla f_{t}(x_{t-h+1:t})^{\top}(x_{t-h+1:t}-y_{t-h+1:t})-\frac{\mu}{2}\left\|x_{t-h+1:t}-y_{t-h+1:t}\right\|^{2}.$ (A.5)
To relate $\nabla f_{t}(x_{t-h+1:t})$ to $\nabla C_{T}(\bm{x})$, we first augment the gradient vector $\nabla f_{t}(x_{t-h+1:t})\in\mathbb{R}^{dh}$ to a higher dimensional vector ~$\tilde{\nabla}f_{t}(x_{t-h+1:t})\in\mathbb{R}^{dT}$ by adding zero entries. By the definition of $C_{T}(\cdot)$, we see that $\nabla C_{T}(\bm{x})=\sum_{t=1}^{T}\tilde{\nabla}f_{t}(x_{t-h+1:t})$. Summing (A.5) over $t\in[T]$ now gives us
$C_{T}(\bm{x})-C_{T}(\bm{y})\leq\sum_{t=1}^{T}\tilde{\nabla}f_{t}(x_{t-h+1:t})^{\top}(\bm{x}-\bm{y})-\frac{\mu}{2}\sum_{t=1}^{T}\left\|x_{t-h+1:y}-y_{t-h+1:t}\right\|^{2},$
[…]
which proves that $C_{T}(\cdot)$ is $\mu$-strongly convex. To show the smoothness of $C_{T}(\cdot)$, we again start from $f_{t}(\cdot)$ and its $\beta$-smoothness gives us [^5] It is well known that a $\beta$-smooth function $f:\mathbb{R}^{d}\to\mathbb{R}$ as per definition 2 satisfies that $f(y)-f(x)\leq\left\langle\nabla f(x),y-x\right\rangle+\frac{\beta}{2}\left\|y-x\right\|^{2},\ \forall x,y\in\mathbb{R}^{d}$ [31].
$f_{t}(y_{t-h+1:t})-f_{t}(x_{t-h+1:t})\leq\nabla f_{t}(x_{t-h+1:t})^{\top}(y_{t-h+1:t}-x_{t-h+1:t})+\frac{\beta}{2}\left\|x_{t-h+1:t}-y_{t-h+1:t}\right\|^{2}.$ (A.6)
Summing (A.6) over $t\in[T]$ gives us
$C_{T}(\bm{y})-C_{T}(\bm{x})\leq\sum_{t=1}^{T}\tilde{\nabla}f_{t}(x_{t-h+1:t})^{\top}(\bm{y}-\bm{x})+\frac{\beta}{2}\sum_{t=1}^{T}\left\|x_{t-h+1:y}-y_{t-h+1:t}\right\|^{2},$
[…]
which proves that $C_{T}(\cdot)$ is $(\beta h)$-smooth. To show that that $C_{T}(\cdot)$ is $G\sqrt{Th}$-Lipschitz, we have
[…]
A.2 Proof of Lemma 2
Proof of part (a) Eq. (6) follows directly from [28, Eq. (26)]. Since $f_{t}^{c}(\cdot)$ is $\beta$-smooth by Assumption 1(a), Eq. (7) follows from [28, Theorem 1].
Proof of part (b) Recalling the definition of $\hat{C}_{T}(\cdot)$, we have
$\hat{C}_{T}(\bm{x})=\sum_{t=1}^{T}\mathbb{E}_{\bm{v}}[f_{t}(\bar{x}_{t-h+1:t})],$ (A.8)
where we set $\bar{x}_{t}=x_{0}$ for all $t\leq 0$. It follows that for any $s\in[T]$,
$\frac{\partial\hat{C}_{T}(\bm{x})}{\partial x_{s}}$ $=\frac{\partial\mathbb{E}_{\bm{v}}\big{[}\sum_{t=1}^{T}f_{t}(\bar{x}_{t-h+1:t})\big{]}}{\partial x_{s}}$
=$\overset{(a)}{=}\mathbb{E}_{\bm{v}}\Big{[}\sum_{t=1}^{T}\frac{\partial f_{t}(\bar{x}_{t-h+1:t})}{\partial x_{s}}\Big{]}$
$=\mathbb{E}_{\bm{v}}\Big{[}\sum_{k=s}^{s+h-1}\frac{\partial f_{k}(\bar{x}_{k-h+1:k})}{\partial x_{s}}\Big{]}$
=$\overset{(b)}{=}\frac{\partial\mathbb{E}_{\bm{v}}\big{[}\sum_{k=s}^{s+h-1}f_{k}(\bar{x}_{k-h+1:k})\big{]}}{\partial x_{s}},$ (A.9)
where we can swap the order of $\mathbb{E}_{\bm{v}}$ and $\frac{\partial}{\partial x_{s}}$ in (a) and (b) since $\frac{\partial f_{t}(\bar{x}_{t-h+1:t})}{\partial x_{s}}$ is continuous in $x_{s}$ for all $t\in[T]$. Now, we notice that $\sum_{k=s}^{s+h-1}f_{k}(x_{k-h+1:k})$ may be viewed as a function of $x_{s-h+1},\dots,x_{s+h-1}$, denoted as $F_{s}(x_{s-h+1:s+h-1})$. Since $\bm{v}=(v_{1},\dots,v_{T})\sim\widetilde{\mathcal{N}}(0,I_{dT})$, we have $v_{t}\sim\widetilde{\mathcal{N}}(0,I_{d})$ for all $t\in[T]$. Letting $\bm{v}_{s}=v_{s-h+1:s+h-1}\in\mathbb{R}^{2h-1}$ be a column vector and also recalling that $\bar{x}_{t}=x_{t}+\delta^{\prime}v_{t}$ and $\tilde{x}_{t}=x_{t}-\delta^{\prime}v_{t}$, we again have from [28, Eq. (26)] that
$\nabla\mathbb{E}_{\bm{v}_{s}}[F_{s}(\bar{x}_{s-h+1:s+h-1})]$ $=\mathbb{E}_{\bm{v}_{s}}\Big{[}\frac{F(\bar{x}_{s-h+1:s+h-1})-F(\tilde{x}_{s-h+1:s+h-1})}{2\delta^{\prime}}\bm{v}_{s}\Big{]}$
$=\mathbb{E}_{\bm{v}_{s}}\Big{[}\frac{\sum_{k=s}^{s+h-1}\big{(}f_{k}(\bar{x}_{k-h+1:k})-f_{k}(\tilde{x}_{k-h+1:k})}{2\delta^{\prime}}\bm{v}_{s}\Big{]}.$ (A.10)
[…]
Appendix B Proofs of Main Results
B.1 Proof of Theorem 1
Proof.
Based on the definition of Algorithm 1, let us further define an auxiliary function
$h_{t}(x)=\hat{f}_{t}^{c}(x)+(\tilde{g}_{t}-\nabla\hat{f}_{t}^{c}(x_{t}^{0}))x,$
where $\tilde{g}_{t}=\frac{1}{2\delta}\big{[}f_{t}^{c}(\bar{x}_{t}^{0})-f^{c}_{t}(\tilde{x}_{t}^{0})\big{]}u_{t}^{0}$. The above definitions imply $\nabla h_{t}(x_{t}^{0})=\tilde{g}_{t}$. Moreover, we have from Lemma 2(a) that $\mathbb{E}_{u_{t}}[\tilde{g}_{t}]=\nabla\hat{f}_{t}^{c}(x_{t}^{0})$ and thus $\mathbb{E}_{u_{t}}[h_{t}(x)]=\hat{f}_{t}^{c}(x)$ (for all $x$ that is independent of $u_{t}$), which further implies that $\mathbb{E}[h_{t}(x_{t}^{0})]=\mathbb{E}[\hat{f}_{t}^{c}(x_{t}^{0})]$ for all $t$. For our analysis in the sequel, we show that for any $t\in[T]$,
$\left\|\nabla h_{t}(x_{t}^{0})\right\|$ $=\big{\lVert}\frac{1}{2\delta}\big{[}f_{t}^{c}(\bar{x}_{t}^{0})-f_{t}^{c}(\tilde{x}_{t}^{0})\big{]}u^{0}\big{\rVert}$
$\leq\frac{1}{2\delta}|f_{t}^{c}(\bar{x}_{t}^{0})-f_{t}^{c}(\tilde{x}_{t}^{0})|\left\|u^{0}\right\|$
$\overset{(a)}{\leq}\frac{1}{2\delta}G\sqrt{h}\left\|\bar{x}_{t}^{0}-\tilde{x}_{t}^{0}\right\|\left\|u^{0}\right\|$
$=\frac{1}{2\delta}G\sqrt{h}\left\|2\delta u^{0}\right\|\left\|u^{0}\right\|=G\sqrt{h}\left\|u^{0}\right\|^{2},$ (B.1)
where $(a)$ follows from the $G$-Lipschitz of $f_{t}(\cdot)$ in Assumption 1(c) and the definition of $f_{t}^{c}(\cdot)$.
Let us denote $e_{t}=e_{t}^{m}+e_{t}^{c}$ with
$e_{t}^{m}=\frac{1}{2\delta}\Big{[}\big{(}l_{t}(\bar{x}_{t-h+1:t}^{0})-l_{t}(\tilde{x}_{t-h+1:t}^{0})\big{)}-\big{(}f_{t}(\bar{x}_{t-h+1:t}^{0})-f_{t}(\tilde{x}_{t-h+1:t}^{0})\big{)}\Big{]}u^{0},$ (B.2)
$e_{t}^{c}=\frac{1}{2\delta}\Big{[}\big{(}f_{t}(\bar{x}_{t-h+1:t}^{0})-f_{t}(\tilde{x}_{t-h+1:t}^{0})\big{)}-\big{(}f_{t}^{c}(\bar{x}_{t}^{0})-f_{t}^{c}(\tilde{x}_{t}^{0})\big{)}\Big{]}u^{0}.$ (B.3)
The update rule in line 7 of Algorithm 1 can then be viewed as performing online gradient descent on the function sequence $h_{1}(\cdot),h_{2}(\cdot),\dots$ with $x_{t+1}^{0}=\mathbf{\Pi}_{\mathcal{X}}(x_{t}^{0}-\eta_{t}(\nabla h_{t}(x_{t}^{0})+e_{t}))$, where $e_{t}$ is an additive error to the true gradient $\nabla h_{t}(x_{t})=\tilde{g}_{t}$. Based on this observation, we show in Appendix C that the following inequality holds:
$\sum_{t=1}^{T}\big{(}h_{t}(x_{t}^{0})-$ $h_{t}(x_{t}^{\star})\big{)}\leq\sum_{t=1}^{T-1}\frac{D\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|}{\eta_{t}}+\sum_{t=1}^{T}D\left\|e_{t}\right\|+\sum_{t=1}^{T}\eta_{t}\left\|\nabla h_{t}(x_{t}^{0})\right\|^{2}+\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|^{2}.$ (B.4)
Bounding (the expectation of) each term on the right-hand-side of (B.4) and noting that $\mathbb{E}[h_{t}(x_{t}^{0})]=\mathbb{E}[\hat{f}_{t}^{c}(x_{t}^{0})]$ and $\mathbb{E}[h_{t}(x_{t}^{\star})]=\mathbb{E}[\hat{f}_{t}^{c}(x_{t}^{\star})]$ as we argued above, we show that
$\mathbb{E}$ $\Big{[}\sum_{t=1}^{T}\hat{f}_{t}^{c}(x_{t}^{0})-\hat{f}_{t}^{c}(x_{t}^{\star})\Big{]}\leq\frac{\sqrt{2}D}{\mu}\sum_{t=1}^{T-1}\mathbb{E}\Big{[}\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|\Big{]}+(\frac{8G^{2}h^{2}+hG^{2}}{2\mu(2h-1)}+\frac{h^{3}G^{2}D}{\delta(2(2h-1))^{1/2}})(1+\ln T)$
$+\frac{2}{\delta^{2}\mu\sqrt{2h-1}}\sum_{t=1}^{T}\varphi_{t}^{2}+(\frac{D}{\delta(2(2h-1))^{1/4}}+\frac{\sqrt{2}h^{2}GD}{2\delta(2(2h-1))^{1/2}})\sum_{t=1}^{T}\varphi_{t}.$ (B.5)
The proof of (B.5) is also included in Appendix C.
Now, we can relate $\hat{f}_{t}^{c}(\cdot)$ to $f_{t}^{c}(\cdot)$ using Lemma 2. Specifically, we have that for any $t\in[T]$,
$|f_{t}^{c}(x_{t}^{0})-\hat{f}_{t}^{c}(x_{t}^{0})|\leq\frac{\delta^{2}\beta d}{2},\ |f_{t}^{c}(x_{t}^{\star})-\hat{f}_{t}^{c}(x_{t}^{\star})|\leq\frac{\delta^{2}\beta d}{2}.$ (B.6)
To complete the proof of the theorem, we relate $f_{t}^{c}(\cdot)$ to $f_{t}(\cdot)$. From the Lipschitz assumption on $f_{t}(\cdot)$, one can show that
$\quad\mathbb{E}\Big{[}\Big{|}\sum_{t=1}^{T}f_{t}(x_{t-h+1:t}^{0})-f_{t}^{c}(x_{t}^{0})\Big{|}\Big{]}$
$\leq G\sum_{t=1}^{T}\sum_{j=1}^{h-1}\mathbb{E}\big{[}\left\|x_{t}^{0}-x_{t-j}^{0}\right\|\big{]}$
$\leq G\sum_{t=1}^{T}\sum_{j=1}^{h-1}\sum_{l=1}^{j}\mathbb{E}\big{[}\left\|x_{t-l+1}^{0}-x_{t-l}^{0}\right\|\big{]}$
$\overset{(a)}{\leq}G\sum_{t=1}^{T}\sum_{j=1}^{h-1}\sum_{l=1}^{j}\eta_{t-l}\mathbb{E}\big{[}\left\|\nabla h(x_{t-l}^{0})\right\|+\left\|e_{t-l}\right\|\big{]}$
$\overset{(b)}{\leq}\frac{1}{\sqrt{2(2h-1)}}h^{5/2}G^{2}\sum_{t=1}^{T}\eta_{t}+Gh^{2}\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|$
$\overset{(c)}{\leq}\frac{h^{5/2}G^{2}}{\sqrt{2(2h-1)}\mu}(1+\ln T)+\frac{\sqrt{2}Gh^{2}}{\delta\mu(2(2h-1))^{1/4}}\sum_{t=1}^{T}\varphi_{t}+\frac{2G^{2}h^{3}}{\mu(2(2h-1))^{1/2}}(1+\ln T)$
$\leq\frac{3G^{2}h^{3}}{\mu(2(2h-1))^{1/2}}(1+\ln T)+\frac{\sqrt{2}Gh^{2}}{\delta\mu(2(2h-1))^{1/4}}\sum_{t=1}^{T}\varphi_{t},$ (B.7)
where (a) follows from Lemma E.1 in Appendix E, (b) follows from (B.1) and the fact that $\lVert u^{0}\rVert^{2}\leq\sqrt{1/2(2h-1)}$, and (c) uses (C.14) in Appendix C.
Similarly, we have
$\Big{|}\sum_{t=1}^{T}f_{t}(x_{t-h+1:t}^{\star})-f_{t}^{c}(x_{t}^{\star})\Big{|}$ $\leq G\sum_{t=1}^{T}\sum_{j=1}^{h-1}\sum_{l=1}^{j}\left\|x_{t-l+1}^{\star}-x_{t-l}^{\star}\right\|$
$\leq Gh^{2}\sum_{t=1}^{T}\left\|x_{t}^{\star}-x_{t-1}^{\star}\right\|$
=$\overset{(a)}{=}Gh^{2}\sum_{t=1}^{T-1}\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|,$ (B.8)
where (a) follows by setting $x_{0}^{\star}=x_{1}^{\star}$ since $t=0$ is vacuous in our analysis. The final result of the theorem then follows by combining (B.5)-(B.8) and noting that $\delta=1/\sqrt{T}\leq 1$. ∎
B.2 Proof of Theorem 2
Proof.
For our analysis in this proof, denote
$\widetilde{\bm{g}}^{j}_{s}=\sum_{k=s}^{s+h-1}\frac{f_{k}(\bar{x}^{j}_{k-h+1:k})-f_{k}(\tilde{x}^{j}_{k-h+1:k})}{2\delta^{\prime}}u_{s}^{j}.$ (B.9)
We then have from Lemma 2 that $\mathbb{E}_{u_{s}^{j}}[\widetilde{\bm{g}}_{s}^{j}]=\partial\hat{C}_{T}(\bm{x}^{j})/\partial x_{s}$ for all $s\in[T]$ and all $j\in\{0,\dots,K-1\}$ and thus $\nabla\hat{C}_{T}(\bm{x}^{j})=\mathbb{E}_{\bm{u}^{j}}[\widetilde{\bm{g}}^{j}]$, where $\bm{u}^{j}=(u_{1}^{j},\dots,u_{T}^{j})\sim\widetilde{\mathcal{N}}(0,I_{dT})$ and $\widetilde{\bm{g}}^{j}=(\widetilde{\bm{g}}_{1}^{j},\dots,\widetilde{\bm{g}}_{T}^{j})$.
To proceed, we define the following potential function based on $C_{T}(\cdot)$:
$V_{j}=(1+\gamma)^{j}(C_{T}(\bm{x}^{j})-C_{T}(\bm{x}^{\star})).$ (B.10)
We then have
$V_{j+1}-V_{j}$ $=(1+\gamma)^{j}\big{(}(\gamma+1)(C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j}))+\gamma(C_{T}(\bm{x}^{j})-C_{T}(\bm{x}^{\star}))\big{)}$
$=(1+\gamma)^{j}\big{(}C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j})+\gamma(C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{\star}))\big{)}.$ (B.11)
To upper bound $V_{j+1}-V_{j}$, we first prove in Appendix D that the following inequality holds for all $j\in\{0,\dots,K-1\}$ and all $\bm{y}\in\mathbb{R}^{dT}$:
$C_{T}(\bm{x}^{j+1})-C_{T}(\bm{y})\leq\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j}-\bm{y}\rangle-\frac{\beta^{\prime}}{2}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}-\frac{\mu}{2}\left\|\bm{x}^{j}-\bm{y}\right\|^{2}\\ +\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle.$ (B.12)
Letting $\bm{y}=\bm{x}^{j}$ and $\bm{y}=\bm{x}^{\star}$ in (B.12), we respectively obtain the following:
$C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j})$ $\leq-\frac{\beta^{\prime}}{2}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{x}^{j}\rangle,$ (B.13)
$C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{\star})$ $\leq\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j}-\bm{x}^{\star}\rangle-\frac{\beta^{\prime}}{2}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}-\frac{\mu}{2}\left\|\bm{x}^{j}-\bm{x}^{\star}\right\|^{2}$
$\qquad\qquad\qquad\qquad\qquad\qquad+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{x}^{\star}\rangle.$ (B.14)
Moreover, for any $\bm{y}=(y_{1},\dots,y_{T})$ with $y_{t}\in\mathcal{X}$ for all $t\in[T]$, we have the following relation proved in Appendix D:
$\mathbb{E}\big{[}\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big{]}\leq\varepsilon.$ (B.15)
Now, substituting (B.13)-(B.14) with the upper bound $\varepsilon$ into (B.11), we show in Appendix D that
$\mathbb{E}_{\bm{u}^{j}}[V_{j+1}-V_{j}]\leq(1+\gamma)^{j+1}\varepsilon.$ (B.16)
From (B.10) and (B.16), we get
$\mathbb{E}_{\bm{u}^{j}}\big{[}C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{\star})\big{]}\leq\frac{1}{1+\gamma}\big{(}C_{T}(\bm{x}^{j})-C_{T}(\bm{x}^{\star})\big{)}+\varepsilon\quad\forall j\in\{0,\dots,K-1\},$
$\Rightarrow$ $\mathbb{E}_{\mathcal{U}_{K-1}}\big{[}C_{T}(\bm{x}^{K})-C_{T}(\bm{x}^{\star})\big{]}\leq\Big{(}\frac{1}{1+\gamma}\Big{)}^{K}\big{(}C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})\big{)}+\sum_{j=0}^{K-1}\Big{(}\frac{1}{1+\gamma}\Big{)}^{j}\varepsilon$
$\Rightarrow$ $\mathbb{E}_{\mathcal{U}_{K-1}}\big{[}C_{T}(\bm{x}^{K})-C_{T}(\bm{x}^{\star})\big{]}\leq\Big{(}\frac{1}{1+\gamma}\Big{)}^{K}\big{(}C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})\big{)}+\frac{\varepsilon}{\gamma},$
where $\mathcal{U}_{K-1}\triangleq(\bm{u}^{0},\dots,\bm{u}^{K-1})$ encapsulates all the randomness of Algorithm 1, completing the proof of the theorem. ∎
Appendix C Omitted Details for the Proof of Theorem 1
C.1 Proof of (B.4)
Recall that $(x_{1}^{\star},\dots,x_{T}^{\star})$ is an optimal solution to (2). Since $\hat{f}_{t}^{c}(\cdot)$ is $\mu$-strongly convex from Lemma 1, we know by the definition that $h_{t}(\cdot)$ is also $\mu$-strongly convex, i.e.,
$h_{t}(x_{t}^{0})-h_{t}(x_{t}^{\star})\leq\nabla h_{t}(x_{t}^{0})^{\top}(x_{t}^{0}-x_{t}^{\star})-\frac{\mu}{2}\left\|x_{t}^{0}-x_{t}^{\star}\right\|^{2}.$ (C.1)
We also have
$\left\|x_{t}^{\star}-x_{t+1}^{0}\right\|^{2}$ $\leq\left\|x_{t}^{\star}-\mathbf{\Pi}_{\mathcal{X}}(x_{t}^{0}-\eta_{t}(\nabla h_{t}(x_{t}^{0})+e_{t}))\right\|^{2}$
$\overset{(a)}{\leq}\left\|x_{t}^{\star}-x_{t}^{0}+\eta_{t}(\nabla h_{t}(x_{t}^{0})+e_{t})\right\|^{2}$
$=\left\|x_{t}^{\star}-x_{t}^{0}\right\|^{2}+2(x_{t}^{\star}-x_{t}^{0})^{\top}\eta_{t}(\nabla h_{t}(x_{t}^{0})+e_{t})+\eta_{t}^{2}\left\|\nabla h_{t}(x_{t}^{0})+e_{t}\right\|^{2},$
where (a) follows from Lemma E.1 in Appendix E. It follows that
$(x_{t}^{0}-x_{t}^{\star})^{\top}\nabla h_{t}(x_{t}^{0})\leq\frac{\left\|x_{t}^{\star}-x_{t}^{0}\right\|^{2}-\left\|x_{t}^{\star}-x_{t+1}^{0}\right\|^{2}}{2\eta_{t}}+(x_{t}^{\star}-x_{t}^{0})^{\top}e_{t}+\frac{\eta_{t}}{2}\left\|\nabla h_{t}(x_{t}^{0})+e_{t}\right\|^{2}.$ (C.2)
Combining (C.1) and (C.2) yields
$h_{t}(x_{t}^{0})-h_{t}(x_{t}^{\star})$ $\leq\frac{\left\|x_{t}^{\star}-x_{t}^{0}\right\|^{2}-\left\|x_{t}^{\star}-x_{t+1}^{0}\right\|^{2}}{2\eta_{t}}+(x_{t}^{\star}-x_{t}^{0})^{\top}e_{t}$
$\qquad+\eta_{t}\left\|\nabla h_{t}(x_{t}^{0})\right\|^{2}+\eta_{t}\left\|e_{t}\right\|^{2}-\frac{\mu}{2}\left\|x_{t}^{0}-x_{t}^{\star}\right\|,$ (C.3)
where we used $\left\|a+b\right\|^{2}\leq 2\left\|a\right\|^{2}+2\left\|b\right\|^{2}$ for $a,b\in\mathbb{R}^{d}$. Meanwhile,
$\left\|x_{t}^{\star}-x_{t+1}^{0}\right\|$ $=\left\|x_{t}^{\star}-x_{t+1}^{\star}+x_{t+1}^{\star}-x_{t+1}^{0}\right\|^{2}$
$=\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|^{2}+\left\|x_{t+1}^{\star}-x_{t+1}^{0}\right\|^{2}+2(x_{t}^{\star}-x_{t+1}^{\star})^{\top}(x_{t+1}^{\star}-x_{t+1}^{0})$
$\geq\left\|x_{t+1}^{\star}-x_{t+1}^{0}\right\|^{2}-2D\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|,$ (C.4)
where the inequality follows from Assumption 2 and the fact that $a^{\top}b\geq-\left\|a\right\|\left\|b\right\|$ for $a,b\in\mathbb{R}^{d}$. We then have from (C.3) and (C.4) that
$h_{t}(x_{t}^{0})-h_{t}(x_{t}^{\star})$ $\leq\frac{\left\|x_{t}^{\star}-x_{t}^{0}\right\|^{2}-\left\|x_{t+1}^{\star}-x_{t+1}^{0}\right\|^{2}}{2\eta_{t}}+\frac{D\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|}{\eta_{t}}$
$\qquad+D\left\|e_{t}\right\|+\eta_{t}\left\|\nabla h_{t}(y_{t})\right\|^{2}+\eta_{t}\left\|e_{t}\right\|^{2}-\frac{\mu}{2}\left\|x_{t}^{0}-x_{t}^{\star}\right\|,$ (C.5)
where we used $(x_{t}^{\star}-x_{t}^{0})^{\top}e_{t}\leq\left\|x_{t}^{\star}-x_{t}^{0}\right\|\left\|e_{t}\right\|\leq D\left\|e_{t}\right\|$. Recalling $\eta_{t}=\frac{1}{t\mu}$, we have
$\sum_{t=1}^{T}\Big{(}\frac{\left\|x_{t}^{\star}-x_{t}^{0}\right\|^{2}-\left\|x_{t+1}^{\star}-x_{t+1}^{0}\right\|^{2}}{2\eta_{t}}-\frac{\mu}{2}\left\|x_{t}^{0}-x_{t}^{\star}\right\|\Big{)}$
$=$ $\sum_{t=1}^{T}(\frac{1}{2\eta_{t}}-\frac{1}{2\eta_{t-1}}-\frac{\mu}{2})\left\|x_{t}^{\star}-x_{t}^{0}\right\|=0,$
[…]
C.2 Proof of (B.5)
First, note that
$\sum_{t=1}^{T-1}\frac{D\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|}{\eta_{t}}\overset{(a)}{\leq}\frac{D}{\mu}\sqrt{(\sum_{t=1}^{T-1}\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|^{2})(\sum_{t=1}^{T-1}\frac{1}{t^{2}})}$
$\overset{(b)}{\leq}$ $\frac{\sqrt{2}D}{\mu}\sqrt{\sum_{t=1}^{T-1}\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|^{2}}$
$\leq$ $\frac{\sqrt{2}D}{\mu}\sum_{t=1}^{T-1}\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|,$ (C.6)
where (a) follows from the CS-inequality and (b) follows from the fact
$\sum_{t=1}^{T-1}\frac{1}{t^{2}}\leq 1+\sum_{t=2}^{T-1}\frac{1}{t(t-1)}=1+\sum_{t=2}^{T-1}(\frac{1}{t-1}-\frac{1}{t})\leq 2.$ (C.7)
Next, we have from (B.1) that
$\sum_{t=1}^{T}\eta_{t}\left\|\nabla h_{t}(x_{t}^{0})\right\|^{2}$ $\leq\frac{hG^{2}}{\mu}\sum_{t=1}^{T}\frac{1}{t}\left\|u_{t}^{0}\right\|^{4}$
$\overset{(a)}{\leq}\frac{hG^{2}}{2(2h-1)\mu}\sum_{t=1}^{T}\frac{1}{t}$
$\overset{(b)}{\leq}\frac{hG^{2}}{2(2h-1)\mu}(1+\ln T),$ (C.8)
where (a) follows from the fact that $\lVert u_{t}^{0}\rVert\leq 1/(2(2h-1))^{1/4}$, and (b) uses the fact $\sum_{t=1}^{T}\frac{1}{t}<1+\ln T$.
To proceed, we upper bound $\left\|e_{t}\right\|$ for all $t\in[T]$, where $e_{t}=e_{t}^{m}+e_{t}^{c}$. We first have from that (B.2)
$\left\|e_{t}^{m}\right\|\leq\frac{\left\|u^{0}\right\|}{\delta}\varphi_{t}\leq\frac{\varphi_{t}}{\delta(2(2h-1))^{1/4}}.$ (C.9)
To upper bound $\left\|e_{t}^{c}\right\|$, we have from (B.3) that
$\left\|e_{t}^{c}\right\|\leq\frac{1}{2\delta}|f_{t}(\bar{x}_{t-h+1:t}^{0})-f_{t}^{c}(\bar{x}_{t}^{0})|\left\|u^{0}\right\|+\frac{1}{2\delta}|f_{t}(\tilde{x}_{t-h+1:t}^{0})-f_{t}^{c}(\tilde{x}_{t}^{0})|\left\|u^{0}\right\|.$ (C.10)
We may further upper bound bound $\left\|e_{t}^{c}\right\|$ as
$\left\|e_{t}^{c}\right\|$ $\leq\frac{1}{2\delta}|f_{t}(\bar{x}_{t-h+1:t}^{0})-f_{t}(\tilde{x}_{t-h+1:t}^{0})|\left\|u^{0}\right\|+\frac{1}{2\delta}|f_{t}^{c}(\bar{x}_{t}^{0})-f_{t}^{c}(\tilde{x}_{t}^{0})|\left\|u^{0}\right\|$
$\overset{(a)}{\leq}\frac{G}{2\delta}\sum_{j=0}^{h-1}\left\|\bar{x}_{t-j}^{0}-\tilde{x}_{t-j}^{0}\right\|\left\|u^{0}\right\|+\frac{Gh}{2\delta}\left\|\bar{x}_{t}^{0}-\tilde{x}_{t}^{0}\right\|\left\|u^{0}\right\|$
$\overset{(b)}{\leq}G\sum_{j=0}^{h-1}\left\|u^{0}\right\|\left\|u^{0}\right\|+Gh\left\|u^{0}\right\|^{2}$
$\leq\frac{2Gh}{\sqrt{2(2h-1)}},$ (C.11)
where (a) again follows from Assumption 1(c), and (b) uses the relation between $\bar{x}_{t}^{0}$ and $\tilde{x}_{t}^{0}$. The above inequality further implies
$\left\|e_{t}^{c}\right\|^{2}\leq\frac{2G^{2}h^{2}}{(2h-1)}.$ (C.12)
We then upper bound $\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|^{2}$ in (B.4). We have
$\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|^{2}$ $\leq\sum_{t=1}^{T}\eta_{t}(2\left\|e_{t}^{m}\right\|^{2}+2\left\|e_{t}^{c}\right\|^{2})$
$\overset{(a)}{\leq}\sum_{t=1}^{T}\eta_{t}\frac{2\varphi_{t}^{2}}{\delta^{2}\sqrt{2(2h-1)}}+4\sum_{t=1}^{T}\eta_{t}\frac{G^{2}h^{2}}{(2h-1)}$
$\overset{(b)}{\leq}\frac{2}{\delta^{2}\mu\sqrt{2(2h-1)}}\sqrt{(\sum_{t=1}^{T}\varphi_{t}^{4})(\sum_{t=1}^{T}\frac{1}{t^{2}})}+\frac{4G^{2}h^{2}}{\mu(2h-1)}\sum_{t=1}^{T}\frac{1}{t}$
$\overset{(c)}{\leq}\frac{2}{\delta^{2}\mu\sqrt{2h-1}}\sum_{t=1}^{T}\varphi_{t}^{2}+\frac{4G^{2}h^{2}}{\mu(2h-1)}(1+\ln T),$ (C.13)
where (a) follows from (C.9) and (C.12), (b) follows from the choice $\eta=\frac{1}{t\mu}$ and the CS-inequality. To obtain (c), we used the fact $\sum_{i=1}^{n}a_{i}^{2}\leq(\sum_{i=1}^{n}a_{i})^{2}$ for $a_{i}\in\mathbb{R}^{+}$, $\sum_{t=1}^{T}1/t^{2}\leq 2$ from (C.7) and $\sum_{t=1}^{T}\frac{1}{t}\leq 1+\ln T$. Similarly, we can upper bound $\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|$ as
$\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|$ $\leq\sum_{t=1}^{T}\eta_{t}(\left\|e_{t}^{m}\right\|+\left\|e_{t}^{c}\right\|)$
$\leq\sum_{t=1}^{T}\frac{1}{t\mu}\frac{\varphi_{t}}{\delta(2(2h-1))^{\frac{1}{4}}}+\sum_{t=1}^{T}\frac{1}{t\mu}\frac{\sqrt{2}Gh}{\sqrt[]{2h-1}}$
$\leq\frac{\sqrt{2}}{\delta\mu(2(2h-1))^{1/4}}\sum_{t=1}^{T}\varphi_{t}+\frac{\sqrt{2}Gh}{\mu\sqrt{2h-1}}(1+\ln T).$ (C.14)
Finally, we upper bound $\sum_{t=1}^{T}D\left\|e_{t}\right\|$ in (B.4). We first note from Assumption 1(c) that
$\quad|f_{t}(\bar{x}_{t-h+1:t}^{0})-f_{t}^{c}(\bar{x}_{t}^{0})|$
$\leq G\sum_{j=1}^{h-1}\left\|\bar{x}_{t}^{0}-\bar{x}_{t-j}^{0}\right\|$
$\leq G\sum_{j=1}^{h-1}\sum_{l=1}^{j}\left\|\bar{x}_{t-l+1}^{0}-\bar{x}_{t-l}^{0}\right\|$
$\leq G\sum_{j=1}^{h-1}\sum_{l=1}^{j}\left\|x_{t-l+1}^{0}+\delta u^{0}-x_{t-l}^{0}-\delta u^{0}\right\|$
$\overset{(a)}{\leq}G\sum_{j=1}^{h-1}\sum_{l=1}^{j}\big{(}\eta_{t-l}\left\|\nabla h_{t-l}(x_{t-l}^{0})\right\|+\eta_{t-l}\left\|e_{t-l}\right\|\big{)}$
[…]
where (a) uses the relation $x_{t-l+1}^{0}=\Pi_{\mathcal{X}}(x_{t-l}^{0}-\eta_{t-l}(\nabla h_{t-l}(x_{t-l}^{0})+e_{t}))$ and Lemma E.1 in Appendix E.
Similarly, one can show that
$|f_{t}(\tilde{x}_{t-h+1:t}^{0})-f_{t}^{c}(\tilde{x}_{t}^{0})|\leq\sqrt{h}G^{2}\sum_{j=1}^{h-1}\sum_{l=1}^{j}\eta_{t-l}\left\|u^{0}\right\|^{2}+G\sum_{j=1}^{h-1}\sum_{l=1}^{j}\eta_{t-l}\left\|e_{t-l}\right\|.$
We then have from (C.10) that
$\left\|e^{c}_{t}\right\|\leq\frac{\sqrt{h}G^{2}}{\delta}\sum_{j=1}^{h-1}\sum_{l=1}^{j}\eta_{t-l}\left\|u^{0}\right\|^{3}+\frac{G}{\delta}\sum_{j=1}^{h-1}\sum_{l=1}^{j}\left\|e_{t-l}\right\|\left\|u^{0}\right\|$
$\quad\quad\leq\frac{\sqrt{h}h^{2}G^{2}}{2\delta}\sum_{t=1}^{T}\eta_{t}\left\|u^{0}\right\|^{3}+\frac{h^{2}G^{2}}{2\delta}\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|\left\|u^{0}\right\|,$
[…]
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{\sqrt{2}Gh}{\mu\sqrt{2h-1}}(1+\ln T)\big{]}$
$\quad\quad\quad\leq\frac{\sqrt{2}h^{2}G}{2\delta(2(2h-1))^{1/2}}\sum_{t=1}^{T}\varphi_{t}+\frac{h^{3}G^{2}}{\delta(2(2h-1))^{1/2}}(1+\ln{T}).$ (C.15)
Combining (C.9) and (C.15) and summing from $t=1$ to $t=T$, we get
$\quad\mathbb{E}\Big{[}\sum_{t=1}^{T}D\left\|e_{t}\right\|\Big{]}$
$\leq(\frac{D}{\delta(2(2h-1))^{1/4}}+\frac{\sqrt{2}h^{2}GD}{2\delta(2(2h-1))^{1/2}})\sum_{t=1}^{T}\varphi_{t}+\frac{h^{3}G^{2}D}{\delta(2(2h-1))^{1/2}}(1+\ln{T}).$ (C.16)
Combining the upper bounds in (C.6), (C.8), (C.13) and (C.16), noting that $\delta=1/\sqrt{T}\leq 1$ and recalling that $\mathbb{E}[h_{t}(x_{t}^{0})]=\mathbb{E}[\hat{f}_{t}^{c}(x_{t}^{0})]$ and $\mathbb{E}[h_{t}(x_{t}^{\star})]=\mathbb{E}[\hat{f}_{t}^{c}(x_{t}^{\star})]$, we obtain (B.5) from (B.4).
Appendix D Omitted Details in the Proof of Theorem 2
D.1 Proof of (B.12)
We first express $C_{T}(\bm{x}^{j+1})-C_{T}(\bm{y})=C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j})+C_{T}(\bm{x}^{j})-C_{T}(\bm{y})$. From $\beta^{\prime}$-smoothness of $C_{T}(\cdot)$, we have
$C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j})\leq\langle\nabla C_{T}(\bm{x}^{j}),\bm{x}^{j+1}-\bm{x}^{j}\rangle+\frac{\beta^{\prime}}{2}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}.$
From the $\mu$-strongly convexity of $C_{T}(\cdot)$, we have
$C_{T}(\bm{x}^{j})-C_{T}(\bm{y})\leq\langle\nabla C_{T}(\bm{x}^{j}),\bm{x}^{j}-\bm{y}\rangle-\frac{\mu}{2}\left\|\bm{x}^{j}-\bm{y}\right\|^{2}.$
Summing the above two inequalities yields
$C_{T}(\bm{x}^{j+1})-C_{T}(\bm{y})\leq\langle\nabla C_{T}(\bm{x}^{j}),\bm{x}^{j+1}-\bm{y}\rangle+\frac{\beta^{\prime}}{2}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}-\frac{\mu}{2}\left\|\bm{x}^{j}-\bm{y}\right\|^{2}.$ (D.1)
As we argued in the main body, denoting $\bm{z}^{j+1}=\bm{x}^{j}-\alpha\bm{g}^{j}$ with $\alpha=1/\beta^{\prime}$, we get that
$\langle\nabla C_{T}(\bm{x}^{j}),\bm{x}^{j+1}-\bm{y}\rangle$ $=\langle\bm{g}^{j}+\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle$
$=\langle\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle$
$=\beta^{\prime}\langle\bm{x}^{j}-\bm{z}^{j+1},\bm{x}^{j+1}-\bm{y}\rangle+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle$
$\leq\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j+1}-\bm{y}\rangle+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle$
$=\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j+1}-\bm{x}^{j}+\bm{x}^{j}-\bm{y}\rangle+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle$
$=-{\beta^{\prime}}^{2}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}+\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j}-\bm{y}\rangle+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle,$
[…]
D.2 Proof of (B.15)
For any $\bm{y}=(y_{1},\dots,y_{T})$ with $y_{t}\in\mathcal{X}$ for all $t\in[T]$, we have
$\mathbb{E}_{\bm{u}^{j}}\big{[}\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big{]}$ $=\mathbb{E}_{\bm{u}^{j}}\big{[}\langle\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}+\widetilde{\bm{g}}^{j}-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big{]},$
$=\underbrace{\mathbb{E}_{\bm{u}^{j}}\big{[}\langle\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big{]}}_{\text{(i)}}+\underbrace{\mathbb{E}_{\bm{u}^{j}}\big{[}\langle\widetilde{\bm{g}}^{j}-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big{]}}_{\text{(ii)}}.$
We first upper bound (ii). Recalling the definition of $\bm{g}^{j}$ (resp., $\widetilde{\bm{g}}^{j}$) given in Algorithm 1 (resp., (B.9)), we get that
$\mathbb{E}_{{u}^{j}_{s}}[\bm{g}^{j}_{s}-\widetilde{\bm{g}}^{j}_{s}]=\mathbb{E}_{u^{j}_{s}}\Big{[}\sum_{k=s}^{s+h-1}\frac{l_{k}(\bar{x}_{k-h+1:k})-f_{k}(\bar{x}_{k-h+1:k})+f_{k}(\tilde{x}_{k-h+1:k})-l_{k}(\tilde{x}_{k-h+1:k})}{2\delta^{\prime}}u^{j}_{s}\Big{]},$
for all $s\in[T]$. By Assumption 3 and the fact that $\|u_{s}^{j}\|\leq 1/2(2h-1))^{1/4}$, we obtain
$\mathbb{E}_{\bm{u}^{j}}\big{[}\lVert\bm{g}_{s}^{j}-\widetilde{\bm{g}}_{s}^{j}\rVert\big{]}\leq\frac{\sum_{k=s}^{s+h-1}\varphi_{k}}{\delta^{\prime}(2(2h-1))^{1/4}}\ \forall s\in[T],$
$\Rightarrow$ $\mathbb{E}_{\bm{u}^{j}}\big{[}\lVert\bm{g}^{j}-\widetilde{\bm{g}}^{j}\rVert\big{]}\leq\frac{\sum_{s=1}^{T}\sum_{k=s}^{s+h-1}\varphi_{k}}{\delta^{\prime}(2(2h-1))^{1/4}}\leq\frac{h\sum_{t=1}^{T}\varphi_{t}}{\delta^{\prime}(2(2h-1))^{1/4}}.$
Thus, we can upper bound (ii) as
[…]
where the second inequality follows from Assumption 2.
We then upper bound (i). Note that
$\mathbb{E}_{\bm{u}^{j}}\big{[}\langle\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big{]}$
$\leq\mathbb{E}_{\bm{u}^{j}}\big{[}\|\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}\|\|\bm{x}^{j+1}-\bm{y}\|\big{]}$
$\overset{(a)}{\leq}D\mathbb{E}_{\bm{u}^{j}}\big{[}\|\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}\|\big{]}$
$\leq D\mathbb{E}_{\bm{u}^{j}}\Big{[}\sqrt{\|\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}\|^{2}}\Big{]}$
$\overset{(b)}{\leq}D\sqrt{\mathbb{E}_{\bm{u}^{j}}\big{[}\|\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}\|^{2}\big{]}}$
$\Rightarrow$ $\frac{1}{D^{2}}\big{(}\mathbb{E}_{\bm{u}^{j}}\big{[}\langle\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big{]}\big{)}^{2}$
$\leq\mathbb{E}_{\bm{u}^{j}}\Big{[}\left\|\nabla C_{T}(\bm{x}^{j})\right\|^{2}+\left\|\widetilde{\bm{g}}^{j}\right\|^{2}-2\nabla C_{T}(\bm{x}^{j})^{\top}\widetilde{\bm{g}}^{j}\Big{]}$
$=\left\|\nabla C_{T}(\bm{x}^{j})\right\|^{2}-2\nabla C_{T}(\bm{x}^{j})^{\top}\nabla\hat{C}_{T}(\bm{x}^{j})+\mathbb{E}_{\bm{u}^{j}}\big{[}\left\|\widetilde{\bm{g}}^{j}\right\|^{2}\big{]}$
$=\nabla C_{T}(\bm{x}^{j})^{\top}\big{(}-\nabla C_{T}(\bm{x}^{j})+2\nabla C_{T}(\bm{x}^{j})-2\nabla\hat{C}_{T}(\bm{x}^{j})\big{)}+\mathbb{E}_{\bm{u}^{j}}\big{[}\left\|\widetilde{\bm{g}}^{j}\right\|^{2}\big{]}$
$=-\left\|\nabla C_{T}(\bm{x}^{j})\right\|^{2}+2\left\|\nabla C_{T}(\bm{x}^{j})\right\|\left\|\nabla C_{T}(\bm{x}^{j})-\nabla\hat{C}_{T}(\bm{x}^{j})\right\|+\mathbb{E}_{\bm{u}^{j}}\big{[}\left\|\widetilde{\bm{g}}^{j}\right\|^{2}\big{]}$
$\overset{(c)}{\leq}-\left\|\nabla C_{T}(\bm{x}^{j})\right\|^{2}+\mathbb{E}_{\bm{u}^{j}}\big{[}\left\|\widetilde{\bm{g}}^{j}\right\|^{2}\big{]}+G\beta hT(Td+3)^{3/2}\delta^{\prime},$
where (a) follows from Assumption 2, (b) follows from Jensen’s inequality and the concavity of $\sqrt{\cdot}$ and (c) follows from Lemma 1(b) and Lemma 2(b). [^6] Note the well-known fact that if a differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ is convex and $G$-Lipschitz, then $\|\nabla f(x)\|\leq G$ for all $x\in\mathbb{R}^{d}$. Thus, we know from Lemma 1(b) that $\|\nabla C_{T}(\bm{x})\|\leq G\sqrt{Th}$ for all $\bm{x}\in\mathbb{R}^{dT}$. It remains to upper bound the term $-\|C_{T}(\bm{x}^{j})\|^{2}+\mathbb{E}_{\bm{u}^{j}}\big{[}\|\widetilde{\bm{g}}^{j}\|^{2}\big{]}$. Noting that $\sum_{k=s}^{s+h-1}f_{k}(x_{k-h+1:k})$ may be viewed as a function of $x_{s-h+1},\dots,x_{s+h-1}$, denoted as $F_{s}(x_{s-h+1:s+h-1})$, we see from (B.9) that $\widetilde{\bm{g}}^{j}_{s}$ can be equivalently written as
$\widetilde{\bm{g}}^{j}_{s}=\frac{1}{2\delta^{\prime}}\big{(}F_{s}(\bar{x}^{j}_{s-h+1:s+h-1})-F_{s}(\tilde{x}^{j}_{s-h+1:s+h-1})\big{)}u_{s}^{j}.$
Moreover, using similar arguments to those in the proof of Lemma 1, one can show that $F_{s}(\cdot)$ is $\beta h$-smooth. Now, following similar arguments to those in the proof of [28, Theorem 4], one can show that
$\left\|\widetilde{\bm{g}}^{j}_{s}\right\|^{2}$ $\leq\frac{\delta^{\prime 2}}{8}\beta^{2}h^{2}\left\|u_{s}^{j}\right\|^{6}+2\left\|u_{s}^{j}\right\|^{4}\left\|\nabla F_{s}(x^{j}_{s-h+1:s+h-1})\right\|^{2}$
$\leq\frac{\delta^{\prime 2}\beta^{2}h^{2}}{8(2(2h-1))^{3/2}}+\frac{1}{(2h-1)}\left\|\nabla F_{s}(x^{j}_{s-h+1:s+h-1})\right\|^{2},$
where the second inequality follows from the fact that $\|u_{s}^{j}\|\leq 1/(2(2h-1))^{1/4}$. One can also show that $\sum_{s=1}^{T}\|\nabla F_{s}(x^{j}_{s-h+1:s+h-1})\|^{2}\leq(2h-1)\|\nabla C_{T}(\bm{x}^{j})\|^{2}$. It follows that
$\left\|\widetilde{\bm{g}}^{j}\right\|^{2}=\sum_{s=1}^{T}\left\|\widetilde{\bm{g}}^{j}_{s}\right\|^{2}\leq\frac{\delta^{\prime 2}\beta^{2}h^{2}T}{8(2(2h-1))^{3/2}}+\left\|\nabla C_{T}(\bm{x}^{j})\right\|^{2}.$
[…]
$\leq\frac{D\delta^{\prime}\beta h\sqrt{T}}{2\sqrt{2}(2(2h-1))^{3/4}}+\sqrt{h\beta GT\delta^{\prime}}D(Td+3)^{3/4}+\frac{Dh\sum_{t=1}^{T}\varphi_{t}}{\delta^{\prime}(2(2h-1))^{1/4}}\triangleq\varepsilon.$
[…]
D.3 Proof of (B.16)
We get from (B.13)-(B.14) and the upper bound $\varepsilon$ in (B.15) that
$\quad C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j})+\gamma(C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{\star}))$
$\leq-\frac{\beta^{\prime}}{2}(1+\gamma)\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|+\gamma\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j}-\bm{x}^{\star}\rangle-\frac{\mu\gamma}{2}\left\|\bm{x}^{j}-\bm{x}^{\star}\right\|^{2}+(1+\gamma)\varepsilon$
$\leq-\frac{1}{\beta^{\prime}-\mu}\left\|\beta^{\prime}(\bm{x}^{j+1}-\bm{x}^{j})-\mu(\bm{x}^{j}-\bm{x}^{\star})\right\|^{2}+(1+\gamma)\varepsilon,$
[…]
Appendix E Technical Lemma
Lemma E.1.
[42, Proposition 2.2] Given a convex set $\mathcal{X}\subseteq\mathbb{R}^{d}$, let $a\in\mathcal{X}$, $b^{\prime}\in\mathbb{R}^{d}$ and $b=\mathbf{\Pi}_{\mathcal{X}}(b^{\prime})$. Then $\langle a-b,b^{\prime}-b\rangle\leq 0$ and $\left\|a-b\right\|^{2}\leq\left\|a-b^{\prime}\right\|^{2}$.