### v3

# Online Convex Optimization with Memory and Limited Predictions

Abstract
This paper addresses an online convex optimization problem where the cost function at each step depends on a history of past decisions (i.e., memory), and the decision maker has access to limited predictions of future cost values within a finite window. The goal is to design an algorithm that minimizes the dynamic regret against the optimal sequence of decisions in hindsight. To this end, we propose a novel predictive algorithm and establish strong theoretical guarantees for its performance. We show that the algorithm’s dynamic regret decays exponentially with the length of the prediction window. Our algorithm comprises two general subroutines of independent interest. The first subroutine solves online convex optimization with memory and bandit feedback, achieving a $\sqrt{TV_{T}}$-dynamic regret, where $V_{T}$ measures the variation of the optimal decision sequence. The second is a zeroth-order method that attains a linear convergence rate for general convex optimization, matching the best achievable rate of first-order methods. The key to our algorithm is a novel truncated Gaussian smoothing technique when querying the decision points to obtain the predictions. We validate our theoretical results with numerical experiments.
1 Introduction
Online decision making is a ubiquitous framework that captures a wide range of real-world applications, such as machine learning, control systems, and finance cesa2006prediction. In this framework, a decision maker interacts with an unknown environment by sequentially choosing a decision point $x_{t}$ at each time step and incurs a cost $f_{t}(x_{t})$ given by some underlying cost function $f_{t}(\cdot)$, and the goal is to minimize the accumulative cost over a horizon $t=1,\dots,T$. For convex cost functions $f_{t}(\cdot)$, the problem is known as Online Convex Optimization (OCO), which has been widely studied in the past decades (see, e.g., hazan2016introduction; shalev2012online for a comprehensive survey). The OCO framework may be further split into two settings. In the full information setting, the decision maker receives complete knowledge of the cost function $f_{t}(\cdot)$ after choosing $x_{t}$ for time step $t$ zinkevich2003online; in the bandit information setting, the decision maker only receives $f_{t}(x_{t})$ (and potentially $f_{t}(\bar{x}_{t})$ at some perturbed points $\bar{x}_{t}$ from $x_{t}$) after choosing $x_{t}$ for time step $t$ flaxman2005online; agarwal2010optimal; saha2011improved. Since the decision maker needs to choose the decision points in an online manner (i.e., before knowing all the cost functions), regret is a typical performance metric used for algorithms proposed to solve OCO, which compares the accumulative cost incurred by the algorithm to the minimum achievable cost in hindsight zinkevich2003online. For the non-stationary environment zhao2022non, the minimum achievable cost typically requires time-varying decisions over the horizon $T$, leading to dynamic regret yang2016tracking; jadbabaie2015online; zhao2021bandit.
In many online decision-making problems, the cost incurred at the current time step depends not only on the decision made for the current time step but also on the decisions made in the past. For example, when the underlying environment is characterized by a dynamical system whose state evolves according to the chosen actions, the cost incurred at the current step is governed by the current action and system state, which in turn depends on past actions and system states. Such a framework of OCO (i.e., OCO with memory) has been studied in arora2012online; anava2015online; zhou2023efficient and specialized to the online control problem of linear dynamical systems agarwal2019online; simchowitz2020improper; yan2023online; ye2024online; ye2024learning. Due to the long-term dependency of the cost on the decision history, i.e., the current decision also influences cost at future steps, there is a line of research that also considers predictions of the cost in a window of future time steps lin2012online; chen2015online; chen2016using; li2020online; li2019control. We refer to this framework as OCO with memory and predictions. However, the works mentioned above have certain limitations. In chen2015online; chen2016using, the cost is assumed to be only a function of decisions in the current step and one step in the past. The classical control algorithm, model predictive control (MPC) rawlings2012postface; zeilinger2011real; angeli2016theoretical, considers the receding horizon optimization for the next $W$ stages. However, similar to li2019control; li2020online, the predictions of the full information of the cost functions are assumed to be available in the MPC algorithms.
Motivated by these limitations in the existing works, we study the problem of OCO with memory and limited (i.e., partial) predictions. At each time step, the decision maker is allowed to only query a limited number of points and obtain potentially noisy predictions of the cost functions from a finite-length prediction window at these points, and the incurred cost has a long-term dependency on the past decisions. Focusing on this partial (or bandit) prediction setting is crucial as it reflects many practical applications where the full cost function or its gradient is unavailable. For instance, when interacting with a complex physical system, a proprietary software, or a black-box simulator, it is often impossible or computationally prohibitive to determine the underlying mathematical model nesterov2017random; balasubramanian2022zeroth. Instead, the decision-maker can only perform queries by inputting a set or sequence of actions and observing the resulting cost. Such “query-and-observe” interaction precisely corresponds to the bandit feedback model studied in classic OCO problems flaxman2005online; hazan2014bandit. We summarize our approaches and contributions.
$\bullet$ Algorithm Design and Regret Analysis. In Section 3, we give an algorithm (Algorithm 1) to solve the problem of OCO with memory and limited predictions. Our algorithm extends that in li2020online for the full information setting to the partial information setting when making the predictions. We prove in Section 4 that the algorithm enjoys a dynamic regret that decays exponentially with the length of the prediction window and matches the dynamic regret of the algorithm in li2020online. Our algorithm contains two general subroutines that can be applied to solve wider classes of problems as we describe next.
$\bullet$ OCO with Memory and Bandit feedback. The first subroutine of Algorithm 1 can be applied to general OCO with memory and bandit feedback. We show in Theorem 1 our algorithm achieves a $\sqrt{TV_{T}}$-dynamic regret, where $V_{T}$ is the variation length of the optimal solution. Our dynamic regret matches that of the algorithm in zhao2022non for OCO with memory and full information of cost functions (up to logarithmic factors in $T$).
$\bullet$ Faster Convergence Rate of Zeroth-Order Method. Our core contribution lies in the second subroutine of Algorithm 1, which can be viewed as a zeroth-order (i.e., derivative-free) method for solving general (offline) convex minimization problems. For smooth and strongly convex cost functions, our algorithm achieves the same linear convergence rate as the first-order method using gradient descent nesterov2013introductory; bubeck2015convex, which improves the convergence rate of the zeroth-order method proposed in nesterov2017random. In particular, the convergence rate of our method is independent of the problem horizon length $T$, while applying the method in nesterov2017random yields a convergence rate that scales as $\mathcal{O}(1/T)$. We achieve this speedup by introducing a novel truncated Gaussian smoothing (formally introduced in Definition 1) when choosing the query points for the function value oracle, i.e., we perturb the decision point $x_{t}$ at time step $t$ by truncated Gaussian random vectors and query the perturbed points to obtain predictions of the function values.
More Related Work. First, Bandit OCO and zeroth-order methods for convex minimization are naturally related to each other since in both scenarios, a general recursive way to update the action has the form: $x_{t+1}=x_{t}-\alpha_{t}g_{t}$, where $g_{t}$ is obtained from the function value oracle and can be viewed as an estimate of the true gradient at $x_{t}$. A major difference is that the performance of algorithms for bandit OCO is characterized by regret while the performance of algorithms for zeroth-order methods is characterized by convergence rate. However, it has been shown that an algorithm for bandit OCO with regret guarantee can be converted to an algorithm for zeroth-order methods with convergence guarantee shamir2013complexity; shamir2017optimal. In this work, we propose an algorithm that combines a bandit OCO algorithm and a zeroth-order algorithm to solve the OCO problem with memory and limited predictions. Second, our algorithm design and analysis rely on the use of truncated Gaussian smoothing. The methods of Gaussian smoothing or smoothing over a sphere have been used in bandit OCO or zeroth-order methods flaxman2005online; nesterov2013introductory; berahas2022theoretical; balasubramanian2022zeroth. We show that the truncated Gaussian smoothing inherits the desired properties of Gaussian smoothing while also allowing us to prove a faster convergence rate of the zeroth-order method for smooth and strongly convex cost functions. Finally, several extensions of the framework in this paper can be considered, including extending to OCO with unbounded memory (i.e., the current cost depends on all the past decisions) kumar2024online and nonsmooth or nonconvex cost functions lin2022gradient.
Notation and terminology. The sets of integers and real numbers are denoted as $\mathbb{Z}$ and $\mathbb{R}$, respectively. For a set $\mathcal{X}$, $\mathcal{X}^{n}$ denotes the corresponding Cartesian product $\mathcal{X}\times\cdots\times\mathcal{X}$. For a real number $a$, let $\lfloor a\rfloor$ be the largest integer that is smaller than or equal to $a$. For a vector $x\in\mathbb{R}^{n}$, let $\left\|x\right\|=\sqrt{x^{\top}x}$ be its euclidean norm. Let $\left\|P\right\|_{F}=\sqrt{\operatorname{\mathrm{tr}}(PP^{\top})}$ be the Frobenius norm of $P\in\mathbb{R}^{n\times m}$. Let $I_{n}$ be an $n$-dimensional identity matrix. Given any integer $n\geq 1$, $n]\triangleq\{1,\dots,n\}$.
2 Problem Formulation
We consider an Online Convex Optimization (OCO) framework over $T\in\mathbb{Z}_{\geq 1}$ time steps. At each $t\in[T]\triangleq\{1,\dots,T\}$, a decision maker (i.e., agent) chooses an action $x_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ and incurs a cost $f_{t}(x_{t-h+1},\dots,x_{t})$, where $f_{t}:\mathbb{R}^{dh}\to\mathbb{R}$ is convex and $h\in\mathbb{Z}_{\geq 1}$. In other words, the cost incurred at any time step $t\in\mathbb{Z}_{\geq 1}$ has memory that depends on the current action $x_{t}$ and the past actions $x_{t-1},x_{t-2},\dots,x_{t-h+1}$. This “cost with memory” structure models a broad class of problems where decisions have long-term effects. Common application scenarios include control systems, where a sequence of actions affects future system states due to the dynamics of the system li2019control. The memory parameter $h$ is an intrinsic characteristic of the problem environment, representing the duration for which past actions remain influential. In practice, its value can be determined by domain knowledge of the specific system or estimated from historical data zhao2022non; zhou2023efficient. To proceed, denote
$C_{T}({\bm{x}})\triangleq\sum_{t=1}^{T}f_{t}(x_{t-h+1:t}),$ (1)
where ${\bm{x}}=(x_{1},\dots,x_{T})$, $x_{t-h+1:t}=(x_{t-h+1},x_{t-h+2},$ $\dots,x_{t-1},x_{t})$ and assume throughout this paper that $x_{t}=\bar{x}_{0}$ for all $t\leq 1$ for some initial condition $\bar{x}_{0}\in\mathcal{X}$. Now, given the complete knowledge of the functions $f_{1}(\cdot),\dots,f_{T}(\cdot)$, one can solve
${\bm{x}}^{\star}\in\operatorname*{arg\,min}_{x_{1},\dots,x_{T}\in\mathcal{X}}C_{T}(\bm{x}),$ (2)
where $\bm{x}^{\star}=(x_{1}^{\star},\dots,x_{T}^{\star})$. In the online framework described above, since the agent does not have the complete knowledge of $f_{1}(\cdot),\dots,f_{T}(\cdot)$ before making the decision $x_{t}\in\mathcal{X}$ at time step $t\in\mathbb{Z}_{\geq 1}$, solving (2) is impossible. The goal is then to design an online algorithm $\mathcal{A}$ for the agent such that the following regret $R(\mathcal{A})$ is minimized:
$R(\mathcal{A})\triangleq C_{T}(\bm{x}^{\mathcal{A}})-C_{T}(\bm{x}^{\star}),$ (3)
where $\bm{x}^{\mathcal{A}}=(x_{1},\dots,x_{T})$ is chosen by the algorithm.
Even though solving (3) requires complete knowledge of $f_{1}(\cdot),\dots,f_{T}(\cdot)$ a priori, in practice one may have access to predictions of $f_{t}(\cdot),\cdots,f_{t+W-1}(\cdot)$ for some $W\in\mathbb{Z}_{\geq 1}$ when choosing $x_{t}\in\mathcal{X}$ at time step $t\in\mathbb{Z}_{\geq 1}$ chen2015online; chen2016using; zeilinger2011real; angeli2016theoretical. However, the available predictions can be inaccurate and scarce (i.e., limited). Unlike existing works as discussed in the introduction, in this paper, we focus on a general framework to capture the inaccuracy and scarcity of the predictions. Specifically, at each time step $t\in\mathbb{Z}_{\geq 1}$, the predictions can be obtained by first running sequences of actions $\{(x_{t}^{j},\dots,x_{t+W-1}^{j})\}_{j=1}^{p}$ on a simulator and then receiving the corresponding costs $\{l_{t}(x_{t-h+1:t}^{j}),\dots,l_{t+W-1}(x_{t+W-h:t+W-1}^{j})\}_{j=1}^{p}$, where $l_{k}:\mathcal{X}^{h}\to\mathbb{R}$ for all $k\in\{t,\dots,t+W-1\}$. In other words, the simulator is associated with a model $l_{t}(\cdot)$ that describes the true cost $f_{t}(\cdot)$ with some model misspecification and can be viewed as a noisy oracle of the function values of $f_{t}(\cdot)$. Based on the available information, the agent then chooses an action $x_{t}\in\mathcal{X}$. Note that for each time step $t\in\mathbb{Z}_{\geq 1}$, one may obtain multiple cost prediction trajectories from the simulator by running multiple sequences of actions on the simulator in parallel (or running a single trajectory on multiple simulators). This prediction model, which allows querying future costs for chosen action sequences, is designed to represent practical scenarios where a high-fidelity simulator or a parallel testing environment is available. For example, in robotics, a planner can query a physics simulator to evaluate the cost of different trajectories, or in energy systems, an operator can simulate various generation schedules to predict their impact before commitment. This type of “query access” to a value-based oracle corresponds to the widely-studied bandit feedback setting in the classic OCO framework (without memory or predictions) flaxman2005online; saha2011improved; hazan2014bandit; hazan2016optimal; zhao2021bandit, since only values of the cost function are available when the agent makes decisions. Our work applies this well-established and practical model to the specific context of OCO with memory and predictions. To summarize, the agent does the following in each time step $t\geq 1$:
  • At the beginning of time step $t\geq 1$, run sequences of actions $\{(x_{t}^{j},\dots,x^{j}_{t+W-1})\}_{j=1}^{p}$ on simulator $l_{t}(\cdot)$ and receive the predictions $\{(l_{t}(x^{j}_{t-h+1:t}),\dots,$ $l_{t+W-1}(x^{j}_{t+W-h:t+W-1})\}_{j=1}^{p}$.
  • Compute $x_{t}$ based on the predictions.
  • Incur the cost $f_{t}(x_{t-h+1:t})$ and proceed to the next time step t+1t+1.111We assume that the predictions at time step $t$ can be obtained before the algorithm enters the next time step.
3 Algorithm Design
To tackle the limited predictions (i.e., bandit information) in our problem setting, we adopt the zeroth-order (i.e., derivative-free) method that has been used for both bandit OCO and (offline) convex optimization problems (e.g., flaxman2005online; nesterov2017random). To this end, we extend the algorithm design in li2019control; lin2012online for the full information setting to the bandit information setting. The detailed steps are summarized in Algorithm 1. Using similar arguments to those in li2020online, one can argue that the overall algorithm can be viewed as using a zeroth-order method for solving (2), which consists of two phases: Initialization (lines 3-7) and zeroth-order method update (lines 8-20). Specifically, let us denote $\bm{x}^{j}\triangleq(x_{1}^{j},\dots,x_{T}^{j})$ for any $j\in[0,\dots,K-1]$. In the first phase, Algorithm 1 uses an OCO subroutine to generate an initial solution sequence $\bm{x}^{0}=(x_{1}^{0},\dots,x_{T}^{0})$ that is competitive to the optimal solution $\bm{x}^{\star}$ given by (2), i.e., this sequence serves as a high-quality “warm-start” for the subsequent steps of Algorithm 1. In the second phase, Algorithm 1 employs a zeroth-order method to iteratively compute the update $\bm{x}^{j+1}=\Pi_{(1-\xi^{\prime})\mathcal{X}}(\bm{x}^{j}-\alpha\bm{g}^{j})$ for $j=0,\dots,K-1$, where $\Pi_{(1-\xi^{\prime})\mathcal{X}}(\cdot)$ denotes the projection onto set $(1-\xi^{\prime})\mathcal{X}^{T}$ and $\bm{g}^{j}=(\bm{g}^{j}_{1},\dots,\bm{g}^{j}_{T})\in\mathbb{R}^{dT}$ with $\bm{g}^{j}_{s}=\sum_{k=s}^{s+h-1}g_{k}^{j}$ for all $s\in[T]$. In particular, the term $\bm{g}^{j}$ is obtained from $l_{1}(\cdot),\dots,l_{T}(\cdot)$ queried at certain points (see lines 6 and 16 of Algorithm 1) which can be viewed as an estimate of the true gradient $\nabla C_{T}(\bm{x}^{j})$. Note that we simply set $f_{t}(\bm{x})=l_{t}(\bm{x})=0$ for all $t\leq 0$, all $t>T$ and all $\bm{x}\in\mathcal{X}^{h}$ in Algorithm 1, and recall that we have set $x_{t}^{0}=\bar{x}_{0}\in\mathcal{X}$ for all $t\leq 1$. We summarize several key features of Algorithm 1.
Feasibility of Queried and Action Points. In lines 7 and 17 of Algorithm 1, we project onto smaller sets $(1-\xi)\mathcal{X}$ and $(1-\xi^{\prime})\mathcal{X}$ compared to the original feasible set $\mathcal{X}$, where $\xi,\xi^{\prime}$ are input parameters to the algorithm. Later in Section 4, we will show that via proper choices of input parameters $\delta,\delta^{\prime},\xi,\xi^{\prime}$ to Algorithm 1, projecting onto these smaller sets ensures that both the action points $x_{t}^{K}$ (line 21) and the perturbed points $\bar{x}_{r}^{0},\tilde{x}_{r}^{0},\bar{x}^{j}_{s+h-1},\tilde{x}^{j}_{s+h-1},\bar{x}_{s}^{j+1},\tilde{x}_{s}^{j+1}$ (lines 4, 13, 19) queried for predictions stay within the feasible set $\mathcal{X}$.
Input: $\bar{x}_{0}\in\mathcal{X},\eta_{t},\alpha,\delta,\delta^{\prime}$.
1 Pick $u^{0}\sim\widetilde{\mathcal{N}}(0,I_{d})$
2 for t=2Wt=2-W to $T$ do
 /* Initialize $x_{t+W}^{0}$ using OCO */
3   Let r=t+W1r=t+W-1.
4   Set $\bar{x}_{r}^{0}=x_{r}^{0}+\delta u^{0}$ and $\tilde{x}_{r}^{0}=x_{r}^{0}-\delta u^{0}$.
5   Receive predictions $l_{r}(\bar{x}^{0}_{r-h+1:r})$ and $l_{r}(\tilde{x}^{0}_{r-h+1:r})$.
6   Set $g_{r}^{0}=\frac{1}{2\delta}\big[l_{r}(\bar{x}_{r-h+1:r}^{0})-l_{s}(\tilde{x}_{r-h+1:r}^{0})\big]u^{0}$.
7   Update $x_{r+1}^{0}=\mathbf{\Pi}_{(1-\xi)\mathcal{X}}(x_{r}^{0}-\eta_{r}g_{r}^{0})$.
 /* Obtain $x_{t}^{K}$ using zeroth-order method */
8   Let $K=\lfloor\frac{W}{h-1}\rfloor$.
9 for $j=0,\dots,K-1$ do
10      Let s=t+W(j+1)(h1)(W(h1)K)=t+(Kj1)(h1)s=t+W-(j+1)(h-1)-(W-(h-1)K)=t+(K-j-1)(h-1).
11    if $j=0$ then
12         Pick $u_{s}^{j}\overset{i.i.d}{\sim}\widetilde{\mathcal{N}}(0,I_{d})$.
13         Set $\bar{x}_{s+h-1}^{j}=x_{s+h-1}^{j}+\delta^{\prime}u_{s}^{j}$ and $\tilde{x}_{s+h-1}^{j}=x_{s+h-1}^{j}-\delta^{\prime}u_{s}^{j}$.
14         Receive predictions $l_{s+h-1}(\bar{x}^{j}_{s:s+h-1})$ and $l_{s+h-1}(\tilde{x}^{j}_{s:s+h-1})$.
15
16    for $k=s,\dots,s+h-1$ do
17         Set $g_{k}^{j}=\frac{1}{2\delta^{\prime}}\big[l_{k}(\bar{x}_{k-h+1:k}^{j})-l_{k}(\tilde{x}_{k-h+1:k}^{j})\big]u_{s}^{j}$.
18     Update $x_{s}^{j+1}=\mathbf{\Pi}_{(1-\xi^{\prime})\mathcal{X}}\big(x_{s}^{j}-\alpha\sum_{k=s}^{s+h-1}g_{k}^{j}\big)$.
19      Pick $u_{s}^{j+1}\overset{i.i.d}{\sim}\widetilde{\mathcal{N}}(0,I_{d})$.
20      Set $\bar{x}_{s}^{j+1}=x_{s}^{j+1}+\delta^{\prime}u_{s}^{j+1}$ and $\tilde{x}_{s}^{j+1}=x_{s}^{j+1}-\delta^{\prime}u_{s}^{j+1}$.
21      Receive $l_{s}(\bar{x}^{j+1}_{s-h+1:s})$ and $l_{s}(\tilde{x}^{j+1}_{s-h+1:s})$.
22
23  Play $x_{t}^{K}$ and incur the cost $f_{t}(x_{t-h+1:t}^{K})$.
Output: $x_{1}^{K},\dots,x_{T}^{K}$.
Algorithm 1 OCO with Memory and Limited Predictions
Time Complexity. The time complexity of Algorithm 1 at each time step $t\geq 1$ is dominated by the second subroutine, which performs $K=\lfloor\frac{W}{h-1}\rfloor$ update steps (line 17). Since these updates are derivative-free, the complexity is primarily determined by the number of function value queries needed for predictions. We will see in Section 4 that these update steps enable the exponential decay in regret with the prediction window size $W$. Finally, the entire algorithm operates without gradient feedback because it is fundamentally designed for the bandit feedback setting. A core premise of our problem is that the decision-maker cannot access the analytical form of the cost functions or their gradients. Interaction is limited to querying the system with an action and observing the resulting cost value. To adhere to this practical yet challenging information constraint, both subroutines are necessarily designed as derivative-free methods.
Two-Point Prediction via $l_{t}(\cdot)$. Note that in line 5, $\bar{x}^{0}_{r-h+1:r}=(x^{0}_{r-h+1},x^{0}_{r-h+2},\dots,x^{0}_{r-1},\bar{x}^{0}_{r})$ and $\tilde{x}^{0}_{r-h+1:r}=(x^{0}_{r-h+1},x^{0}_{r-h+2},\dots,x^{0}_{r-1},\tilde{x}^{0}_{r})$, where $\bar{x}^{0}_{r}$, $\tilde{x}^{0}_{r}$ are set in line 4. Similar arguments hold for lines 14 and 20. For each call of line 5, line 14 or line 20, Algorithm 1 will query two points for $l_{t}(\cdot)$ to obtain the corresponding predictions. For any $t\in\{2-W,\dots,T\}$, one can check that to obtain $x_{t}^{K}$ in line 21 of Algorithm 1, the algorithm needs to use function values of $l_{t}(\cdot)$ contained in the set $\mathcal{I}_{t}$ defined as
$\mathcal{I}_{t}\triangleq\big\{l_{k}(\bar{x}_{k-h+1:k}^{j}),l_{k}(\tilde{x}_{k-h+1:k}^{j}):k\in\{t+(K-j-1)h,\dots,t+W-1\},j\in\{0,\dots,K-1\}\big\}.$ (4)
One can also check that all the function values in $\mathcal{I}_{t}$ are available via the predictions in lines 5, 14 and 20 of Algorithm 1. Moreover, at any $t\in\{2-W,\dots,T\}$, Algorithm 1 only needs to maintain in its memory those function values in $\mathcal{I}_{t}$ (that are needed to obtain $x_{t}^{K}$) and remove the other function values from its memory. In other words, the memory of Algorithm 1 may be recursively updated through receiving fresh predictions via $l_{t}(\cdot)$ and removing the function values of $l_{t}(\cdot)$ that are no longer useful. Finally, we note that for $t=2-W,\dots,T$, the points associated with the predictions in lines 5, 14 and 20 of Algorithm 1 correspond to consecutive points on $2(K+1)$ trajectories given by $\{\bar{x}_{1}^{j},\bar{x}_{2}^{j},\dots\}_{j=0}^{K}$ and $\{\tilde{x}_{1}^{j},\tilde{x}_{2}^{j},\dots\}_{j=0}^{K}$, and thus we do not need to restart the simulator $l_{t}(\cdot)$ during the course of Algorithm 1.
Truncated Gaussian Smoothing. In Algorithm 1 (lines 4, 13 and 19), we use a truncated Gaussian random vector denoted as $u_{s}^{j}\sim\widetilde{\mathcal{N}}(0,I_{d})$ to perturb the point $x_{s}^{j}$ to obtain $\bar{x}_{s}^{j}$ and $\tilde{x}_{s}^{j}$ which are then fed to the simulator $l_{t}(\cdot)$. Specifically, for any $i\in[d]$, we let the $i$th margin $u_{s,i}^{j}$ of $u_{s}^{j}$ to be a truncated Gaussian random variable obtained from the normal distribution such that $u_{s,i}^{j}$ belongs to the interval $(1/2d^{2}(2h-1))^{1/4},1/(2d^{2}(2h-1))^{1/4}]$ and the probability density function (pdf) of $u_{s,i}^{j}$ (i.e., the $i$th margin of $\widetilde{\mathcal{N}}(0,I_{d})$) is given by
$p(x)=\frac{1}{\sqrt{2\pi}\kappa}e^{-\frac{x^{2}}{2}}\ \text{and}\ \kappa\ \text{is such that}\int_{-1/(2d^{2}(2h-1))^{\frac{1}{4}}}^{1/(2d^{2}(2h-1))^{\frac{1}{4}}}p(x)=1.$ (5)
It follows that $\lVert u_{s}^{j}\rVert\leq 1/(2(2h-1))^{1/4}\leq 1$, since $h>1$. Formally, we introduce the following johnson1995continuous.
Definition 1.
We use $u\sim\widetilde{\mathcal{N}}(0,I_{d})$ to represent an $n$-dimensional truncated Gaussian random vector whose $i$th margin $u_{i}$ satisfies $u_{i}\in[-1/(2d^{2}(2h-1))^{1/4},1/(2d^{2}(2h-1))^{1/4}]$ and the pdf of $u_{i}$ is given by (5) for all $i\in[d]$.
There exist efficient methods for sampling from a truncated Gaussian distribution (see e.g. robert1995simulation). As we discussed in the introduction, using the truncated Gaussian allows us to prove the desired regret guarantee of Algorithm 1 in the next section.
4 Regret Analysis
We now analyze the regret of Algorithm 1. To proceed, we introduce the following definitions and assumptions, which are standard in the OCO literature agarwal2010optimal; agarwal2019online; zhao2022non; flaxman2005online.
Definition 2.
A continuous differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $\beta$-smooth if
$\left\|\nabla f(x)-\nabla f(y)\right\|\leq\beta\left\|x-y\right\|,\ \forall x,y\in\mathbb{R}^{d}.\\$
Definition 3.
A differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $\mu$-strongly convex if
$f(x)-f(y)\leq\langle x-y,\nabla f(x)\rangle-\frac{\mu}{2}\left\|x-y\right\|^{2},\ \forall x,y\in\mathbb{R}^{d}.$
Definition 4.
A function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $G$-Lipschitz if
$|f(x)-f(y)|\leq G\left\|x-y\right\|,\ \forall x,y\in\mathbb{R}^{d}.$
Assumption 1.
For any $t\in[T]$, $f_{t}(\cdot)$ is (a) $\beta$-smooth, (b) $\mu$-strongly convex and (c) $G$-Lipschitz.
Assumption 2.
The feasible set $\mathcal{X}\subseteq\mathbb{R}^{d}$ is convex, contains the origin, and satisfies $\max_{x,y\in\mathcal{X}}\left\|x-y\right\|\leq D$.
Remark 1.
As we discussed in Section 3, Algorithm 1 (lines 7 and 17) projects onto smaller sets $(1-\xi)\mathcal{X}$ and $(1-\xi^{\prime})\mathcal{X}$, where $\xi,\xi^{\prime}\in\mathbb{R}_{\geq 0}$. By Assumption 2, we know that $(1-\xi)\mathcal{X}$ is also convex and contains the origin. Moreover, for any two $y,z\in(1-\xi)\mathcal{X}$, we can write $y=(1-\xi)x_{1}$ and $z=(1-\xi)x_{2}$ for some $x_{1},x_{2}\in\mathcal{X}$, which implies via Assumption 2 that $\left\|y-z\right\|=\left\|(1-\xi)(x_{1}-x_{2})\right\|=(1-\xi)\left\|x_{1}-x_{2}\right\|\leq(1-\xi)D$. Similar arguments hold for the set $(1-\xi^{\prime})\mathcal{X}$. Now, we formally show how to ensure that perturbed used points in lines 4,13,19 of Algorithm 1 stay within the set $\mathcal{X}$ as promised in Section 3. To this end, consider a perturbed point, e.g., $\bar{x}=x+\delta u$, where $x$ is kept within the smaller set $(1-\xi)\mathcal{X}$ via projection and $u\sim\tilde{\mathcal{N}}(0,I_{d})$ as per Definition 1. Recall that $\mathcal{X}$ contains the origin by Assumption 2, we know that there exists a ball $r\mathbb{B}=\{x:\left\|x\right\|\leq r\}$ with radius $r>0$ such that $r\mathbb{B}\subseteq\mathcal{X}$, which implies that any point $y\in(1-\xi)\mathcal{X}$ has a distance of at least $\xi r$ to the boundary of $\mathcal{X}$. Recalling that $\left\|u\right\|\leq 1$ by definition, to guarantee that $\bar{x}=x+\delta u$ remains within $\mathcal{X}$, we can choose $\delta\leq\xi r$. Finally, it is worth mentioning that in the two phases of Algorithm 1, we use potentially different $\xi,\xi^{\prime}$ to obtain the sets $(1-\xi)\mathcal{X},(1-\xi^{\prime})\mathcal{X}$, and potentially different $\delta,\delta^{\prime}$ to obtain the perturbed points. Such choices of $\xi,\xi^{\prime},\delta,\delta^{\prime}$ facilitate the subsequent proof of the algorithm’s regret upper bounds.
We will also assume that the model error of $l_{t}(\cdot)$ is uniformly bounded.
Assumption 3.
For any $t\in[T]$, there exists $\varphi_{t}\in\mathbb{R}_{\geq 0}$ such that $|l_{t}(\bm{x})-f_{t}(\bm{x})|\leq\varphi_{t}$ for all $\bm{x}\in\mathcal{X}^{h}$.
Remark 2.
Note that the noisy model in Assumption 3 is equivalent to assuming that $l_{t}(\bm{x})=f_{t}(\bm{x})+\phi_{t}$ for all $\bm{x}\in\mathcal{X}^{h}$ with $\phi_{t}\in\mathbb{R}$ and $\left|\phi_{t}\right|\leq\varphi_{t}$, where $\phi_{t}$ can be either deterministic or stochastic. The bounded noise model has also been considered in e.g. berahas2022theoretical. The potentially unbounded stochastic noise model has been considered in e.g. agarwal2011stochastic, and we leave the extension to such a noise model to future work.
We define several auxiliary functions that will be useful in our later analysis.
Definition 5.
(a) For any $x\in\mathbb{R}^{d}$ and any $t\in[T]$, define the $h$-unitary function $f_{t}^{c}(x)\triangleq f_{t}(x,\dots,x)$ and its smoothed version $\hat{f}_{t}^{c}(x)\triangleq\mathbb{E}_{v}[f_{t}^{c}(x+\delta v)]$, where $v\sim\widetilde{\mathcal{N}}(0,I_{d})$ and $\delta\in\mathbb{R}_{>0}$. [^2] Throughout the paper, we use $\mathbb{E}_{x}[\cdot]$ (resp., $\mathbb{E}[\cdot]$) to denote the expectation with respect to a certain random variable (or vector) $x$ (resp., all the randomness in Algorithm 1).
(b) For any $\bm{x}=(x_{1},\dots,x_{T})\in\mathbb{R}^{dT}$, $\hat{C}_{T}(\bm{x})\triangleq\mathbb{E}_{\bm{v}}[C(\bm{x}+\delta^{\prime}\bm{v})]$, where $\bm{v}\sim\widetilde{\mathcal{N}}(0,I_{dT})$ and $\delta^{\prime}\in\mathbb{R}_{>0}$.
The following result shows that the functions $\hat{f}_{t}^{c}(\cdot)$ and $\hat{C}_{T}(\cdot)$ inherit the properties from $f_{t}(\cdot)$; the proof is included in Appendix A.
Lemma 1.
Suppose Assumption 1 holds for $f_{t}(\cdot)$ for all $t\in[T]$. Then, (a) $\hat{f}_{t}^{c}(\cdot)$ is $\beta$-smooth and $\mu$-strongly convex; (b) $C_{T}(\cdot)$ is $\beta h$-smooth, $\mu$-strongly convex and $G\sqrt{Th}$-Lipschitz; and (c) $\hat{C}_{T}(\cdot)$ is $\beta h$-smooth and $\mu$-strongly convex.
More importantly, one can form unbiased gradient estimators of $\hat{f}_{t}^{c}(\cdot)$ and $\hat{C}_{T}(\cdot)$ using only the function values. Formally, we have the following result; the proof can be found in Appendix A.
Lemma 2.
Under Assumption 1(a), the following hold: (a) For $f_{t}^{c}(\cdot)$ and $\hat{f}_{t}^{c}(\cdot)$ given by Definition 5(a), any $x\in\mathbb{R}^{d}$ and any $\delta\in\mathbb{R}_{>0}$,
$\nabla\hat{f}_{t}^{c}(x)=\mathbb{E}_{v}\Big[\frac{f_{t}^{c}(x+\delta v)-f_{t}^{c}(x-\delta v)}{2\delta}v\Big]$ (6)
and
$|f_{t}^{c}(x)-\hat{f}_{t}^{c}(x)|\leq\frac{\delta^{2}}{2}\beta d.$ (7)
(b)For $C_{T}(\cdot)$ and $\hat{C}_{T}(\cdot)$ given by Definition 5(b), any $\bm{x}=(x_{1},\dots,x_{T})\in\mathbb{R}^{dT}$ and any $\delta^{\prime}\in\mathbb{R}_{>0}$,
$\frac{\partial\hat{C}_{T}(\bm{x})}{\partial x_{s}}=$ $\mathbb{E}_{v_{s}}\Big[\sum_{k=s}^{s+h-1}\frac{f_{k}(\bar{x}_{k-h+1:k})-f_{k}(\tilde{x}_{k-h+1:k})}{2\delta^{\prime}}v_{s}\Big]\qquad\forall s\in[T],$ (8)
$|C_{T}(\bm{x})-\hat{C}_{T}(\bm{x})|\leq\frac{\delta^{\prime 2}}{2}\beta hTd,$ (9)
and
$\left\|\nabla\hat{C}_{T}(\bm{x})-\nabla C_{T}(\bm{x})\right\|\leq\frac{\delta^{\prime}}{2}\beta h(Td+3)^{3/2},$ (10)
where $\bm{v}=(v_{1},\dots,v_{T})\sim\widetilde{\mathcal{N}}(0,I_{dT})$, $\bar{x}_{t}=x_{t}+\delta^{\prime}v_{t}$ and $\tilde{x}_{t}=x_{t}-\delta^{\prime}v_{t}$.
Recall from Eq. (3) that the regret of Algorithm 1 is given by $R(\mathcal{A})=C_{T}(\bm{x}^{\mathcal{A}})-C_{T}(\bm{x}^{\star})$, where $\bm{x}^{\mathcal{A}}=(x_{1}^{K},\dots,x_{T}^{K})$ is chosen by the algorithm and $\bm{x}^{\star}=(x_{1}^{\star},\dots,x_{T}^{\star})$ is the optimal solution to (1). Since Algorithm 1 contains two phases, we will split our analysis accordingly.
Initialization Phase Regret. We first upper bound the contribution of the initialization step in Algorithm 1 to $R(\mathcal{A})$, where we note from the definition of Algorithm 1 that $x_{1}^{0},\dots,x_{T}^{0}$ are chosen by the initializing phase of the algorithm.
Note that the modification of projecting onto a smaller set $(1-\xi)\mathcal{X}$ ensures the feasibility of query points in Remark 1. To analyze the regret while avoiding dependencies on the variation (also called path length) of the specific optimal solution within the smaller set, we utilize a scaled version of the true optimal sequence as our comparator. Let $\bm{x}^{\star}=(x_{1}^{\star},\dots,x_{T}^{\star})$ be the optimal solution to (1). We define a scaled comparator sequence $\hat{\bm{x}}^{\star}=(\hat{x}_{1}^{\star},\dots,\hat{x}_{T}^{\star})$ where $\hat{x}_{t}^{\star}=(1-\xi)x_{t}^{\star}$. Since $\mathcal{X}$ is convex and contains the origin, $\hat{\bm{x}}^{\star}$ lies within the smaller feasible set $((1-\xi)\mathcal{X})^{T}$ as discussed in Remark 1. We then decompose the initialization regret as follows:
$R(\mathcal{A}_{\mathrm{init}})$ $=C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})$
$=\underbrace{\left(C_{T}(\bm{x}^{0})-C_{T}(\hat{\bm{x}}^{\star})\right)}_{\text{Regret against scaled comparator}}+\underbrace{\left(C_{T}(\hat{\bm{x}}^{\star})-C_{T}(\bm{x}^{\star})\right)}_{\text{Approximation Error}}.$
The following theorem bounds the regret in the initialization phase for Algorithm 1, utilizing the fact that the path length of the scaled comparator is bounded by the path length of the true optimum. The detailed proof can be found in Appendix B.
Theorem 1.
Suppose Assumptions 1-3 hold. Let $\eta_{t}=\frac{1}{t\mu}$ in Algorithm 1. Then, the initialization step of Algorithm 1 satisfies that
$\mathbb{E}[R(\mathcal{A}_{\text{init}})]$ $\leq\Big(\frac{\sqrt{2}(1-\xi)^{2}D}{\mu}+(1-\xi)Gh^{2}\Big)V_{T}+T\delta^{2}\beta d$
$\quad+\Big(\frac{8G^{2}h^{2}+hG^{2}}{2\mu(2h-1)}+\frac{h^{3}G^{2}(1-\xi)D}{\delta(2(2h-1))^{1/2}}+\frac{3G^{2}h^{3}}{\mu(2(2h-1))^{1/2}}\Big)(1+\ln T)$
$\quad+\frac{2}{\delta^{2}\mu\sqrt{2h-1}}\sum_{t=1}^{T}\varphi_{t}^{2}+\Big(\frac{\sqrt{2}h^{2}G(1-\xi)D}{2\delta(2(2h-1))^{1/2}}+\frac{\sqrt{2}Gh^{2}+(1-\xi)D\mu}{\delta\mu(2(2h-1))^{1/4}}\Big)\sum_{t=1}^{T}\varphi_{t}$
$\quad+\xi GT\sqrt{h}D,$ (11)
with $V_{T}\triangleq\mathbb{E}\big[\sum_{t=1}^{T-1}\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|\big]$, where $\mathbb{E}[\cdot]$ denotes the expectation with respect to the randomness in Algorithm 1 and $\bm{x}^{\star}$ is the global optimal solution defined in (2).
Recalling the definition of Algorithm 1, the initialization phase of Algorithm 1 can be viewed as being applied to the cost functions $f_{1}(\cdot),\dots,f_{T}(\cdot)$ (without using any prediction) and returning $x_{1}^{0},\dots,x_{T}^{0}$. To be more specific, the initialization phase of Algorithm 1 solves the problem of OCO with memory, where the cost functions are given by $f_{t}(\cdot),\dots,f_{T}(\cdot)$ and one can check that by line 7 of Algorithm 1, the decision points $x_{1}^{0},\dots,x_{T}^{0}$ are obtained as
$x_{t+1}^{0}=\Pi_{(1-\xi)\mathcal{X}}(x_{t}^{0}-\eta_{t}g_{t}^{0}),$
[…]
Zeroth-Order Update Phase. Next, we upper bound the contribution of the zeroth-order method used in Algorithm 1 to $R(\mathcal{A})$. As we argued before in Section 3, the iterates $\bm{x}^{j}=(x_{1}^{j},\dots,x_{T}^{j})$ for $j=0,\dots,K-1$ in Algorithm 1 are obtained from the update rule $\bm{x}^{j+1}=\mathbf{\Pi}_{(1-\xi^{\prime})\mathcal{X}}(\bm{x}^{j}-\alpha\bm{g}^{j})$ starting from the initial point $\bm{x}^{0}$ (given by the initialization phase of Algorithm 1), where $\bm{g}^{j}=(\bm{g}^{j}_{1},\dots,\bm{g}^{j}_{T})\in\mathbb{R}^{dT}$ with $\bm{g}^{j}_{s}=\sum_{k=s}^{s+h-1}g_{k}^{j}$ for all $s\in[T]$ and $\bm{g}^{j}$ can be viewed as an estimate of the true gradient $\nabla C_{T}(\bm{x}^{j})$ based on the function values of $l_{t}(\cdot)$ evaluated at certain points perturbed by the truncated Gaussian smoothing (see our arguments in Section 3).
The following result characterizes the total dynamic regret of Algorithm 1 by combining the initialization regret and the convergence properties of the zeroth-order update. The detailed proof, including the decomposition of regret and the convergence analysis, is provided in Appendix B.
Theorem 2.
Consider the same setting as Theorem 1 and further let $\alpha=1/\beta^{\prime}$ in Algorithm 1, where $\beta^{\prime}=\beta h$ with $\beta$ given by Assumption 1. Then, the overall regret $R(\mathcal{A})$ of Algorithm 1 defined in (3) satisfies:
$\mathbb{E}[R(\mathcal{A})]\leq\Big(\frac{1}{1+\gamma}\Big)^{K}$ $\mathbb{E}[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})]+\frac{\varepsilon}{\gamma}+\xi^{\prime}GT\sqrt{h}D,$ (12)
where $\mathbb{E}[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})]$ is the initialization regret given in Theorem 1, $\gamma\triangleq\mu/(\beta^{\prime}-\mu)$ [^3] For any $\beta$-smooth and $\mu$-strongly convex function $f(\cdot)$, it is known that $\beta>\mu$ holds (e.g., nesterov2013introductory)., and
$\varepsilon\triangleq\frac{(1-\xi^{\prime})D\delta^{\prime}\beta h\sqrt{T}}{2\sqrt{2}(2(2h-1))^{3/4}}+\sqrt{h\beta GT\delta^{\prime}}(1-\xi^{\prime})D(Td+3)^{3/4}+\frac{(1-\xi^{\prime})Dh\sum_{t=1}^{T}\varphi_{t}}{\delta^{\prime}(2(2h-1))^{1/4}}.$
We see from Theorem 2 that for the zeroth-order update phase (lines 8-20) to be effective, the number of iterations $K$ must be at least $1$. Since $K=\lfloor W/(h-1)\rfloor$, a minimum requirement is that $W\geq h-1$. Note that we do not treat $W$ to be an input parameter to Algorithm 1; rather, the value of $W$ depends on the specific problem setup. In practice, $W$ captures the ability to obtain predictions using the simulator $l_{t}(\cdot)$. Specifically, Theorem 2 shows that the dynamic regret decays exponentially with $K$, and thus with $W$, which suggests a larger $W$ is always beneficial for achieving lower regret. However, in many real-world applications, obtaining reliable predictions becomes increasingly difficult as the prediction horizon grows, since the quality and availability of future information may degrade significantly for a longer prediction window.
Moreover, the convergence of the first term on the right-hand side of (12) achieves the same linear convergence rate (i.e., $(1-\mu/\beta^{\prime})^{K}$) as the standard first-order method based on gradient descent for smooth and strongly convex cost functions nesterov2013introductory. Supposing $\varphi_{t}=0$ for all $t\geq 0$, the second term on the right-hand side of (12) can be made arbitrarily small by setting $\delta^{\prime}$. A careful inspection of the proof of Theorem 2 reveals that the result in Theorem 2 still holds if we remove Assumption 1(c). The reason is that in the proof, we only need an upper bound on $\nabla C_{T}(\bm{x})$ for all $\bm{x}\in\mathcal{X}^{dT}$. Since $C_{T}(\cdot)$ is $\beta\sqrt{2h}$-smooth from Lemma 1, i.e., $\nabla C_{T}(\cdot)$ is continuous, and $\mathcal{X}^{dT}$ is bounded by Assumption 2, such an upper bound on $\nabla C_{T}(\cdot)$ always exists. Based on the above arguments, under the smoothness and strong convexity assumption and setting $\varphi_{t}=0$ for all $t\geq 0$, our convergence rate given in Theorem 2 for the zeroth-order subroutine in Algorithm 1 improves that of the zeroth-order method provided in nesterov2017random. [^4] Since nesterov2017random considers the setting of $\varphi_{t}=0$, we also $\varphi_{t}=0$ for comparison. Specifically, the zeroth-order method in nesterov2017random also has the form $\bm{x}^{j+1}=\bm{x}^{j}-\alpha\bm{g}^{j}$ as we described above, where $\bm{g}^{j}$ is obtained based on function values at certain points perturbed by a Gaussian smoothing, i.e., $u_{s}^{j+1}$ in line 12 of Algorithm 1 is drawn from a Gaussian distribution $\mathcal{N}(0,I_{d})$. Setting the step size to satisfy $\alpha=\frac{1}{4(Td+4)\beta^{\prime}}$, where $Td$ is the dimension of the variable $\bm{x}^{j}$, nesterov2017random provides a convergence
$\mathbb{E}\big[C_{T}(\bm{x}^{K})-C_{T}(\bm{x}^{\star})\big]\leq\Big(1-\frac{\mu}{8(Td+4)\beta^{\prime}})\Big)^{K}\big(C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})\big)+\mathcal{O}(\frac{\delta^{\prime}T^{2}d^{2}\beta^{\prime}}{\mu}).$ (13)
Comparing (13) to (12) and noting that $\gamma=\mu/(\beta^{\prime}-\mu)$, we see that Algorithm 1 achieves a faster convergence around a neighborhood of the optimal solution $C_{T}(\bm{x}^{\star})$ and the convergence rate of Algorithm 1 does not degrade as the horizon length $T$ increases.
Furthermore, the zeroth-order method based on Gaussian smoothing considered in nesterov2017random provides no guarantee that its perturbed query points (e.g., $\bar{x}=x+\delta u$ with $u\sim\mathcal{N}(0,I_{d})$) remain within the feasible set $\mathcal{X}$, due to the fact that the norm of the Gaussian random vector $u$ may not be bounded. In contrast, as we argued in Remark 1, we can guarantee that the perturbed query points (e.g., $\bar{x}=x+\delta u$ with $u\sim\tilde{N}(0,I_{d})$) remain within the feasible set $\mathcal{X}$, by projecting $x$ onto smaller set $(1-\xi)\mathcal{X}$ and leverging the fact that the truncated Gaussian random vector $u$ always has bounded norm.
Overall Bound on the Regret $R(\mathcal{A})$ of Algorithm 1. Combining the result in Theorems 1-2 gives the desired dynamic regret of Algorithm 1 for solving OCO with memory and limited predictions. The total regret consists of an exponentially decaying term of the initialization regret with respect to $K=\lfloor W/(h-1)\rfloor$, plus the error terms $\varepsilon/\gamma$ and $\xi^{\prime}GT\sqrt{h}D$. Supposing $\varphi_{t}=0$ for all $t\geq 0$, the two error terms can be made arbitrarily small by setting $\delta^{\prime}$, and the overall regret Algorithm 1 reduces to the exponential decay term of the initialization regret, which matches that of the algorithm proposed in li2020online for the full information setting with one-step memory (i.e., $h=1$).
5 Numerical Experiments
This section provides numerical results to complement our theoretical analysis and demonstrate the performance of our proposed algorithm.
5.1 Experimental Setup
We consider an unconstrained quadratic programming problem, a standard benchmark in convex optimization, where the total cost is given by $C(x)=\sum_{t=1}^{T}f_{t}(x_{t-h+1:t})$. The cost at each time step is $f_{t}(x_{t-h+1:t})=\frac{1}{2}x_{t-h+1:t}^{\top}A_{t}x_{t-h+1:t}+B_{t}^{\top}x_{t-h+1:t}$. For all experiments, we consider a scalar decision variable $x_{t}\in\mathbb{R}$ (i.e., $d=1$) and a memory of $h=2$. The matrices $A_{t}\in\mathbb{R}^{2\times 2}$ and vectors $B_{t}\in\mathbb{R}^{2}$ are randomly generated at each run, with $A_{t}$ being constructed as positive definite to ensure that the strong convexity and smoothness properties from Assumption 2 hold. The algorithm parameters are set according to our theoretical analysis: $x_{0}=0.5,\eta_{t}=0.2/t,\alpha=0.05,\delta=0.2,\delta^{\prime}=0.0001$. We assume a perfect simulator, i.e., $l_{t}(x)=f_{t}(x)$ for all $t$ ($\phi_{t}=0$). We evaluate the performance of our algorithm under three different smoothing distributions used for the two-point feedback gradient estimation:
  • (a)
    Gaussian distribution: The random direction $u$ is drawn from a standard normal distribution, $u\sim\mathcal{N}(0,1)$.
  • (b)
    Truncated Gaussian distribution: The random direction $u$ is drawn from a normal distribution truncated to the interval [2,2][-2,2], which reduces the variance of the gradient estimate.
  • (c)
    Bernoulli distribution: The random direction $u$ is drawn from a Bernoulli distribution, taking values in $\{-1,1\}$ with equal probability.
All results are generated by averaging over 100 independent experimental runs. The plots show the median regret, with the shaded areas representing the interquartile range (from the 25th to the 75th percentile).
5.2 Validation of Initialization Phase Regret (Theorem 1)
First, we empirically validate the performance of the initialization phase of our algorithm (lines 3-8 in Algorithm 1), which is an Online Convex Optimization (OCO) procedure with memory. Theorem 1 establishes a dynamic regret bound of $\mathcal{O}(TlnT)$ for this phase. To verify this, we run the initialization phase over a long time horizon, varying $T$ from 5 to 500, and compute the regret of the generated sequence $x_{0}$. Figure 1 displays the normalized regret for all three smoothing distributions. The subplots show that the regret grows sub-linearly with $T$, as evidenced by the decreasing curves in the $Reg/T$ plots. Furthermore, the $Reg/T$ plots show a slowly growing trend, consistent with the theoretical $\ln T$ factor. These results provide strong empirical support for the regret bound established in Theorem 1. Among the smoothing methods, the lower-variance distributions (Truncated Gaussian and Bernoulli) tend to exhibit slightly better and more stable performance than the standard Gaussian.
(a) Gaussian
(b) Truncated Gaussian
(c) Bernoulli
Figure 1:
5.3 Regret Decay with Prediction Window (Theorem 2)
Next, we investigate the core claim of our paper: the total regret of the full algorithm decays exponentially with the length of the prediction window, $W$. This behavior is a direct consequence of the linear convergence of the zeroth-order phase, as analyzed in Theorem 2. For this experiment, we fix the time horizon at $T=400$ and vary the prediction window $W$ from 4 to 50. The regret is calculated based on the final output sequence $x_{K}$, where $K=\lfloor W/(h-1)\rfloor$ is the number of zeroth-order iterations. Figure 2 plots the natural logarithm of the regret, $\ln(Regret)$, against $W$. As depicted, all three smoothing methods show a clear, approximately linear decrease on this semi-log plot. This linear trend provides strong empirical evidence of the exponential decay of regret with respect to the prediction window size. This result validates the convergence analysis of Theorem 2, demonstrating that as more future information is made available (larger $W$), the algorithm performs more optimization iterations and rapidly converges to a better solution, drastically reducing the overall regret. Finally, the performances of using the Bernoulli random variable and the Gaussian random variable for smoothing are comparable to the case of using the truncated Gaussian random variable; however, there are no theoretical guarantees of the regret for Bernoulli smoothing or Gaussian smoothing, which are left to future work.
Figure 2: Log of regret for the full algorithm versus the prediction window size $W$, with $T=400$. The linear downward trend for all smoothing methods empirically confirms the theoretical exponential decay of regret, a key result derived from Theorem 2.
5.4 Convergence Rate Comparison with nesterov2017random
Finally, to highlight the advantages of our zeroth-order method based on truncated Gaussian smoothing, we directly compare its convergence rate against the zeroth-order method proposed by nesterov2017random based on Gaussian smoothing. In this experiment, we fix the horizon to be $T=200$. Both algorithms start from the same initial sequence $x_{0}$, generated by our OCO initialization phase. Figure 3 aggregates the performance curves under different step-size settings into a single plot for direct comparison. As we argued in Section 4, our algorithm allows for a large step size $\alpha=1/\beta^{\prime}$ compared to the step size $\alpha=1/(T\beta^{\prime})$ (here $\beta^{\prime}=20$ according to the setup in Section 5.1) that can be chosen for the algorithm in nesterov2017random, which aligns with the results in Figure 3. Moreover, when choosing the step size to be $\alpha=1/\beta^{\prime}$ for the algorithm in nesterov2017random, we see from Figure 3 that the algorithm does not converge. This divergence occurs because the method in nesterov2017random theoretically requires a step size scaling as $\mathcal{O}(1/T)$ to ensure stability, whereas our method robustly supports a much larger step size (scaling as $\mathcal{O}(1)$). These results empirically validate that our approach not only achieves faster convergence but also offers better stability under a more aggressive step size.
Figure 3: Convergence Rate Comparison (T=200)
6 Conclusion
We studied the problem of OCO with memory and predictions over a horizon $T$. We proposed an algorithm to solve this problem and showed that the dynamic regret of the algorithm decays exponentially with the prediction window length. The first subroutine of the algorithm can be used to solve general online convex optimization with memory and bandit feedback with $\sqrt{T}$-dynamic regret with respect to $T$. The second subroutine of the algorithm is a zeroth-order method that can be used to solve general convex optimization problems with a linear convergence rate that improves the rate of zeroth-order algorithms and matches the rate of first-order algorithms for convex optimization problems. Our algorithm design and analysis rely on truncated Gaussian smoothing when querying the decision points for obtaining the predictions. Future work includes extending the algorithm design and analysis to more general (i.e., nonconvex or nonsmooth) cost functions and exploring other possible smoothing methods for zeroth-order algorithms.
Appendix A Proofs of Preliminary Results
A.1 Proof of Lemma 1
We first prove the properties of $\hat{f}_{t}^{c}(\cdot)$. Note that the $\mu$-strongly convexity of $f^{c}_{t}(\cdot)$ follows directly from that of $f_{t}(\cdot)$. Thus, for any $x,y\in\mathbb{R}^{d}$,
$f_{t}^{c}(x+\delta v)-f_{t}^{c}(y+\delta v)\leq\nabla f_{t}^{c}(x+\delta v)^{\top}(x-y)-\frac{\mu}{2}\left\|x-y\right\|^{2}.$
Taking the expectation with respect to $v\sim\widetilde{\mathcal{N}}(0,I_{d})$ on both sides of the above inequality, we get
$\hat{f}_{t}^{c}(x)-\hat{f}_{t}^{c}(y)$ $\leq\mathbb{E}_{v}\big[\nabla f_{t}^{c}(x+\delta v)\big]^{\top}(x-y)-\frac{\mu}{2}\left\|x-y\right\|^{2}$
$\leq\nabla\hat{f}_{t}^{c}(x)^{\top}(x-y)-\frac{\mu}{2}\left\|x-y\right\|^{2},$ (A.1)
where we can swap the order of $\mathbb{E}_{v}$ and $\nabla$ since $\nabla f_{t}^{c}(\cdot)$ is continuous in $x$. It follows from (A.1) that $\hat{f}_{t}^{c}(\cdot)$ is $\mu$-strongly convex. Similarly, the $\beta$-smoothness of $f_{t}^{c}(\cdot)$ follows from that of $f_{t}(\cdot)$. For any $x,y\in\mathbb{R}^{d}$, we have
$\left\|\nabla f_{t}^{c}(x+\delta v)-\nabla f_{t}^{c}(y+\delta v)\right\|\leq\beta\left\|x-y\right\|.$
Again, taking the expectation $\mathbb{E}_{v}[\cdot]$ on both sides of the above inequality yields
$\mathbb{E}_{v}\big[\left\|\nabla f_{t}^{c}(x+\delta v)-\nabla f_{t}^{c}(y+\delta v)\right\|\big]\leq\beta\left\|x-y\right\|,$
$\overset{(a)}{\Rightarrow}$ $\left\|\mathbb{E}_{v}[\nabla f_{t}^{c}(x+\delta v)]-\mathbb{E}_{v}[\nabla f_{t}^{c}(y+\delta v)]\right\|\leq\beta\left\|x-y\right\|,$
$\Rightarrow$ $\left\|\nabla\hat{f}_{t}^{c}(x)-\nabla\hat{f}_{t}^{c}(y)\right\|\leq\beta\left\|x-y\right\|,$
where (a) follows from Jensen’s inequality (i.e., $\mathbb{E}[\left\|a\right\|]\geq\left\|\mathbb{E}[a]\right\|$ for any $a\in\mathbb{R}^{d}$), which shows that $\hat{f}_{t}^{c}(\cdot)$ is $\beta$-smooth.
Next, we prove the properties of $C_{T}(\cdot)$. Recall that $C_{T}(\bm{x})=\sum_{t=1}^{T}f_{t}(x_{t-h+1:t})$, where $\bm{x}=(x_{1},\dots,x_{T})\in\mathbb{R}^{dT}$ and $x_{t}=x_{0}$ for all $t\leq 0$. Considering any $\bm{x}=(x_{1},\dots,x_{T})\in\mathbb{R}^{dT}$ and any $\bm{y}=(y_{1},\dots,y_{T})\in\mathbb{R}^{dT}$, we have
$\left\|\bm{x}-\bm{y}\right\|^{2}\leq\sum_{t=1}^{T}\left\|x_{t-h+1:t}-y_{t-h+1:t}\right\|^{2}.$ (A.2)
Moreover, we have
$\sum_{t=1}^{T}\left\|x_{t-h+1:t}-y_{t-h+1:t}\right\|^{2}\leq h\left\|\bm{x}-\bm{y}\right\|^{2},$ (A.3)
$\Rightarrow$ $\sqrt{\sum_{t=1}^{T}\left\|x_{t-h+1:t}-y_{t-h+1:t}\right\|^{2}}\leq\sqrt{h}\left\|\bm{x}-\bm{y}\right\|,$
$\overset{(a)}{\Rightarrow}$ $\sum_{t=1}^{T}\left\|x_{t-h+1:t}-y_{t-h+1:t}\right\|\leq\sqrt{Th}\left\|\bm{x}-\bm{y}\right\|,$ (A.4)
where (a) uses the fact that $\sqrt{\sum_{t=1}^{T}a_{t}^{2}}\geq\frac{\sum_{t=1}^{T}a_{t}}{\sqrt{T}}$ for any $a_{t}\in\mathbb{R}$. By the $\mu$-strongly convexity of $f_{t}(\cdot)$, for any $t\in[T]$,
$f_{t}(x_{t-h+1:t})-f_{t}(y_{t-h+1:t})\leq\nabla f_{t}(x_{t-h+1:t})^{\top}(x_{t-h+1:t}-y_{t-h+1:t})-\frac{\mu}{2}\left\|x_{t-h+1:t}-y_{t-h+1:t}\right\|^{2}.$ (A.5)
To relate $\nabla f_{t}(x_{t-h+1:t})$ to $\nabla C_{T}(\bm{x})$, we first augment the gradient vector $\nabla f_{t}(x_{t-h+1:t})\in\mathbb{R}^{dh}$ to a higher dimensional vector $\tilde{\nabla}f_{t}(x_{t-h+1:t})\in\mathbb{R}^{dT}$ by adding zero entries. By the definition of $C_{T}(\cdot)$, we see that $\nabla C_{T}(\bm{x})=\sum_{t=1}^{T}\tilde{\nabla}f_{t}(x_{t-h+1:t})$. Summing (A.5) over $t\in[T]$ now gives us
$C_{T}(\bm{x})-C_{T}(\bm{y})\leq\sum_{t=1}^{T}\tilde{\nabla}f_{t}(x_{t-h+1:t})^{\top}(\bm{x}-\bm{y})-\frac{\mu}{2}\sum_{t=1}^{T}\left\|x_{t-h+1:y}-y_{t-h+1:t}\right\|^{2},$
[…]
which proves that $C_{T}(\cdot)$ is $\mu$-strongly convex. To show the smoothness of $C_{T}(\cdot)$, we again start from $f_{t}(\cdot)$ and its $\beta$-smoothness gives us [^5] It is well known that a $\beta$-smooth function $f:\mathbb{R}^{d}\to\mathbb{R}$ as per definition 2 satisfies that $f(y)-f(x)\leq\left\langle\nabla f(x),y-x\right\rangle+\frac{\beta}{2}\left\|y-x\right\|^{2},\ \forall x,y\in\mathbb{R}^{d}$ [bubeck2015convex].
$f_{t}(y_{t-h+1:t})-f_{t}(x_{t-h+1:t})\leq\nabla f_{t}(x_{t-h+1:t})^{\top}(y_{t-h+1:t}-x_{t-h+1:t})+\frac{\beta}{2}\left\|x_{t-h+1:t}-y_{t-h+1:t}\right\|^{2}.$ (A.6)
Summing (A.6) over $t\in[T]$ gives us
$C_{T}(\bm{y})-C_{T}(\bm{x})\leq\sum_{t=1}^{T}\tilde{\nabla}f_{t}(x_{t-h+1:t})^{\top}(\bm{y}-\bm{x})+\frac{\beta}{2}\sum_{t=1}^{T}\left\|x_{t-h+1:y}-y_{t-h+1:t}\right\|^{2},$
[…]
which proves that $C_{T}(\cdot)$ is $(\beta h)$-smooth. To show that that $C_{T}(\cdot)$ is $G\sqrt{Th}$-Lipschitz, we have
[…]
A.2 Proof of Lemma 2
Proof of part (a) Eq. (6) follows directly from [nesterov2017random, Eq. (26)]. Since $f_{t}^{c}(\cdot)$ is $\beta$-smooth by Assumption 1(a), Eq. (7) follows from [nesterov2017random, Theorem 1].
Proof of part (b) Recalling the definition of $\hat{C}_{T}(\cdot)$, we have
$\hat{C}_{T}(\bm{x})=\sum_{t=1}^{T}\mathbb{E}_{\bm{v}}[f_{t}(\bar{x}_{t-h+1:t})],$ (A.8)
where we set $\bar{x}_{t}=x_{0}$ for all $t\leq 0$. It follows that for any $s\in[T]$,
$\frac{\partial\hat{C}_{T}(\bm{x})}{\partial x_{s}}$ $=\frac{\partial\mathbb{E}_{\bm{v}}\big[\sum_{t=1}^{T}f_{t}(\bar{x}_{t-h+1:t})\big]}{\partial x_{s}}$
$\overset{(a)}{=}\mathbb{E}_{\bm{v}}\Big[\sum_{t=1}^{T}\frac{\partial f_{t}(\bar{x}_{t-h+1:t})}{\partial x_{s}}\Big]$
$=\mathbb{E}_{\bm{v}}\Big[\sum_{k=s}^{s+h-1}\frac{\partial f_{k}(\bar{x}_{k-h+1:k})}{\partial x_{s}}\Big]$
$\overset{(b)}{=}\frac{\partial\mathbb{E}_{\bm{v}}\big[\sum_{k=s}^{s+h-1}f_{k}(\bar{x}_{k-h+1:k})\big]}{\partial x_{s}},$ (A.9)
where we can swap the order of $\mathbb{E}_{\bm{v}}$ and $\frac{\partial}{\partial x_{s}}$ in (a) and (b) since $\frac{\partial f_{t}(\bar{x}_{t-h+1:t})}{\partial x_{s}}$ is continuous in $x_{s}$ for all $t\in[T]$. Now, we notice that $\sum_{k=s}^{s+h-1}f_{k}(x_{k-h+1:k})$ may be viewed as a function of $x_{s-h+1},\dots,x_{s+h-1}$, denoted as $F_{s}(x_{s-h+1:s+h-1})$. Since $\bm{v}=(v_{1},\dots,v_{T})\sim\widetilde{\mathcal{N}}(0,I_{dT})$, we have $v_{t}\sim\widetilde{\mathcal{N}}(0,I_{d})$ for all $t\in[T]$. Letting $\bm{v}_{s}=v_{s-h+1:s+h-1}\in\mathbb{R}^{2h-1}$ be a column vector and also recalling that $\bar{x}_{t}=x_{t}+\delta^{\prime}v_{t}$ and $\tilde{x}_{t}=x_{t}-\delta^{\prime}v_{t}$, we again have from [nesterov2017random, Eq. (26)] that
$\nabla\mathbb{E}_{\bm{v}_{s}}[F_{s}(\bar{x}_{s-h+1:s+h-1})]$ $=\mathbb{E}_{\bm{v}_{s}}\Big[\frac{F(\bar{x}_{s-h+1:s+h-1})-F(\tilde{x}_{s-h+1:s+h-1})}{2\delta^{\prime}}\bm{v}_{s}\Big]$
$=\mathbb{E}_{\bm{v}_{s}}\Big[\frac{\sum_{k=s}^{s+h-1}\big(f_{k}(\bar{x}_{k-h+1:k})-f_{k}(\tilde{x}_{k-h+1:k})}{2\delta^{\prime}}\bm{v}_{s}\Big].$ (A.10)
[…]
Appendix B Proofs of Main Results
B.1 Proof of Theorem 1
As outlined in the main text, the regret of the initialization phase, $R(\mathcal{A}_{\text{init}})$, is defined against the true optimal solution $\bm{x}^{\star}=(x_{1}^{\star},\dots,x_{T}^{\star})$. To avoid dependencies on the specific optimum of the constrained set $(1-\xi)\mathcal{X}$, we construct a comparison sequence $\hat{\bm{x}}^{\star}=(\hat{x}_{1}^{\star},\dots,\hat{x}_{T}^{\star})$ by scaling the global optimal solution:
$\hat{x}_{t}^{\star}=(1-\xi)x_{t}^{\star},\quad\forall t\in[T].$ (B.1)
Since we assume the feasible set $\mathcal{X}$ contains the origin (or is centered such that this scaling is valid), and $\mathcal{X}$ is convex, it holds that $\hat{x}_{t}^{\star}\in(1-\xi)\mathcal{X}$. We decompose the total regret as follows:
$\mathbb{E}[R(\mathcal{A}_{\text{init}})]$ $=\mathbb{E}\Big[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})\Big]$
[…]
We will now bound each part separately.
Part 1: Bounding the Regret against $\hat{\bm{x}}^{\star}$. This part bounds the term $\mathbb{E}\big[C_{T}(\bm{x}^{0})-C_{T}(\hat{\bm{x}}^{\star})\big]$. The proof proceeds in several steps.
step 1: Preparatory Terminology. Based on the definition of Algorithm 1, let us further define an auxiliary function
$h_{t}(x)=\hat{f}_{t}^{c}(x)+(\tilde{g}_{t}-\nabla\hat{f}_{t}^{c}(x_{t}^{0}))x,$
where $\tilde{g}_{t}=\frac{1}{2\delta}\big[f_{t}^{c}(\bar{x}_{t}^{0})-f^{c}_{t}(\tilde{x}_{t}^{0})\big]u_{t}^{0}$. The above definitions imply $\nabla h_{t}(x_{t}^{0})=\tilde{g}_{t}$. Moreover, we have from Lemma 2(a) that $\mathbb{E}_{u_{t}}[\tilde{g}_{t}]=\nabla\hat{f}_{t}^{c}(x_{t}^{0})$ and thus $\mathbb{E}_{u_{t}}[h_{t}(x)]=\hat{f}_{t}^{c}(x)$ (for all $x$ that is independent of $u_{t}$), which further implies that $\mathbb{E}[h_{t}(x_{t}^{0})]=\mathbb{E}[\hat{f}_{t}^{c}(x_{t}^{0})]$ for all $t$. For our analysis in the sequel, we show that for any $t\in[T]$,
$\left\|\nabla h_{t}(x_{t}^{0})\right\|$ $=\big\lVert\frac{1}{2\delta}\big[f_{t}^{c}(\bar{x}_{t}^{0})-f_{t}^{c}(\tilde{x}_{t}^{0})\big]u^{0}\big\rVert$
$\leq\frac{1}{2\delta}|f_{t}^{c}(\bar{x}_{t}^{0})-f_{t}^{c}(\tilde{x}_{t}^{0})|\left\|u^{0}\right\|$
$\overset{(a)}{\leq}\frac{1}{2\delta}G\sqrt{h}\left\|\bar{x}_{t}^{0}-\tilde{x}_{t}^{0}\right\|\left\|u^{0}\right\|$
$=\frac{1}{2\delta}G\sqrt{h}\left\|2\delta u^{0}\right\|\left\|u^{0}\right\|=G\sqrt{h}\left\|u^{0}\right\|^{2},$ (B.2)
where $(a)$ follows from the $G$-Lipschitz of $f_{t}(\cdot)$ in Assumption 1(c) and the definition of $f_{t}^{c}(\cdot)$. Let us denote $e_{t}=e_{t}^{m}+e_{t}^{c}$ with
$e_{t}^{m}=\frac{1}{2\delta}\Big[\big(l_{t}(\bar{x}_{t-h+1:t}^{0})-l_{t}(\tilde{x}_{t-h+1:t}^{0})\big)-\big(f_{t}(\bar{x}_{t-h+1:t}^{0})-f_{t}(\tilde{x}_{t-h+1:t}^{0})\big)\Big]u^{0},$ (B.3)
$e_{t}^{c}=\frac{1}{2\delta}\Big[\big(f_{t}(\bar{x}_{t-h+1:t}^{0})-f_{t}(\tilde{x}_{t-h+1:t}^{0})\big)-\big(f_{t}^{c}(\bar{x}_{t}^{0})-f_{t}^{c}(\tilde{x}_{t}^{0})\big)\Big]u^{0}.$ (B.4)
Step 2: Relate Total Cost $C_{T}(\cdot)$ to Unitary Cost $f_{t}^{c}(\cdot)$. First, we relate the algorithmic regret in terms of $C_{T}(\cdot)$ to the regret in terms of the $h$-unitary function $f_{t}^{c}(x)=f_{t}(x,\dots,x)$.
$C_{T}(\bm{x}^{0})-C_{T}(\hat{\bm{x}}^{\star})=\sum_{t=1}^{T}\Big(f_{t}(x_{t-h+1:t}^{0})-f_{t}(\hat{x}_{t-h+1:t}^{\star})\Big).$
We can rewrite this as:
$\sum_{t=1}^{T}\Big(f_{t}^{c}(x_{t}^{0})-f_{t}^{c}(\hat{x}_{t}^{\star})\Big)+\sum_{t=1}^{T}\Big(f_{t}(x_{t-h+1:t}^{0})-f_{t}^{c}(x_{t}^{0})\Big)-\sum_{t=1}^{T}\Big(f_{t}(\hat{x}_{t-h+1:t}^{\star})-f_{t}^{c}(\hat{x}_{t}^{\star})\Big).$
We bound the expectation of the second term using the Lipschitz property (Assumption 1(c)):
$\quad\mathbb{E}\Big[\Big|\sum_{t=1}^{T}f_{t}(x_{t-h+1:t}^{0})-f_{t}^{c}(x_{t}^{0})\Big|\Big]$
$\leq G\sum_{t=1}^{T}\sum_{j=1}^{h-1}\mathbb{E}\big[\left\|x_{t}^{0}-x_{t-j}^{0}\right\|\big]\leq G\sum_{t=1}^{T}\sum_{j=1}^{h-1}\sum_{l=1}^{j}\mathbb{E}\big[\left\|x_{t-l+1}^{0}-x_{t-l}^{0}\right\|\big]$
$\overset{(a)}{\leq}G\sum_{t=1}^{T}\sum_{j=1}^{h-1}\sum_{l=1}^{j}\eta_{t-l}\mathbb{E}\big[\left\|\nabla h(x_{t-l}^{0})\right\|+\left\|e_{t-l}\right\|\big]$
$\overset{(b)}{\leq}\frac{1}{\sqrt{2(2h-1)}}h^{5/2}G^{2}\sum_{t=1}^{T}\eta_{t}+Gh^{2}\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|$
$\overset{(c)}{\leq}\frac{h^{5/2}G^{2}}{\sqrt{2(2h-1)}\mu}(1+\ln T)+\frac{\sqrt{2}Gh^{2}}{\delta\mu(2(2h-1))^{1/4}}\sum_{t=1}^{T}\varphi_{t}+\frac{2G^{2}h^{3}}{\mu(2(2h-1))^{1/2}}(1+\ln T)$
$\leq\frac{3G^{2}h^{3}}{\mu(2(2h-1))^{1/2}}(1+\ln T)+\frac{\sqrt{2}Gh^{2}}{\delta\mu(2(2h-1))^{1/4}}\sum_{t=1}^{T}\varphi_{t},$ (B.5)
where (a) follows from Lemma E.1, (b) from (B.2) and $\lVert u^{0}\rVert^{2}\leq\sqrt{1/2(2h-1)}$, and (c) from (C.9). Then, we bound the third term (which is deterministic) similarly:
$\Big|\sum_{t=1}^{T}f_{t}(\hat{x}_{t-h+1:t}^{\star})-f_{t}^{c}(\hat{x}_{t}^{\star})\Big|$ $\leq G\sum_{t=1}^{T}\sum_{j=1}^{h-1}\sum_{l=1}^{j}\left\|\hat{x}_{t-l+1}^{\star}-\hat{x}_{t-l}^{\star}\right\|$
$\leq Gh^{2}\sum_{t=1}^{T}\left\|\hat{x}_{t}^{\star}-\hat{x}_{t-1}^{\star}\right\|$
[…]
where $V_{T}=\sum_{t=1}^{T-1}\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|$ is the path-length discussed before.
Step 3: Relate Unitary Cost $f_{t}^{c}(\cdot)$ to Smoothed Cost $\hat{f}_{t}^{c}(\cdot)$. Next, we relate $f_{t}^{c}(\cdot)$ to its smoothed version $\hat{f}_{t}^{c}(\cdot)$ using Lemma 2(a), equation (7):
$|f_{t}^{c}(x)-\hat{f}_{t}^{c}(x)|\leq\frac{\delta^{2}}{2}\beta d,$
which implies:
$f_{t}^{c}(x_{t}^{0})-f_{t}^{c}(\hat{x}_{t}^{\star})$ $=\big(f_{t}^{c}(x_{t}^{0})-\hat{f}_{t}^{c}(x_{t}^{0})\big)+\big(\hat{f}_{t}^{c}(x_{t}^{0})-\hat{f}_{t}^{c}(\hat{x}_{t}^{\star})\big)+\big(\hat{f}_{t}^{c}(\hat{x}_{t}^{\star})-f_{t}^{c}(\hat{x}_{t}^{\star})\big)$
$\leq\frac{\delta^{2}\beta d}{2}+\big(\hat{f}_{t}^{c}(x_{t}^{0})-\hat{f}_{t}^{c}(\hat{x}_{t}^{\star})\big)+\frac{\delta^{2}\beta d}{2}.$
Summing over $t$, we get:
$\sum_{t=1}^{T}\Big(f_{t}^{c}(x_{t}^{0})-f_{t}^{c}(\hat{x}_{t}^{\star})\Big)\leq\sum_{t=1}^{T}\Big(\hat{f}_{t}^{c}(x_{t}^{0})-\hat{f}_{t}^{c}(\hat{x}_{t}^{\star})\Big)+T\delta^{2}\beta d.$ (B.7)
Step 4: Analyze the OCO Update on the Smoothed Cost. We now analyze the core OCO regret term $\sum_{t=1}^{T}\big(\hat{f}_{t}^{c}(x_{t}^{0})-\hat{f}_{t}^{c}(\hat{x}_{t}^{\star})\big)$ in the right-hand side of (B.7). We use the auxiliary function (which has been defined in step 1.0) $h_{t}(x)=\hat{f}_{t}^{c}(x)+(\tilde{g}_{t}-\nabla\hat{f}_{t}^{c}(x_{t}^{0}))x$, where $\tilde{g}_{t}=\frac{1}{2\delta}\big[f_{t}^{c}(\bar{x}_{t}^{0})-f^{c}_{t}(\tilde{x}_{t}^{0})\big]u_{t}^{0}$. By construction, $\mathbb{E}_{u_{t}}[h_{t}(x)]=\hat{f}_{t}^{c}(x)$ for any $x$ independent of $u_{t}$. Thus, $\mathbb{E}[h_{t}(x_{t}^{0})]=\mathbb{E}[\hat{f}_{t}^{c}(x_{t}^{0})]$ and $\mathbb{E}[h_{t}(\hat{x}_{t}^{\star})]=\mathbb{E}[\hat{f}_{t}^{c}(\hat{x}_{t}^{\star})]$. We therefore need to bound $\mathbb{E}\Big[\sum_{t=1}^{T}\big(h_{t}(x_{t}^{0})-h_{t}(\hat{x}_{t}^{\star})\big)\Big]$.
The update rule is $x_{t+1}^{0}=\mathbf{\Pi}_{(1-\xi)\mathcal{X}}(x_{t}^{0}-\eta_{t}(\nabla h_{t}(x_{t}^{0})+e_{t}))$, where $e_{t}=g_{t}^{0}-\tilde{g}_{t}=e_{t}^{m}+e_{t}^{c}$. Since $\hat{f}_{t}^{c}(\cdot)$ is $\mu$-strongly convex (Lemma 1), $h_{t}(\cdot)$ is also $\mu$-strongly convex. By definition of strong convexity:
$h_{t}(x_{t}^{0})-h_{t}(\hat{x}_{t}^{\star})\leq\nabla h_{t}(x_{t}^{0})^{\top}(x_{t}^{0}-\hat{x}_{t}^{\star})-\frac{\mu}{2}\left\|x_{t}^{0}-\hat{x}_{t}^{\star}\right\|^{2}.$ (B.8)
From the update rule and Lemma E.1 (since $\hat{x}_{t}^{\star}\in(1-\xi)\mathcal{X}$):
$\left\|x_{t+1}^{0}-\hat{x}_{t}^{\star}\right\|^{2}$ $=\left\|\mathbf{\Pi}_{(1-\xi)\mathcal{X}}(x_{t}^{0}-\eta_{t}(\nabla h_{t}(x_{t}^{0})+e_{t}))-\hat{x}_{t}^{\star}\right\|^{2}$
$\leq\left\|x_{t}^{0}-\eta_{t}(\nabla h_{t}(x_{t}^{0})+e_{t})-\hat{x}_{t}^{\star}\right\|^{2}$
$=\left\|x_{t}^{0}-\hat{x}_{t}^{\star}\right\|^{2}-2\eta_{t}(x_{t}^{0}-\hat{x}_{t}^{\star})^{\top}(\nabla h_{t}(x_{t}^{0})+e_{t})+\eta_{t}^{2}\left\|\nabla h_{t}(x_{t}^{0})+e_{t}\right\|^{2}.$
Rearranging this gives:
$(x_{t}^{0}-\hat{x}_{t}^{\star})^{\top}\nabla h_{t}(x_{t}^{0})\leq\frac{\left\|x_{t}^{0}-\hat{x}_{t}^{\star}\right\|^{2}-\left\|x_{t+1}^{0}-\hat{x}_{t}^{\star}\right\|^{2}}{2\eta_{t}}+(\hat{x}_{t}^{\star}-x_{t}^{0})^{\top}e_{t}+\frac{\eta_{t}}{2}\left\|\nabla h_{t}(x_{t}^{0})+e_{t}\right\|^{2}.$ (B.9)
Substitute (B.9) into (B.8):
$h_{t}(x_{t}^{0})-h_{t}(\hat{x}_{t}^{\star})$ $\leq\frac{\left\|x_{t}^{0}-\hat{x}_{t}^{\star}\right\|^{2}-\left\|x_{t+1}^{0}-\hat{x}_{t}^{\star}\right\|^{2}}{2\eta_{t}}+(\hat{x}_{t}^{\star}-x_{t}^{0})^{\top}e_{t}$
$\quad+\frac{\eta_{t}}{2}\left\|\nabla h_{t}(x_{t}^{0})+e_{t}\right\|^{2}-\frac{\mu}{2}\left\|x_{t}^{0}-\hat{x}_{t}^{\star}\right\|^{2}$
$\leq\frac{\left\|x_{t}^{0}-\hat{x}_{t}^{\star}\right\|^{2}-\left\|x_{t+1}^{0}-\hat{x}_{t}^{\star}\right\|^{2}}{2\eta_{t}}+(1-\xi)D\left\|e_{t}\right\|$
$\quad+\eta_{t}\left\|\nabla h_{t}(x_{t}^{0})\right\|^{2}+\eta_{t}\left\|e_{t}\right\|^{2}-\frac{\mu}{2}\left\|x_{t}^{0}-\hat{x}_{t}^{\star}\right\|^{2},$ (B.10)
where we used $\left\|a+b\right\|^{2}\leq 2\left\|a\right\|^{2}+2\left\|b\right\|^{2}$ and $\left\|\hat{x}_{t}^{\star}-x_{t}^{0}\right\|\leq(1-\xi)D$ from Remark 1.
Now we account for the dynamic comparator $\hat{x}_{t}^{\star}$. Use the fact that $\left\|a-b\right\|^{2}\geq\left\|a-c\right\|^{2}-2\langle a-b,c-b\rangle\geq\left\|a-c\right\|^{2}-2\left\|a-c\right\|\left\|c-b\right\|$:
$\left\|x_{t+1}^{0}-\hat{x}_{t}^{\star}\right\|^{2}$ $=\left\|x_{t+1}^{0}-\hat{x}_{t+1}^{\star}+\hat{x}_{t+1}^{\star}-\hat{x}_{t}^{\star}\right\|^{2}$
$\geq\left\|x_{t+1}^{0}-\hat{x}_{t+1}^{\star}\right\|^{2}-2\left\|x_{t+1}^{0}-\hat{x}_{t+1}^{\star}\right\|\left\|\hat{x}_{t+1}^{\star}-\hat{x}_{t}^{\star}\right\|$
$\geq\left\|x_{t+1}^{0}-\hat{x}_{t+1}^{\star}\right\|^{2}-2(1-\xi)D\left\|\hat{x}_{t}^{\star}-\hat{x}_{t+1}^{\star}\right\|.$ (B.11)
Substituting (B.11) into (B.10):
(specifically, $\left\|x_{t+1}^{0}-\hat{x}_{t}^{\star}\right\|^{2}\leq-\left\|x_{t+1}^{0}-\hat{x}_{t+1}^{\star}\right\|^{2}+2(1-\xi)D\left\|\hat{x}_{t}^{\star}-\hat{x}_{t+1}^{\star}\right\|$)
$h_{t}(x_{t}^{0})-h_{t}(\hat{x}_{t}^{\star})$ $\leq\frac{\left\|x_{t}^{0}-\hat{x}_{t}^{\star}\right\|^{2}-\left\|x_{t+1}^{0}-\hat{x}_{t+1}^{\star}\right\|^{2}}{2\eta_{t}}+\frac{(1-\xi)D\left\|\hat{x}_{t}^{\star}-\hat{x}_{t+1}^{\star}\right\|}{\eta_{t}}$
$\quad+(1-\xi)D\left\|e_{t}\right\|+\eta_{t}\left\|\nabla h_{t}(x_{t}^{0})\right\|^{2}+\eta_{t}\left\|e_{t}\right\|^{2}-\frac{\mu}{2}\left\|x_{t}^{0}-\hat{x}_{t}^{\star}\right\|^{2}.$
Now, we sum from $t=1$ to $T$. With the choice $\eta_{t}=1/(t\mu)$, we have $\frac{1}{2\eta_{t}}-\frac{\mu}{2}=\frac{t\mu}{2}-\frac{\mu}{2}=\frac{(t-1)\mu}{2}=\frac{1}{2\eta_{t-1}}$. The potential function terms form a telescoping sum that is non-positive:
$\sum_{t=1}^{T}\Big(\frac{\left\|x_{t}^{0}-\hat{x}_{t}^{\star}\right\|^{2}}{2\eta_{t}}$ $-\frac{\left\|x_{t+1}^{0}-\hat{x}_{t+1}^{\star}\right\|^{2}}{2\eta_{t}}-\frac{\mu}{2}\left\|x_{t}^{0}-\hat{x}_{t}^{\star}\right\|^{2}\Big)$
$=\sum_{t=1}^{T}\Big((\frac{1}{2\eta_{t}}-\frac{\mu}{2})\left\|x_{t}^{0}-\hat{x}_{t}^{\star}\right\|^{2}-\frac{\left\|x_{t+1}^{0}-\hat{x}_{t+1}^{\star}\right\|^{2}}{2\eta_{t}}\Big)$
$=\sum_{t=1}^{T}\Big(\frac{\left\|x_{t}^{0}-\hat{x}_{t}^{\star}\right\|^{2}}{2\eta_{t-1}}-\frac{\left\|x_{t+1}^{0}-\hat{x}_{t+1}^{\star}\right\|^{2}}{2\eta_{t}}\Big)$
$=\frac{\left\|x_{1}^{0}-x_{1,\xi}^{\star}\right\|^{2}}{2\eta_{0}}-\frac{\left\|x_{T+1}^{0}-\hat{x}_{t+1}^{\star}\right\|^{2}}{2\eta_{T}}\leq 0,$
where we set $1/\eta_{0}=0$ and $\hat{x}_{t+1}^{\star}=\hat{x}_{t}^{\star}$ since $t=T+1$ is vacuous in our analysis.
Thus, we arrive at the inequality:
$\sum_{t=1}^{T}\big(h_{t}(x_{t}^{0})-h_{t}(\hat{x}_{t}^{\star})\big)$ $\leq\sum_{t=1}^{T-1}\frac{(1-\xi)D\left\|\hat{x}_{t}^{\star}-\hat{x}_{t+1}^{\star}\right\|}{\eta_{t}}+\sum_{t=1}^{T}(1-\xi)D\left\|e_{t}\right\|$
$\quad+\sum_{t=1}^{T}\eta_{t}\left\|\nabla h_{t}(x_{t}^{0})\right\|^{2}+\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|^{2}.$ (B.12)
Step 5: Bound the Terms in the Inequality (B.12). We now take the expectation of (B.12) and use the bounds derived in Appendix C (which we cite here):
  • $\mathbb{E}\Big[\sum_{t=1}^{T-1}\frac{(1-\xi)D\left\|\hat{x}_{t}^{\star}-\hat{x}_{t+1}^{\star}\right\|}{\eta_{t}}\Big]\leq\frac{\sqrt{2}(1-\xi)D}{\mu}\mathbb{E}\Big[\sum_{t=1}^{T-1}\left\|\hat{x}_{t}^{\star}-\hat{x}_{t+1}^{\star}\right\|\Big]=\frac{\sqrt{2}(1-\xi)^{2}D}{\mu}V_{T}$ (from (C.1)).
  • $\mathbb{E}\Big[\sum_{t=1}^{T}\eta_{t}\left\|\nabla h_{t}(x_{t}^{0})\right\|^{2}\Big]\leq\frac{hG^{2}}{2(2h-1)\mu}(1+\ln T)$ (from (C.3)).
  • $\mathbb{E}\Big[\sum_{t=1}^{T}(1-\xi)D\left\|e_{t}\right\|\Big]\leq(\frac{(1-\xi)D}{\delta(2(2h-1))^{1/4}}+\frac{\sqrt{2}h^{2}G(1-\xi)D}{2\delta(2(2h-1))^{1/2}})\sum_{t=1}^{T}\varphi_{t}+\frac{h^{3}G^{2}(1-\xi)D}{\delta(2(2h-1))^{1/2}}(1+\ln{T})$ (from (C.11)).
  • $\mathbb{E}\Big[\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|^{2}\Big]\leq\frac{2}{\delta^{2}\mu\sqrt{2h-1}}\sum_{t=1}^{T}\varphi_{t}^{2}+\frac{4G^{2}h^{2}}{\mu(2h-1)}(1+\ln T)$ (from (C.8)).
Summing these bounds gives the bound for $\mathbb{E}\Big[\sum_{t=1}^{T}\big(\hat{f}_{t}^{c}(x_{t}^{0})-\hat{f}_{t}^{c}(\hat{x}_{t}^{\star})\big)\Big]$ (note that $\mathbb{E}[h_{t}(x_{t}^{0})]=\mathbb{E}[\hat{f}_{t}^{c}(x_{t}^{0})]$ and $\mathbb{E}[h_{t}(\hat{x}_{t}^{\star})]=\mathbb{E}[\hat{f}_{t}^{c}(\hat{x}_{t}^{\star})]$ which have been argued before):
$\mathbb{E}\Big[\sum_{t=1}^{T}\big(\hat{f}_{t}^{c}(x_{t}^{0})$ $-\hat{f}_{t}^{c}(\hat{x}_{t}^{\star})\big)\Big]$
$\leq\frac{\sqrt{2}(1-\xi)^{2}D}{\mu}V_{T}+\Big(\frac{hG^{2}}{2(2h-1)\mu}+\frac{h^{3}G^{2}(1-\xi)D}{\delta(2(2h-1))^{1/2}}+\frac{4G^{2}h^{2}}{\mu(2h-1)}\Big)(1+\ln T)$
$\quad+\frac{2}{\delta^{2}\mu\sqrt{2h-1}}\sum_{t=1}^{T}\varphi_{t}^{2}+\Big(\frac{(1-\xi)D}{\delta(2(2h-1))^{1/4}}+\frac{\sqrt{2}h^{2}G(1-\xi)D}{2\delta(2(2h-1))^{1/2}}\Big)\sum_{t=1}^{T}\varphi_{t}$
$\leq\frac{\sqrt{2}(1-\xi)^{2}D}{\mu}V_{T}+\Big(\frac{8G^{2}h^{2}+hG^{2}}{2\mu(2h-1)}+\frac{h^{3}G^{2}(1-\xi)D}{\delta(2(2h-1))^{1/2}}\Big)(1+\ln T)$
$\quad+\frac{2}{\delta^{2}\mu\sqrt{2h-1}}\sum_{t=1}^{T}\varphi_{t}^{2}+\Big(\frac{(1-\xi)D}{\delta(2(2h-1))^{1/4}}+\frac{\sqrt{2}h^{2}G(1-\xi)D}{2\delta(2(2h-1))^{1/2}}\Big)\sum_{t=1}^{T}\varphi_{t}.$ (B.13)
Combine All Pieces for Algorithmic Regret. We now combine the results from Steps 2, 3, and 5:
$\mathbb{E}\Big[C_{T}(\bm{x}^{0})-C_{T}(\hat{\bm{x}}^{\star})\Big]$ $\leq\mathbb{E}\Big[\sum_{t=1}^{T}\Big(f_{t}^{c}(x_{t}^{0})-f_{t}^{c}(\hat{x}_{t}^{\star})\Big)\Big]+\mathbb{E}\Big[\Big|\sum_{t=1}^{T}f_{t}(x_{t-h+1:t}^{0})-f_{t}^{c}(x_{t}^{0})\Big|\Big]$
$\qquad\qquad+\mathbb{E}\Big[\Big|\sum_{t=1}^{T}f_{t}(\hat{x}_{t-h+1:t}^{\star})-f_{t}^{c}(\hat{x}_{t}^{\star})\Big|\Big]$
[…]
$\qquad\qquad+\mathbb{E}\Big[\Big|\sum_{t=1}^{T}f_{t}(\hat{x}_{t-h+1:t}^{\star})-f_{t}^{c}(\hat{x}_{t}^{\star})\Big|\Big].$
Plugging everything in:
[…]
$\leq\Big[\frac{\sqrt{2}(1-\xi)^{2}D}{\mu}V_{T}+\Big(\frac{8G^{2}h^{2}+hG^{2}}{2\mu(2h-1)}+\frac{h^{3}G^{2}(1-\xi)D}{\delta(2(2h-1))^{1/2}}\Big)(1+\ln T)$
$\quad+\frac{2}{\delta^{2}\mu\sqrt{2h-1}}\sum_{t=1}^{T}\varphi_{t}^{2}+\Big(\frac{(1-\xi)D}{\delta(2(2h-1))^{1/4}}+\frac{\sqrt{2}h^{2}G(1-\xi)D}{2\delta(2(2h-1))^{1/2}}\Big)\sum_{t=1}^{T}\varphi_{t}\Big]$
$\quad+T\delta^{2}\beta d$
$\quad+\Big[\frac{3G^{2}h^{3}}{\mu(2(2h-1))^{1/2}}(1+\ln T)+\frac{\sqrt{2}Gh^{2}}{\delta\mu(2(2h-1))^{1/4}}\sum_{t=1}^{T}\varphi_{t}\Big]$
$\quad+(1-\xi)Gh^{2}V_{T}.$
Rearranging terms gives the bound for the Algorithmic Regret:
$\mathbb{E}\Big[C_{T}(\bm{x}^{0})$ $-C_{T}(\hat{\bm{x}}^{\star})\Big]\leq\Big(\frac{\sqrt{2}(1-\xi)^{2}D}{\mu}+(1-\xi)Gh^{2}\Big)V_{T}+T\delta^{2}\beta d$
$\quad+\Big(\frac{8G^{2}h^{2}+hG^{2}}{2\mu(2h-1)}+\frac{h^{3}G^{2}(1-\xi)D}{\delta(2(2h-1))^{1/2}}+\frac{3G^{2}h^{3}}{\mu(2(2h-1))^{1/2}}\Big)(1+\ln T)$
$\quad+\frac{2}{\delta^{2}\mu\sqrt{2h-1}}\sum_{t=1}^{T}\varphi_{t}^{2}$
$\quad+\Big(\frac{(1-\xi)D}{\delta(2(2h-1))^{1/4}}+\frac{\sqrt{2}h^{2}G(1-\xi)D}{2\delta(2(2h-1))^{1/2}}+\frac{\sqrt{2}Gh^{2}}{\delta\mu(2(2h-1))^{1/4}}\Big)\sum_{t=1}^{T}\varphi_{t}$
$\leq\Big(\frac{\sqrt{2}(1-\xi)D}{\mu}+Gh^{2}\Big)V_{T,\xi}+T\delta^{2}\beta d$
$\quad+\Big(\frac{8G^{2}h^{2}+hG^{2}}{2\mu(2h-1)}+\frac{h^{3}G^{2}(1-\xi)D}{\delta(2(2h-1))^{1/2}}+\frac{3G^{2}h^{3}}{\mu(2(2h-1))^{1/2}}\Big)(1+\ln T)$
$\quad+\frac{2}{\delta^{2}\mu\sqrt{2h-1}}\sum_{t=1}^{T}\varphi_{t}^{2}+\Big(\frac{\sqrt{2}h^{2}G(1-\xi)D}{2\delta(2(2h-1))^{1/2}}+\frac{\sqrt{2}Gh^{2}+(1-\xi)D\mu}{\delta\mu(2(2h-1))^{1/4}}\Big)\sum_{t=1}^{T}\varphi_{t}.$ (B.14)
Part 2: Bounding the Approximation Error. This part bounds $C_{T}(\hat{\bm{x}}^{\star})-C_{T}(\bm{x}^{\star})$. Since $C_{T}(\cdot)$ is $G\sqrt{Th}$-Lipschitz (from Lemma 1(b)):
$|C_{T}(\hat{\bm{x}}^{\star})-C_{T}(\bm{x}^{\star})|$ $\leq G\sqrt{Th}\left\|\hat{\bm{x}}^{\star}-\bm{x}^{\star}\right\|$
$=G\sqrt{Th}\sqrt{\sum_{t=1}^{T}\left\|(1-\xi)x_{t}^{\star}-x_{t}^{\star}\right\|^{2}}$
$=G\sqrt{Th}\sqrt{\sum_{t=1}^{T}\xi^{2}\left\|x_{t}^{\star}\right\|^{2}}$
$\leq G\sqrt{Th}\cdot\xi\sqrt{TD^{2}}\quad(\text{since }\left\|x_{t}^{\star}\right\|\leq D\text{ as }x_{t}^{\star}\in\mathcal{X})$
$=\xi GT\sqrt{h}D.$ (B.15)
Final Bound. Combining the bounds for Part 1 (Eq. (B.14)) and Part 2 (Eq. (B.15)), we get the final bound for the total initialization regret $\mathbb{E}[R(\mathcal{A}_{\text{init}})]$:
$\mathbb{E}[R(\mathcal{A}_{\text{init}})]$ $=\mathbb{E}\Big[C_{T}(\bm{x}^{0})-C_{T}(\hat{\bm{x}}^{\star})\Big]+\Big(C_{T}(\hat{\bm{x}}^{\star})-C_{T}(\bm{x}^{\star})\Big)$
$\leq\Big(\frac{\sqrt{2}(1-\xi)^{2}D}{\mu}+(1-\xi)Gh^{2}\Big)V_{T}+T\delta^{2}\beta d$
$\quad+\Big(\frac{8G^{2}h^{2}+hG^{2}}{2\mu(2h-1)}+\frac{h^{3}G^{2}(1-\xi)D}{\delta(2(2h-1))^{1/2}}+\frac{3G^{2}h^{3}}{\mu(2(2h-1))^{1/2}}\Big)(1+\ln T)$
$\quad+\frac{2}{\delta^{2}\mu\sqrt{2h-1}}\sum_{t=1}^{T}\varphi_{t}^{2}+\Big(\frac{\sqrt{2}h^{2}G(1-\xi)D}{2\delta(2(2h-1))^{1/2}}+\frac{\sqrt{2}Gh^{2}+(1-\xi)D\mu}{\delta\mu(2(2h-1))^{1/4}}\Big)\sum_{t=1}^{T}\varphi_{t}$
$\quad+\xi GT\sqrt{h}D,$
which completes the proof of Theorem 1.
B.2 Proof of Theorem 2
For our analysis in this proof, denote
$\widetilde{\bm{g}}^{j}_{s}=\sum_{k=s}^{s+h-1}\frac{f_{k}(\bar{x}^{j}_{k-h+1:k})-f_{k}(\tilde{x}^{j}_{k-h+1:k})}{2\delta^{\prime}}u_{s}^{j}.$ (B.16)
We then have from Lemma 2 that $\mathbb{E}_{u_{s}^{j}}[\widetilde{\bm{g}}_{s}^{j}]=\partial\hat{C}_{T}(\bm{x}^{j})/\partial x_{s}$ for all $s\in[T]$ and all $j\in\{0,\dots,K-1\}$ and thus $\nabla\hat{C}_{T}(\bm{x}^{j})=\mathbb{E}_{\bm{u}^{j}}[\widetilde{\bm{g}}^{j}]$, where $\bm{u}^{j}=(u_{1}^{j},\dots,u_{T}^{j})\sim\widetilde{\mathcal{N}}(0,I_{dT})$ and $\widetilde{\bm{g}}^{j}=(\widetilde{\bm{g}}_{1}^{j},\dots,\widetilde{\bm{g}}_{T}^{j})$.
Let $\bm{x}_{\xi^{\prime}}^{\star}\in\operatorname*{arg\,min}_{\bm{x}\in((1-\xi^{\prime})\mathcal{X})^{T}}C_{T}(\bm{x})$. To proceed, we first establish the convergence to the proxy optimal solution $\bm{x}_{\xi^{\prime}}^{\star}$. Define the potential function:
$W_{j}=(1+\gamma)^{j}(C_{T}(\bm{x}^{j})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})).$ (B.17)
We then have
$W_{j+1}-W_{j}$ $=(1+\gamma)^{j}\big((\gamma+1)(C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j}))+\gamma(C_{T}(\bm{x}^{j})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star}))\big)$
$=(1+\gamma)^{j}\big(C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j})+\gamma(C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star}))\big).$ (B.18)
To upper bound $W_{j+1}-W_{j}$, we first prove in Appendix D that the following inequality holds for all $j\in\{0,\dots,K-1\}$ and all $\bm{y}\in((1-\xi^{\prime})\mathcal{X})^{T}$:
$C_{T}(\bm{x}^{j+1})-C_{T}(\bm{y})\leq\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j}-\bm{y}\rangle-\frac{\beta^{\prime}}{2}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}-\frac{\mu}{2}\left\|\bm{x}^{j}-\bm{y}\right\|^{2}\\ +\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle.$ (B.19)
Letting $\bm{y}=\bm{x}^{j}$ (which holds for $j\geq 1$, and we analyze $j=0$ separately if needed, though the proof structure allows it) and $\bm{y}=\bm{x}_{\xi^{\prime}}^{\star}$ in (B.19) (note that $\bm{x}_{\xi^{\prime}}^{\star}\in((1-\xi^{\prime})\mathcal{X})^{T}$), we respectively obtain the following:
$C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j})$ $\leq-\frac{\beta^{\prime}}{2}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{x}^{j}\rangle,$ (B.20)
$C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})$ $\leq\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j}-\bm{x}_{\xi^{\prime}}^{\star}\rangle-\frac{\beta^{\prime}}{2}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}-\frac{\mu}{2}\left\|\bm{x}^{j}-\bm{x}_{\xi^{\prime}}^{\star}\right\|^{2}$
$\qquad\qquad\qquad\qquad\qquad\qquad+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{x}_{\xi^{\prime}}^{\star}\rangle.$ (B.21)
Moreover, for any $\bm{y}=(y_{1},\dots,y_{T})$ with $y_{t}\in(1-\xi^{\prime})\mathcal{X}$ for all $t\in[T]$, we have the following relation proved in Appendix D:
$\mathbb{E}\big[\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big]\leq\varepsilon.$ (B.22)
Now, substituting (B.20)-(B.21) with the upper bound $\varepsilon$ into (B.18), we show in Appendix D that
$\mathbb{E}_{\bm{u}^{j}}[W_{j+1}-W_{j}]\leq(1+\gamma)^{j+1}\varepsilon.$ (B.23)
From (B.17) and (B.23), we get
$\mathbb{E}_{\bm{u}^{j}}\big[C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]\leq\frac{1}{1+\gamma}\big(C_{T}(\bm{x}^{j})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big)+\varepsilon\quad\forall j\in\{0,\dots,K-1\},$
$\Rightarrow$ $\mathbb{E}_{\mathcal{U}_{K-1}}\big[C_{T}(\bm{x}^{K})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]\leq\Big(\frac{1}{1+\gamma}\Big)^{K}\big(C_{T}(\bm{x}^{0})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big)+\sum_{j=0}^{K-1}\Big(\frac{1}{1+\gamma}\Big)^{j}\varepsilon$
$\Rightarrow$ $\mathbb{E}_{\mathcal{U}_{K-1}}\big[C_{T}(\bm{x}^{K})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]\leq\Big(\frac{1}{1+\gamma}\Big)^{K}\mathbb{E}\big[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]+\frac{\varepsilon}{\gamma},$
where $\mathcal{U}_{K-1}\triangleq(\bm{u}^{0},\dots,\bm{u}^{K-1})$ encapsulates all the randomness of Algorithm 1.
Overall Bound on the Regret $R(\mathcal{A})$. We now combine this result with the initialization regret to bound the total dynamic regret against the true optimal $\bm{x}^{\star}$. We decompose the regret as:
$\mathbb{E}[R(\mathcal{A})]=\mathbb{E}\big[C_{T}(\bm{x}^{K})-C_{T}(\bm{x}^{\star})\big]=\underbrace{\mathbb{E}\big[C_{T}(\bm{x}^{K})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]}_{\text{Term A}}+\underbrace{\Big(C_{T}(\bm{x}_{\xi^{\prime}}^{\star})-C_{T}(\bm{x}^{\star})\Big)}_{\text{Term B}}.$
Bounding Term B (Approximation Error): Since $\bm{x}^{\star}$ is in $\mathcal{X}^{T}$ and contains the origin, $(1-\xi^{\prime})\bm{x}^{\star}\in((1-\xi^{\prime})\mathcal{X})^{T}$. The approximation error is bounded by:
Term B $=C_{T}(\bm{x}_{\xi^{\prime}}^{\star})-C_{T}(\bm{x}^{\star})$
$\overset{\text{(a)}}{\leq}C_{T}((1-\xi^{\prime})\bm{x}^{\star})-C_{T}(\bm{x}^{\star})$
$\overset{\text{(b)}}{\leq}\xi^{\prime}GT\sqrt{h}D,$
where (a) is due to $\bm{x}_{\xi^{\prime}}^{\star}\in\operatorname*{arg\,min}_{\bm{x}\in((1-\xi^{\prime})\mathcal{X})^{T}}C_{T}(\bm{x})$; (b) follows the same proof logic as (B.15).
Bounding Term A and Combining: We substitute the bound for Term A derived before:
$\text{Term A}\leq\Big(\frac{1}{1+\gamma}\Big)^{K}\mathbb{E}\big[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]+\frac{\varepsilon}{\gamma}.$
We bound the initial error term $\mathbb{E}\big[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]$ by relating it to the total initialization regret. Since $C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\geq C_{T}(\bm{x}^{\star})$, we have:
$\mathbb{E}\big[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star})\big]$ $=\mathbb{E}\big[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})\big]-\big(C_{T}(\bm{x}_{\xi^{\prime}}^{\star})-C_{T}(\bm{x}^{\star})\big)$
$\leq\mathbb{E}[R(\mathcal{A}_{\text{init}})].$
Substituting these back into the regret decomposition, we obtain the final result:
$\mathbb{E}[R(\mathcal{A})]\leq\Big(\frac{1}{1+\gamma}\Big)^{K}\mathbb{E}[R(\mathcal{A}_{\text{init}})]+\frac{\varepsilon}{\gamma}+\xi^{\prime}GT\sqrt{h}D.$
This completes the proof of Theorem 2.
Appendix C Omitted Details for the Proof of Theorem 1
C.1 Proof of the bounds of the four terms in (B.12)
First, note that
$\sum_{t=1}^{T-1}\frac{(1-\xi)D\left\|\hat{x}_{t}^{\star}-\hat{x}_{t+1}^{\star}\right\|}{\eta_{t}}\overset{(a)}{\leq}\frac{(1-\xi)D}{\mu}\sqrt{(\sum_{t=1}^{T-1}\left\|\hat{x}_{t}^{\star}-\hat{x}_{t+1}^{\star}\right\|^{2})(\sum_{t=1}^{T-1}\frac{1}{t^{2}})}$
$\overset{(b)}{\leq}$ $\frac{\sqrt{2}(1-\xi)D}{\mu}\sqrt{\sum_{t=1}^{T-1}\left\|\hat{x}_{t}^{\star}-\hat{x}_{t+1}^{\star}\right\|^{2}}$
$\leq$ $\frac{\sqrt{2}(1-\xi)D}{\mu}\sum_{t=1}^{T-1}\left\|\hat{x}_{t}^{\star}-\hat{x}_{t+1}^{\star}\right\|=\frac{\sqrt{2}(1-\xi)^{2}D}{\mu}V_{T},$ (C.1)
where (a) follows from the CS-inequality and (b) follows from the fact
$\sum_{t=1}^{T-1}\frac{1}{t^{2}}\leq 1+\sum_{t=2}^{T-1}\frac{1}{t(t-1)}=1+\sum_{t=2}^{T-1}(\frac{1}{t-1}-\frac{1}{t})\leq 2.$ (C.2)
Next, we have from (B.2) that
$\sum_{t=1}^{T}\eta_{t}\left\|\nabla h_{t}(x_{t}^{0})\right\|^{2}$ $\leq\frac{hG^{2}}{\mu}\sum_{t=1}^{T}\frac{1}{t}\left\|u_{t}^{0}\right\|^{4}$
$\overset{(a)}{\leq}\frac{hG^{2}}{2(2h-1)\mu}\sum_{t=1}^{T}\frac{1}{t}$
$\overset{(b)}{\leq}\frac{hG^{2}}{2(2h-1)\mu}(1+\ln T),$ (C.3)
where (a) follows from the fact that $\lVert u_{t}^{0}\rVert\leq 1/(2(2h-1))^{1/4}$, and (b) uses the fact $\sum_{t=1}^{T}\frac{1}{t}<1+\ln T$.
To proceed, we upper bound $\left\|e_{t}\right\|$ for all $t\in[T]$, where $e_{t}=e_{t}^{m}+e_{t}^{c}$. We first have from (B.3) that
$\left\|e_{t}^{m}\right\|\leq\frac{\left\|u^{0}\right\|}{\delta}\varphi_{t}\leq\frac{\varphi_{t}}{\delta(2(2h-1))^{1/4}}.$ (C.4)
To upper bound $\left\|e_{t}^{c}\right\|$, we have from (B.4) that
$\left\|e_{t}^{c}\right\|\leq\frac{1}{2\delta}|f_{t}(\bar{x}_{t-h+1:t}^{0})-f_{t}^{c}(\bar{x}_{t}^{0})|\left\|u^{0}\right\|+\frac{1}{2\delta}|f_{t}(\tilde{x}_{t-h+1:t}^{0})-f_{t}^{c}(\tilde{x}_{t}^{0})|\left\|u^{0}\right\|.$ (C.5)
We may further upper bound bound $\left\|e_{t}^{c}\right\|$ as
$\left\|e_{t}^{c}\right\|$ $\leq\frac{1}{2\delta}|f_{t}(\bar{x}_{t-h+1:t}^{0})-f_{t}(\tilde{x}_{t-h+1:t}^{0})|\left\|u^{0}\right\|+\frac{1}{2\delta}|f_{t}^{c}(\bar{x}_{t}^{0})-f_{t}^{c}(\tilde{x}_{t}^{0})|\left\|u^{0}\right\|$
$\overset{(a)}{\leq}\frac{G}{2\delta}\sum_{j=0}^{h-1}\left\|\bar{x}_{t-j}^{0}-\tilde{x}_{t-j}^{0}\right\|\left\|u^{0}\right\|+\frac{Gh}{2\delta}\left\|\bar{x}_{t}^{0}-\tilde{x}_{t}^{0}\right\|\left\|u^{0}\right\|$
$\overset{(b)}{\leq}G\sum_{j=0}^{h-1}\left\|u^{0}\right\|\left\|u^{0}\right\|+Gh\left\|u^{0}\right\|^{2}$
$\leq\frac{2Gh}{\sqrt{2(2h-1)}},$ (C.6)
where (a) again follows from Assumption 1(c), and (b) uses the relation between $\bar{x}_{t}^{0}$ and $\tilde{x}_{t}^{0}$. The above inequality further implies
$\left\|e_{t}^{c}\right\|^{2}\leq\frac{2G^{2}h^{2}}{(2h-1)}.$ (C.7)
We then upper bound $\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|^{2}$ in (B.12). We have
$\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|^{2}$ $\leq\sum_{t=1}^{T}\eta_{t}(2\left\|e_{t}^{m}\right\|^{2}+2\left\|e_{t}^{c}\right\|^{2})$
$\overset{(a)}{\leq}\sum_{t=1}^{T}\eta_{t}\frac{2\varphi_{t}^{2}}{\delta^{2}\sqrt{2(2h-1)}}+4\sum_{t=1}^{T}\eta_{t}\frac{G^{2}h^{2}}{(2h-1)}$
$\overset{(b)}{\leq}\frac{2}{\delta^{2}\mu\sqrt{2(2h-1)}}\sqrt{(\sum_{t=1}^{T}\varphi_{t}^{4})(\sum_{t=1}^{T}\frac{1}{t^{2}})}+\frac{4G^{2}h^{2}}{\mu(2h-1)}\sum_{t=1}^{T}\frac{1}{t}$
$\overset{(c)}{\leq}\frac{2}{\delta^{2}\mu\sqrt{2h-1}}\sum_{t=1}^{T}\varphi_{t}^{2}+\frac{4G^{2}h^{2}}{\mu(2h-1)}(1+\ln T),$ (C.8)
where (a) follows from (C.4) and (C.7), (b) follows from the choice $\eta=\frac{1}{t\mu}$ and the CS-inequality. To obtain (c), we used the fact $\sum_{i=1}^{n}a_{i}^{2}\leq(\sum_{i=1}^{n}a_{i})^{2}$ for $a_{i}\in\mathbb{R}^{+}$, $\sum_{t=1}^{T}1/t^{2}\leq 2$ from (C.2) and $\sum_{t=1}^{T}\frac{1}{t}\leq 1+\ln T$. Similarly, we can upper bound $\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|$ as
$\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|$ $\leq\sum_{t=1}^{T}\eta_{t}(\left\|e_{t}^{m}\right\|+\left\|e_{t}^{c}\right\|)$
$\leq\sum_{t=1}^{T}\frac{1}{t\mu}\frac{\varphi_{t}}{\delta(2(2h-1))^{\frac{1}{4}}}+\sum_{t=1}^{T}\frac{1}{t\mu}\frac{\sqrt{2}Gh}{\sqrt[]{2h-1}}$
$\leq\frac{\sqrt{2}}{\delta\mu(2(2h-1))^{1/4}}\sum_{t=1}^{T}\varphi_{t}+\frac{\sqrt{2}Gh}{\mu\sqrt{2h-1}}(1+\ln T).$ (C.9)
Finally, we upper bound $\sum_{t=1}^{T}D\left\|e_{t}\right\|$ in (B.12). We first note from Assumption 1(c) that
$\quad|f_{t}(\bar{x}_{t-h+1:t}^{0})-f_{t}^{c}(\bar{x}_{t}^{0})|$ $\leq G\sum_{j=1}^{h-1}\left\|\bar{x}_{t}^{0}-\bar{x}_{t-j}^{0}\right\|$
$\leq G\sum_{j=1}^{h-1}\sum_{l=1}^{j}\left\|\bar{x}_{t-l+1}^{0}-\bar{x}_{t-l}^{0}\right\|$
$\leq G\sum_{j=1}^{h-1}\sum_{l=1}^{j}\left\|x_{t-l+1}^{0}+\delta u^{0}-x_{t-l}^{0}-\delta u^{0}\right\|$
$\overset{(a)}{\leq}G\sum_{j=1}^{h-1}\sum_{l=1}^{j}\big(\eta_{t-l}\left\|\nabla h_{t-l}(x_{t-l}^{0})\right\|+\eta_{t-l}\left\|e_{t-l}\right\|\big)$
[…]
where (a) uses the relation $x_{t-l+1}^{0}=\Pi_{(1-\xi)\mathcal{X}}(x_{t-l}^{0}-\eta_{t-l}(\nabla h_{t-l}(x_{t-l}^{0})+e_{t}))$ and Lemma E.1 in Appendix E.
Similarly, one can show that
$|f_{t}(\tilde{x}_{t-h+1:t}^{0})-f_{t}^{c}(\tilde{x}_{t}^{0})|\leq\sqrt{h}G^{2}\sum_{j=1}^{h-1}\sum_{l=1}^{j}\eta_{t-l}\left\|u^{0}\right\|^{2}+G\sum_{j=1}^{h-1}\sum_{l=1}^{j}\eta_{t-l}\left\|e_{t-l}\right\|.$
We then have from (C.5) that
$\left\|e^{c}_{t}\right\|\leq\frac{\sqrt{h}G^{2}}{\delta}\sum_{j=1}^{h-1}\sum_{l=1}^{j}\eta_{t-l}\left\|u^{0}\right\|^{3}+\frac{G}{\delta}\sum_{j=1}^{h-1}\sum_{l=1}^{j}\left\|e_{t-l}\right\|\left\|u^{0}\right\|$
$\quad\quad\leq\frac{\sqrt{h}h^{2}G^{2}}{2\delta}\sum_{t=1}^{T}\eta_{t}\left\|u^{0}\right\|^{3}+\frac{h^{2}G^{2}}{2\delta}\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|\left\|u^{0}\right\|,$
[…]
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{\sqrt{2}Gh}{\mu\sqrt{2h-1}}(1+\ln T)\big]$
$\quad\quad\quad\leq\frac{\sqrt{2}h^{2}G}{2\delta(2(2h-1))^{1/2}}\sum_{t=1}^{T}\varphi_{t}+\frac{h^{3}G^{2}}{\delta(2(2h-1))^{1/2}}(1+\ln{T}).$ (C.10)
Combining (C.4) and (C.10) and summing from $t=1$ to $t=T$, we get
$\quad\mathbb{E}\Big[\sum_{t=1}^{T}D\left\|e_{t}\right\|\Big]$
$\leq(\frac{D}{\delta(2(2h-1))^{1/4}}+\frac{\sqrt{2}h^{2}GD}{2\delta(2(2h-1))^{1/2}})\sum_{t=1}^{T}\varphi_{t}+\frac{h^{3}G^{2}D}{\delta(2(2h-1))^{1/2}}(1+\ln{T}).$ (C.11)
[…]
Appendix D Omitted Details in the Proof of Theorem 2
D.1 Proof of (B.19)
We first express $C_{T}(\bm{x}^{j+1})-C_{T}(\bm{y})=C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j})+C_{T}(\bm{x}^{j})-C_{T}(\bm{y})$. From $\beta^{\prime}$-smoothness of $C_{T}(\cdot)$, we have
$C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j})\leq\langle\nabla C_{T}(\bm{x}^{j}),\bm{x}^{j+1}-\bm{x}^{j}\rangle+\frac{\beta^{\prime}}{2}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}.$
From the $\mu$-strongly convexity of $C_{T}(\cdot)$, we have
$C_{T}(\bm{x}^{j})-C_{T}(\bm{y})\leq\langle\nabla C_{T}(\bm{x}^{j}),\bm{x}^{j}-\bm{y}\rangle-\frac{\mu}{2}\left\|\bm{x}^{j}-\bm{y}\right\|^{2}.$
Summing the above two inequalities yields
$C_{T}(\bm{x}^{j+1})-C_{T}(\bm{y})\leq\langle\nabla C_{T}(\bm{x}^{j}),\bm{x}^{j+1}-\bm{y}\rangle+\frac{\beta^{\prime}}{2}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}-\frac{\mu}{2}\left\|\bm{x}^{j}-\bm{y}\right\|^{2}.$ (D.1)
As we argued in the main body, denoting $\bm{z}^{j+1}=\bm{x}^{j}-\alpha\bm{g}^{j}$ with $\alpha=1/\beta^{\prime}$, we get that
$\langle\nabla C_{T}(\bm{x}^{j}),\bm{x}^{j+1}-\bm{y}\rangle$ $=\langle\bm{g}^{j}+\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle$
$=\langle\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle$
$=\beta^{\prime}\langle\bm{x}^{j}-\bm{z}^{j+1},\bm{x}^{j+1}-\bm{y}\rangle+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle$
$\leq\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j+1}-\bm{y}\rangle+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle$
$=\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j+1}-\bm{x}^{j}+\bm{x}^{j}-\bm{y}\rangle+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle$
$=-\beta^{\prime}\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|^{2}+\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j}-\bm{y}\rangle+\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle,$
[…]
D.2 Proof of (B.22)
For any $\bm{y}=(y_{1},\dots,y_{T})$ with $y_{t}\in(1-\xi^{\prime})\mathcal{X}$ for all $t\in[T]$, we have
$\mathbb{E}_{\bm{u}^{j}}\big[\langle\nabla C_{T}(\bm{x}^{j})-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big]$ $=\mathbb{E}_{\bm{u}^{j}}\big[\langle\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}+\widetilde{\bm{g}}^{j}-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big],$
$=\underbrace{\mathbb{E}_{\bm{u}^{j}}\big[\langle\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big]}_{\text{(i)}}+\underbrace{\mathbb{E}_{\bm{u}^{j}}\big[\langle\widetilde{\bm{g}}^{j}-\bm{g}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big]}_{\text{(ii)}}.$
We first upper bound (ii). Recalling the definition of $\bm{g}^{j}$ (resp., $\widetilde{\bm{g}}^{j}$) given in Algorithm 1 (resp., (B.16)), we get that
$\mathbb{E}_{{u}^{j}_{s}}[\bm{g}^{j}_{s}-\widetilde{\bm{g}}^{j}_{s}]=\mathbb{E}_{u^{j}_{s}}\Big[\sum_{k=s}^{s+h-1}\frac{l_{k}(\bar{x}_{k-h+1:k})-f_{k}(\bar{x}_{k-h+1:k})+f_{k}(\tilde{x}_{k-h+1:k})-l_{k}(\tilde{x}_{k-h+1:k})}{2\delta^{\prime}}u^{j}_{s}\Big],$
for all $s\in[T]$. By Assumption 3 and the fact that $\|u_{s}^{j}\|\leq 1/2(2h-1))^{1/4}$, we obtain
$\mathbb{E}_{\bm{u}^{j}}\big[\lVert\bm{g}_{s}^{j}-\widetilde{\bm{g}}_{s}^{j}\rVert\big]\leq\frac{\sum_{k=s}^{s+h-1}\varphi_{k}}{\delta^{\prime}(2(2h-1))^{1/4}}\ \forall s\in[T],$
$\Rightarrow$ $\mathbb{E}_{\bm{u}^{j}}\big[\lVert\bm{g}^{j}-\widetilde{\bm{g}}^{j}\rVert\big]\leq\frac{\sum_{s=1}^{T}\sum_{k=s}^{s+h-1}\varphi_{k}}{\delta^{\prime}(2(2h-1))^{1/4}}\leq\frac{h\sum_{t=1}^{T}\varphi_{t}}{\delta^{\prime}(2(2h-1))^{1/4}}.$
Thus, we can upper bound (ii) as
$\text{(ii)}\leq\mathbb{E}_{\bm{u}^{j}}\big[\|\bm{g}^{j}-\widetilde{\bm{g}}^{j}\|\|\bm{x}^{j+1}-\bm{y}\|\big]\leq\frac{(1-\xi^{\prime})Dh\sum_{t=1}^{T}\varphi_{t}}{\delta^{\prime}(2(2h-1))^{1/4}},$
where the second inequality follows from Remark 1 (noting $\bm{x}^{j+1},\bm{y}\in((1-\xi^{\prime})\mathcal{X})^{T}$).
We then upper bound (i). Note that
$\mathbb{E}_{\bm{u}^{j}}\big[\langle\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big]$
$\leq\mathbb{E}_{\bm{u}^{j}}\big[\|\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}\|\|\bm{x}^{j+1}-\bm{y}\|\big]$
$\overset{(a)}{\leq}(1-\xi^{\prime})D\mathbb{E}_{\bm{u}^{j}}\big[\|\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}\|\big]$
$\leq(1-\xi^{\prime})D\mathbb{E}_{\bm{u}^{j}}\Big[\sqrt{\|\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}\|^{2}}\Big]$
$\overset{(b)}{\leq}(1-\xi^{\prime})D\sqrt{\mathbb{E}_{\bm{u}^{j}}\big[\|\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j}\|^{2}\big]}$
$\Rightarrow$ $\frac{1}{(1-\xi^{\prime})^{2}D^{2}}\big(\mathbb{E}_{\bm{u}^{j}}\big[\langle\nabla C_{T}(\bm{x}^{j})-\widetilde{\bm{g}}^{j},\bm{x}^{j+1}-\bm{y}\rangle\big]\big)^{2}$
$\leq\mathbb{E}_{\bm{u}^{j}}\Big[\left\|\nabla C_{T}(\bm{x}^{j})\right\|^{2}+\left\|\widetilde{\bm{g}}^{j}\right\|^{2}-2\nabla C_{T}(\bm{x}^{j})^{\top}\widetilde{\bm{g}}^{j}\Big]$
$=\left\|\nabla C_{T}(\bm{x}^{j})\right\|^{2}-2\nabla C_{T}(\bm{x}^{j})^{\top}\nabla\hat{C}_{T}(\bm{x}^{j})+\mathbb{E}_{\bm{u}^{j}}\big[\left\|\widetilde{\bm{g}}^{j}\right\|^{2}\big]$
$=\nabla C_{T}(\bm{x}^{j})^{\top}\big(-\nabla C_{T}(\bm{x}^{j})+2\nabla C_{T}(\bm{x}^{j})-2\nabla\hat{C}_{T}(\bm{x}^{j})\big)+\mathbb{E}_{\bm{u}^{j}}\big[\left\|\widetilde{\bm{g}}^{j}\right\|^{2}\big]$
$=-\left\|\nabla C_{T}(\bm{x}^{j})\right\|^{2}+2\left\|\nabla C_{T}(\bm{x}^{j})\right\|\left\|\nabla C_{T}(\bm{x}^{j})-\nabla\hat{C}_{T}(\bm{x}^{j})\right\|+\mathbb{E}_{\bm{u}^{j}}\big[\left\|\widetilde{\bm{g}}^{j}\right\|^{2}\big]$
$\overset{(c)}{\leq}-\left\|\nabla C_{T}(\bm{x}^{j})\right\|^{2}+\mathbb{E}_{\bm{u}^{j}}\big[\left\|\widetilde{\bm{g}}^{j}\right\|^{2}\big]+G\beta hT(Td+3)^{3/2}\delta^{\prime},$
where (a) follows from Remark 1, (b) follows from Jensen’s inequality and the concavity of $\sqrt{\cdot}$ and (c) follows from Lemma 1(b) and Lemma 2(b). [^6] Note the well-known fact that if a differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ is convex and $G$-Lipschitz, then $\|\nabla f(x)\|\leq G$ for all $x\in\mathbb{R}^{d}$. Thus, we know from Lemma 1(b) that $\|\nabla C_{T}(\bm{x})\|\leq G\sqrt{Th}$ for all $\bm{x}\in\mathbb{R}^{dT}$. It remains to upper bound the term $\|C_{T}(\bm{x}^{j})\|^{2}+\mathbb{E}_{\bm{u}^{j}}\big[\|\widetilde{\bm{g}}^{j}\|^{2}\big]$. Noting that $\sum_{k=s}^{s+h-1}f_{k}(x_{k-h+1:k})$ may be viewed as a function of $x_{s-h+1},\dots,x_{s+h-1}$, denoted as $F_{s}(x_{s-h+1:s+h-1})$, we see from (B.16) that $\widetilde{\bm{g}}^{j}_{s}$ can be equivalently written as
$\widetilde{\bm{g}}^{j}_{s}=\frac{1}{2\delta^{\prime}}\big(F_{s}(\bar{x}^{j}_{s-h+1:s+h-1})-F_{s}(\tilde{x}^{j}_{s-h+1:s+h-1})\big)u_{s}^{j}.$
Moreover, using similar arguments to those in the proof of Lemma 1, one can show that $F_{s}(\cdot)$ is $\beta h$-smooth. Now, following similar arguments to those in the proof of [nesterov2017random, Theorem 4], one can show that
$\left\|\widetilde{\bm{g}}^{j}_{s}\right\|^{2}$ $\leq\frac{\delta^{\prime 2}}{8}\beta^{2}h^{2}\left\|u_{s}^{j}\right\|^{6}+2\left\|u_{s}^{j}\right\|^{4}\left\|\nabla F_{s}(x^{j}_{s-h+1:s+h-1})\right\|^{2}$
$\leq\frac{\delta^{\prime 2}\beta^{2}h^{2}}{8(2(2h-1))^{3/2}}+\frac{1}{(2h-1)}\left\|\nabla F_{s}(x^{j}_{s-h+1:s+h-1})\right\|^{2},$
where the second inequality follows from the fact that $\|u_{s}^{j}\|\leq 1/(2(2h-1))^{1/4}$. One can also show that $\sum_{s=1}^{T}\|\nabla F_{s}(x^{j}_{s-h+1:s+h-1})\|^{2}\leq(2h-1)\|\nabla C_{T}(\bm{x}^{j})\|^{2}$. It follows that
$\left\|\widetilde{\bm{g}}^{j}\right\|^{2}=\sum_{s=1}^{T}\left\|\widetilde{\bm{g}}^{j}_{s}\right\|^{2}\leq\frac{\delta^{\prime 2}\beta^{2}h^{2}T}{8(2(2h-1))^{3/2}}+\left\|\nabla C_{T}(\bm{x}^{j})\right\|^{2}.$
Thus, we can upper bound $\text{(i)}+\text{(ii)}$ as
$\text{(i)}+$ $\text{(ii)}\leq\sqrt{\frac{(1-\xi^{\prime})^{2}D^{2}\delta^{\prime 2}\beta^{2}h^{2}T}{8(2(2h-1))^{3/2}}+(1-\xi^{\prime})^{2}D^{2}Gh\beta T(Td+3)^{3/2}\delta^{\prime}}+\frac{(1-\xi^{\prime})Dh\sum_{t=1}^{T}\varphi_{t}}{\delta^{\prime}(2(2h-1))^{1/4}}$
$\leq\frac{(1-\xi^{\prime})D\delta^{\prime}\beta h\sqrt{T}}{2\sqrt{2}(2(2h-1))^{3/4}}+\sqrt{h\beta GT\delta^{\prime}}(1-\xi^{\prime})D(Td+3)^{3/4}+\frac{(1-\xi^{\prime})Dh\sum_{t=1}^{T}\varphi_{t}}{\delta^{\prime}(2(2h-1))^{1/4}}\triangleq\varepsilon.$
[…]
D.3 Proof of (B.23)
We get from (B.20)-(B.21) and the upper bound $\varepsilon$ in (B.22) that
$\quad C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}^{j})+\gamma(C_{T}(\bm{x}^{j+1})-C_{T}(\bm{x}_{\xi^{\prime}}^{\star}))$
$\leq-\frac{\beta^{\prime}}{2}(1+\gamma)\left\|\bm{x}^{j+1}-\bm{x}^{j}\right\|+\gamma\beta^{\prime}\langle\bm{x}^{j}-\bm{x}^{j+1},\bm{x}^{j}-\bm{x}_{\xi^{\prime}}^{\star}\rangle-\frac{\mu\gamma}{2}\left\|\bm{x}^{j}-\bm{x}_{\xi^{\prime}}^{\star}\right\|^{2}+(1+\gamma)\varepsilon$
$\leq-\frac{1}{\beta^{\prime}-\mu}\left\|\beta^{\prime}(\bm{x}^{j+1}-\bm{x}^{j})-\mu(\bm{x}^{j}-\bm{x}_{\xi^{\prime}}^{\star})\right\|^{2}+(1+\gamma)\varepsilon\leq(1+\gamma)\varepsilon,$
[…]
Appendix E Technical Lemma
Lemma E.1.
[bansal2019potential, Proposition 2.2] Given a convex set $\mathcal{X}\subseteq\mathbb{R}^{d}$, let $a\in\mathcal{X}$, $b^{\prime}\in\mathbb{R}^{d}$ and $b=\mathbf{\Pi}_{\mathcal{X}}(b^{\prime})$. Then $\langle a-b,b^{\prime}-b\rangle\leq 0$ and $\left\|a-b\right\|^{2}\leq\left\|a-b^{\prime}\right\|^{2}$.